### v1

# TIPO: Text to Image with Text Presampling for Prompt Optimization

Abstract
TIPO (Text to Image with text pre-sampling for Prompt Optimization) is an innovative framework designed to enhance text-to-image (T2I) generation by language model (LM) for automatic prompt engineering. By refining and extending user-provided prompts, TIPO bridges the gap between simple inputs and the detailed prompts required for high-quality image generation.
Unlike previous approaches that rely on Large Language Models (LLMs) or reinforcement learning (RL), TIPO adjusts user input prompts with the distribution of a trained prompt dataset, eliminating the need for complex runtime cost via lightweight model. This pre-sampling approach enables efficient and scalable prompt optimization. Experimental results demonstrate TIPO’s effectiveness in improving aesthetic scores, reducing image corruption, and better aligning generated images with dataset distributions. These findings highlight the critical role of prompt engineering in T2I systems and open avenues for broader applications of automatic prompt refinement.
1 Introduction
Recent advancements in Text-to-Image (T2I) generative models have revolutionized creative applications [36, 6, 19, 41, 38, 39, 43, 40, 37, 42, 12, 11, 28, 20, 7]. These models tend to perform better when provided with longer, more detailed prompts, which describes specific elements like style, composition, and context. However, this reliance on highly descriptive inputs is a significant bottleneck for generating high-quality images, especially for users who prefer or require simpler inputs. The need for intricate, nuanced prompts often results in a higher barrier to entry, limiting accessibility and ease of use for those unfamiliar with the complexities of prompt engineering. The previous works [31] tried to utilize LLMs to lenghthen prompt to solve tendency, however many ignore the root cause of why some prompts result in worse images and rely heavily on manually engineered prompts, which require substantial resources.
To address these challenges, we introduce TIPO (Text to Image with text pre-sampling for Prompt Optimization), an innovative framework designed to enhance T2I generative models. TIPO introduces ‘text pre-sampling’ preprocess procedure in inference pipeline, letting text-to-image models reproduce a more accurate distribution of corresponding user prompts without additional manual engineering.
Unlike previous approaches that focus on extending prompts with manually collected prompt sets [4, 15, 45, 21], directly use LLM to modify the prompts [26, 51] or utilize reinforcement learning like method which update model’s input or model itself from text-to-image feedback [22, 13, 31] without addressing underlying issues, TIPO introduces a method that optimizes prompts efficiently as shown Section 4.
Our work makes the following key contributions:
  1. 1.
    We introduce an universal prompt optimization framework that is designed to work with any text-to-image model, providing a versatile solution for improving image generation across different architectures.
  2. 2.
    We provide comprehensive evaluation methodology and results including using both standard and non-standard metrics. Our experiments demonstrate that TIPO’s prompt engineering method can improve image quality in terms of mathematical measures and user preferences, validating its effectiveness from multiple perspectives.
Figure 1: Comparison of prompt optimization methods. (a) uses instructions for prompt refinement but lacks understanding of T2I-specific prompts. (b) improves detail but limits variety. (c) optimizes for model feedback but are computationally expensive. (d) TIPO (ours) aligns prompts with the trained distribution of T2I models, ensuring prompt detail, consistency with training data, and controlled variety within model scope.
This paper presents TIPO’s theoretical foundations, implementation details, and experimental results demonstrating its effectiveness in improving aesthetic scores [17], reducing image corruption [32, 33], and better aligning generated images with dataset distributions [24, 35]. Our findings highlight the critical role of prompt optimization in advancing T2I technology and open avenues for broader applications of automatic prompt refinement.
2 Related Works
2.1 Direct Prompt Refinement Through LLMs
With the advent of advanced LLMs, several projects have leveraged their few-shot in-context learning capabilities to optimize prompts for Text-to-Image (T2I) models. For instance, CogView3 [51] and the work by Lee et al. [26] utilize GPT-J and Text Style Transfer (TST) techniques for prompt refinement. However, this approach relies heavily on the LLM’s inherent knowledge of visual content descriptions, which can be unpredictable and may not always align with specific T2I model requirements.
2.2 Extending Prompts with Manually Collected Data
In the open-source T2I community, a common approach involves using curated datasets of “good prompts” to fine-tune or train LLMs for prompt extension [4, 45, 15]. These models are typically trained on image-text pairs collected from users of proprietary T2I services like Midjourney [^1] https://www.midjourney.com or Dall-E 3 [6]. While this method is straightforward and somewhat effective, it often lacks the versatility to adapt to diverse T2I models beyond the one used in the original dataset collection.
2.3 Image Output-Based Tuning
Recent works have explored more sophisticated prompt optimization techniques based on image outputs.
Reinforcement Learning Approaches: Promptist [22] introduced a two-stage framework combining supervised fine-tuning with reinforcement learning (RL). The RL stage incorporates aesthetic and relevance scores as environmental feedback, enabling the model to learn optimal prompt refinement strategies.
Iterative Optimization: OPT2I [31] proposed a backpropagation-free prompt optimization framework. This method uses scorers and LLMs to iteratively update input prompts, aiming to improve aesthetic quality and coherence.
These RL-based or RL-like approaches allow for prompt refinement tailored to specific model properties or user preferences. However, they often require extensive computational resources and may not generalize well across different T2I models or datasets.
While these methods have shown promise, they often face challenges such as difficulty in maintaining alignment with original user intent, challenges in collecting manually engineered prompts, high computational costs for RL-based methods, and limited generalization across diverse T2I models. By contrast, TIPO offers an efficient approach to prompt optimization by directly leveraging the training dataset, removing the need for manual prompt engineering and runtime tuning, and achieving versatile, contextually aligned image generation.
Figure 2: Comparison of prompt optimization methods for a T2I model [7]. (a) yields a basic image with limited detail. (b) provides specific output but requires manual engineering. (c) introduce nonsensical elements (red boxes) in the teal region, where the model’s actual output distribution for a given prompt (green) and all the image distribution (blue) overlap outside the user’s expected output distribution (red). (d) captures user intent accurately, balancing detail and diversity.
3 Prompt Optimization
3.1 Motivation and Challenges
Nowadays text-conditioned models [7, 37, 19] generate outputs based on user-provided text inputs. However, when users seek high-quality outputs or wish to explore diverse outputs while preserving key aspects of their input, they often need to manually expand their prompts. This manual prompt expansion can be time-consuming and labor-intensive. Moreover, simply creating a longer prompt does not guarantee improved results; it must be crafted in line with the model’s specific prompting conventions and interpretive nuances.
To address these challenges, previous works have aimed to automate the prompt optimization process. Some methods attempt to expand prompts using the model’s familiar vocabulary and stylistic tendencies [6], while others rely on direct feedback from the model to iteratively improve prompt alignment [22]. However, these approaches have limitations: if the structure or language does not fully align with the model’s prompting conventions, performance may degrade, and feedback-based approaches often involve significant computational costs.
Therefore, effective prompt optimization for text-conditioned models requires accurately capturing user intent and expanding the prompt in a way that is both precise and aligned with the model’s conventions for interpreting text inputs. This need for precision and alignment motivates a systematic approach to prompt formulation, which we define in the following section.
3.2 Formulation
To formalize prompt extension, we define $\mathcal{P}$ as the set of all possible prompts, $\mathcal{N}$ as the Gaussian noise vector space, and $\mathcal{I}$ as the set of all possible images. A T2I model then acts as a mapping function:
$f(p):\mathcal{N}\rightarrow\mathcal{I}_{p}$
where, for any prompt $p\in\mathcal{P}$, the model maps noise vectors from $\mathcal{N}$ to a subset of images $\mathcal{I}_{p}\subseteq\mathcal{I}$ associated with the prompt $p$.
Now we define a pair of prompts: simple prompt $p_{s}$ and detailed prompt $p_{d}$. Formally, for a simple prompt $p_{s}$ and a detailed prompt $p_{d}$ that extends $p_{s}$, we have the following relationship:
$f(p_{d})(\mathcal{N})\subseteq f(p_{s})(\mathcal{N})$
This formulation implies that the image subset generated by the detailed prompt $p_{d}$ is contained within the subset generated by the simple prompt $p_{s}$, thereby preserving the relevance of the initial prompt while adding specificity.
To further encapsulate the prompt generation process, we define the prompt extension function, $\mathcal{E}$. This function takes a simple prompt $p_{s}$ and additional metadata $\mathcal{M}$ (such as tags) as inputs, and produces a detailed prompt $p_{d}$ as output:
$\mathcal{E}:\mathcal{P}_{s}\times\mathcal{M}\rightarrow\mathcal{P}_{d}$
where $\mathcal{P}_{s}$ is the space of simple prompts and $\mathcal{P}_{d}$ is the space of detailed prompts. Here, $\mathcal{M}$ provides auxiliary information that guides the extension of $p_{s}$ into a more refined prompt $p_{d}$.
For each instance of prompt extension, we apply the function as follows:
$p_{d}=\mathcal{E}(p_{s},M)$
where $M$ could represent a specific metadata component, such as $T_{s}$, or other forms of contextual information depending on the input type.
The simplest approach to extend a simple prompt is by adding random words. However, this approach has limitations, as it often introduces internally conflicting information that results in out-of-distribution (OOD) prompt inputs, leading the model to generate nonsensical images. In such cases, although each added word may be meaningful in isolation, these random combinations can create prompts that are internally contradictory (and are not simply creative combinations). Thus, while random word extension typically yields unreliable results, it may occasionally produce a usable prompt by chance. To represent this random extension method, let $p_{r_{i}}$ denote the $i^{th}$ prompt created by adding random words to $p_{s}$. The resulting union of distributions fails to approximate the desired distribution.
$\bigcup_{i=0}^{n}f(p_{r_{i}})(\mathcal{N})\supseteq\mathcal{I}\supseteq\mathcal{I}_{p_{s}}$
Which means:
$\bigcup_{i=0}^{n}f(p_{r_{i}})(\mathcal{N})\not\approx\mathcal{I}_{p_{s}}$
This scenario is illustrated in Figure 2, where “Simple Prompt + Random Words” shows a scattered distribution extending beyond the bounds of the desired distribution or even the “image” distribution itself.
4 TIPO Framework
4.1 Overview
TIPO introduces a dataset-driven approach to prompt optimization by training directly on the existing dataset. This avoids reliance on manually engineered prompts [4, 45, 15] or on computationally intensive RL techniques [22, 31]. Methods using LLMs rely on general language knowledge, which may misalign with specific T2I model needs [6, 51]. In contrast to previous approaches, TIPO captures the model’s training data distribution closely.
4.2 Distribution-Aligned Prompt Extension
Central to TIPO’s method is an extension function $\mathcal{E}$, which iteratively refines prompts to enhance both the diversity and quality of generated images while remaining aligned with the intent of the original user-provided prompt and the training data distribution. By applying $\mathcal{E}$ to a simple prompt $p_{s}$ to generate a detailed prompt $p_{d}$, TIPO aims to approximate the target distribution:
$\bigcup_{i=0}^{n}f(p_{t_{i}})(\mathcal{N})\approx\mathcal{I}_{p_{s}}$
where $p_{t_{i}}$ denotes the $i^{th}$ TIPO-generated prompt derived from $p_{s}$, and $\mathcal{I}_{p_{s}}$ represents the target distribution of image outputs associated with $p_{s}$. Through this systematic prompt refinement, Figure 2 “Simple Prompt + TIPO” illustrates the broader image space exploration enabled by TIPO while retaining relevance to the original prompt, effectively balancing diversity with coherence.
4.3 Prompt Structure
The construction of simple prompts, $p_{s}$, and their corresponding detailed prompts, $p_{d}$, varies depending on the dataset and input modality. We consider two primary scenarios: tag-based captions and natural language (NL) captions.
Tags are used to describe the content of images and are commonly formatted as a comma-separated string when input into T2I models. For a given image-tag pair, let $T_{n}=\{t_{1},t_{2},t_{3},\ldots,t_{n}\}$ represent the complete set of tags for an image. We derive the simple tag set $T_{s}=\{t_{1},t_{2},\ldots,t_{m}\}$ where $m<n$, and define the detailed tag set as $T_{d}=T_{n}$. The simple and detailed tag prompts are then constructed as:
$p_{s}=\text{concat}(T_{s}),\quad p_{d}=\text{concat}(T_{d})$
NL captions are typically generated by LMs with vision modality [29, 1, 2, 27, 5, 14, 48, 16] and may contain multiple sentences. Each image is associated with either two types of NL captions—a short NL caption $S_{s}$ providing a brief, general description, and a long NL caption $S_{d}$ offering an extended, detailed description—or a single long caption that can be segmented. For images with a single long caption, let
$S_{d}=\{\text{sentence}_{1},\text{sentence}_{2},\text{sentence}_{3},\ldots,\text{sentence}_{n}\}$
represent the complete set of sentences in $p_{d}$. We derive the simple sentence set
$S_{s}=\{\text{sentence}_{1},\text{sentence}_{2},\ldots,\text{sentence}_{m}\}$
where $m<n$, and define the detailed sentence set as $S_{d}=S_{n}$. The simple and detailed NL prompts are then constructed as:
$p_{s}=\text{concat}(S_{s}),\quad p_{d}=\text{concat}(S_{d})$
These ensure that $p_{d}$ is a superset of $p_{s}$, aligning with the concept that detailed prompts provide a more specific description.
To generate $p_{d}$ from $p_{s}$, TIPO employs two distinct formats based on the relationship between $p_{s}$ and $p_{d}$:
  1. 1.
    When $p_{s}$ is a substring of $p_{d}$: The format <meta> <p_d> is used. As TIPO model is an autoregressive LM, $p_{d}$ is directly used after the <meta> tokens, allowing the model to generate subsequent tokens based on any substring of $p_{d}$.
  2. 2.
    When $p_{s}$ is not a substring of $p_{d}$: The format <meta> <p_s> <p_d> is employed. Since $p_{s}$ is not contained within $p_{d}$, both are included separately after the <meta> tokens to guide the generation process effectively.
4.4 Prompt Formatting
In TIPO, we define a simple and consistent format for the content of the <meta> token or the simple prompt $p_{s}$. The format is as follows: <Category>: <Content>. Metadata categories might include artist, copyright, aspect ratio, quality, and year. For instance,
quality: masterpiece, year: newest
Importantly, in certain cases, input tag or NL prompts can be treated as metadata <Content> as well, when the task involves generating tags or NL prompts conditioned on each other.
4.5 An Example Scenario and Workflow
Figure 3: (a) A generated image and prompts. (b) Prompt optimization of TIPO, starting from user input (short tag prompt $p_{s}$) to final output (detail prompt $p_{d}$). The shading from gray to light sky blue represents an increase in context richness in the prompt.
Figure 3 illustrates a scenario where both tag captions $T_{s}$ and short NL captions $S_{s}$ are available. In such cases, TIPO processes each input type separately to maintain clarity and coherence. During each generation cycle, one input type is treated as the primary prompt which is target prompt for extending or refining, while the other types are used as <Content> of metadata to enrich the context. Figure 3 provides an example of this setup, where:
$\begin{array}[]{ll}S_{s}&=\{\textsf{A young girl with long hair...}\},\quad\verb|<meta>|=\emptyset{}\\ T_{s}&=\left\{\textsf{outdoors, scenery, water, wind, landscape, ...}\right\}\\ \end{array}$
The generation proceeds sequentially as follows:
  1. 1.
    short_to_tag: TIPO uses $T_{s}$ as the primary prompt to generate a detailed tag sequence $T_{d}$
  2. 2.
    tag_to_long: $T_{d}$ is incorporated into the metadata, and TIPO produces a refined short NL prompt $S_{s}$ based on $T_{d}$.
  3. 3.
    short_tag_to_long: With both $T_{d}$ and $S_{s}$ in the metadata, TIPO generates a comprehensive long NL prompt $S_{d}$, ensuring a more detailed prompt.
  4. 4.
    TIPO aggregates $T_{d}$, $S_{d}$, and any additional metadata to construct a context rich prompt, $p_{d}$.
This example highlights how TIPO leverages an autoregressive LM to perform prompt extension in a stepwise, sequential processes, through short_to_tag, tag_to_long, and short_to_tag_to_long tasks within a single generation cycle. More detailed task definitions for TIPO can be found in Section 5.2. By iteratively refining prompts in this structured way, TIPO addresses the limitations of simplistic or randomly extended prompts, enabling it to generate high-quality, diverse images that closely align with user intent.
4.6 Versatile and Scalable Design of TIPO
The example in Section 4.5 demonstrates how TIPO’s structured approach enables it to create coherent, intent-aligned prompts. Building on this foundation, TIPO’s design extends smoothly to a wide range of models and datasets, showcasing its versatility and scalability.
A single TIPO model trained on dataset D can be applied to all text-to-image models trained on the same dataset, enabling broad applicability across different architectures and versions that share the same training data. Similarly, a TIPO model trained using caption method M (e.g. tag, NL captions) can be used with any T2I model trained with captions from method M. Experimental results in Section 6 further illustrate this adaptability, as TIPO successfully applies across multiple model architectures and training configurations, confirming its broad applicability and robustness.
5 Training TIPO model
In this section, we detail how to build the TIPO model based on the TIPO framework. For various T2I models [36, 6, 19, 41, 38, 39, 43, 40, 37, 42, 12, 11, 28, 20, 7], tag and NL captions are widely used captioning methods. To allow TIPO model to seamlessly transform captions between different formats or refine captions within the same format, we define specific tasks that enable flexible and precise prompt construction.
5.1 Modeling
TIPO adopts a causal autoregressive LM framework as prompt extension function, $\mathcal{E}$, training to extend and refine user-provided prompts into detailed, content-rich prompts $p_{d}$ from simpler inputs $p_{s}$, thereby effectively balancing specificity with alignment to user intent. This process captures a subset of outputs closely aligned with the user’s original intent, ensuring relevance. Through use of existing datasets for prompt optimization, along with its iterative extension function and application of LM, TIPO achieves significant coverage and diversity in generating high-quality images aligned with training distributions, all without the need for manual prompt engineering.
5.2 Task Definitions
TIPO encompasses three primary tasks: extending tag sequences, extending NL prompts, and generating refined NL prompts. To facilitate these tasks, we define several specific task types and corresponding formatting methods during training:
  1. 1.
    tag_to_long: Use tags as metadata to generate a new NL prompt or extend a user-provided NL prompt.
  2. 2.
    long_to_tag: Use an NL prompt as metadata to extend a tag sequence.
  3. 3.
    short_to_tag: Use the simple prompt $p_{s}$ as metadata to extend a tag sequence.
  4. 4.
    short_to_long: Use a user-input NL prompt as metadata to generate a refined detailed prompt $p_{d}$.
  5. 5.
    short_to_tag_to_long: Use a user-input NL prompt or tag sequence as metadata to generate a refined detailed prompt $p_{d}$.
  6. 6.
    short_to_long_to_tag: Use a user-input NL prompt or generated NL prompt as metadata to extend a tag sequence.
  7. 7.
    tag_to_short_to_long: Use user-input tags or NL prompts as metadata to generate a refined detailed prompt $p_{d}$.
By defining these task types, we can achieve special cases where, for instance, inputting a short description or a short tag sequence results in a full-size tag sequence and NL prompt in a single pass (i.e., generating $T_{d}$ and $p_{d}$ simultaneously) as shown in Section 4.5.
5.3 Model and Training Details
We utilize the LLaMA architecture [46, 47] with models of 200 million and 500 million parameters:
  • TIPO-200M: Pretrained on the Danbooru2023 [34, 49] and GBC10M [25] datasets for 5 epochs, then fine-tuned on the Danbooru2023, GBC10M, and CoyoHD11M [9] datasets for 3 epochs, resulting in a total of approximately 40 billion non-padding tokens seen.
  • TIPO-500M: Pretrained on the Danbooru2023, GBC10M, and CoyoHD11M datasets for 5 epochs, resulting in a total of approximately 30 billion non-padding tokens seen.
Note that we calculate the token count based on non-padding tokens only. Given the relatively short and variable lengths of data entries, the effective token count may be lower than anticipated.
5.4 Training Procedure
In our experiments, each training pass generates a new prompt of a single type. To cover various prompt types—such as a refined prompt $p_{d}$, an extended NL prompt $S_{d}$, and an extended tag sequence $T_{d}$—the model requires at least three passes. For each dataset entry, we randomly select one of the seven task types to apply and randomly decide how to split the simple prompt from the detailed prompt ($p_{s}$ from $p_{d}$) and the simple tag set from the detailed tag set ($T_{s}$ from $T_{d}$).
This randomization in task selection and prompt splitting effectively increases the training dataset’s size, allowing the model to encounter a broader range of inputs and outputs. This approach enhances the model’s ability to generalize across the various tasks defined in TIPO.
(a) Direct LLM
(b) Prompt DB
(c) RL (Promptist)
(d) TIPO (Ours)
Figure 4: Scenery test examples from 4 different methods.
6 Evaluation
In this section, we detail the experimental setup used to evaluate the TIPO framework. We outline the formatting conventions for prompts and metadata, define the various tasks employed during training, and describe the models and datasets utilized in our experiments.
6.1 Experimental Setup
To assess the quality of generated prompts and resulting images, we employed the following evaluation metrics.
FDD (Frechet Dino Distance): Traditionally, the Frechet Inception Distance (FID) [24] has been used to measure the distribution distance between datasets and generated images. However, recent works have shown that FID is not always aligned with human preferences [44]. Therefore, we measure the Frechet Distance using DinoV2 [35] outputs. In our experiments, we used the pretrained ViT-g [18] model from DinoV2 and used the normalized embedding vector to measure the Frechet Distance result.
Aesthetic Score: We compute the Aesthetic Score using the Aesthetic Predictor V2.5 [17]. Higher aesthetic scores indicate images that are perceived as more visually appealing, well-composed, and of higher quality. This metric aims to quantify the subjective visual appeal of generated images, with higher scores suggesting outputs that are more likely to be judged favorably by human viewers.
AI Corrupt Score: The AI Corrupt Score is obtained from the AICorruptMetrics [32] in sdeval [33]. This metric is trained on AI-generated images with human-annotated “corrupt or not” labels. A higher value indicates a higher likelihood of the image being “correct” or “complete”.
All results presented in this section are tested on the TIPO-200M model.
6.2 Scenery Tag Test
In this test, we randomly select 32,768 entries from Danbooru2023 [50, 34] that include the “scenery” tag. We set up the following inputs:
  • “Scenery” + Meta:
    • Retain all metadata categories as described earlier.
    • Only include “scenery” as the content tag.
  • “Scenery” + Meta + TIPO:
    • Use “scenery” + meta as input.
    • Extend $T_{s}$ (scenery) to $T_{d}$ and generate $p_{d}$ from $T_{d}$.
For the scenery tag test, we generate one image from each prompt using the Kohaku-XL-zeta [^2] https://huggingface.co/KBlueLeaf/Kohaku-XL-Zeta model, which is a fine-tuned SDXL [37] model on the Danbooru dataset [34].
Figure 4 illustrates the comparative visual result by various methods. The results reveal significant biases in Prompt DB-based [21] and RL-based [22] approaches. The former consistently generates snowy mountain scenes, while the latter produces near-future fantasy scenery, often resulting in corrupted images. The Direct LLM method, utilizing GPT-4o [2], shows improved results but still exhibits notable bias; out of 9 generated prompts, 8 exclusively describe natural scenes only without any human-made structures.
In contrast, TIPO demonstrates superior performance in reproducing the dataset’s distribution. While primarily generating nature scenes, it also incorporates architectural elements and artifacts, more accurately reflecting the diverse content of the original dataset. This balanced output suggests that TIPO effectively mitigates the biases observed in other methods.
Metric Original LLM Prompt DB Promptist TIPO (Ours)
FDD $\downarrow$ 0.3558 0.5414 0.3247 0.2350 0.2282
Aesthetic Score $\uparrow$ 5.0569 6.3676 6.1609 5.9468 6.2571
AI Corrupt Score $\uparrow$ 0.4257 0.7490 0.5024 0.5669 0.9195
Table 1: Metric results on Scenery Tag Test.
As shown in Table 1, applying TIPO significantly improves FDD performance and AI Corrupt score. The improvement in FDD score suggests that TIPO-generated prompts lead to images that more closely align with the original dataset distribution. Furthermore, the enhanced AI Corrupt score indicates that TIPO aids in generating more coherent images, potentially reducing corruption or distortion in the output.
The results from the direct LLM approach (GPT4o-mini) exemplify a key concern in prompt engineering: while achieving high aesthetic scores, it performs poorly in terms of FDD. This discrepancy underscores the potential pitfall of using LLMs for prompt generation without considering dataset alignment. Such models may produce well-constructed prompts that yield visually appealing images but fail to capture the nuanced distribution of the text-to-image dataset. In contrast, TIPO demonstrates a balance between aesthetic quality and dataset fidelity, offering a more robust solution for prompt optimization in text-to-image generation tasks.
6.3 Short/Truncated Long Test
This test evaluates TIPO’s performance on enhancing natural language prompts which are already contains some useful info for model to generate reasonable images. We randomly selected 10k entries each from [9] and [25]. For each entry, we extracted the short caption and the truncated long caption (limited to a maximum of two sentences). These captions serve as user input prompts for our test. Using these prompts, we generated one image per prompt with the SDXL-1.0-base [^3] https://huggingface.co/stabilityai/stable-diffusion-xl-base-1.0 model [37].
Short Original LLM Prompt DB Promptist TIPO(Ours)
FDD $\downarrow$ 0.0957 0.1668 0.0980 0.1783 0.1168
Aesthetic Score $\uparrow$ 5.8370 6.0589 5.8213 5.7963 5.8531
AI Corrupt Score $\uparrow$ 0.7113 0.6985 0.7064 0.6314 0.7131
Truncated Long Original LLM Prompt DB Promptist TIPO(Ours)
FDD $\downarrow$ 0.0955 0.1683 0.1247 0.2096 0.1210
Aesthetic Score $\uparrow$ 5.7497 6.0168 5.8191 5.7759 5.8364
AI Corrupt Score $\uparrow$ 0.6868 0.6712 0.6741 0.5925 0.7130
Table 2: Metric result in Short/Truncated-long test of prompt generation
Table 2 presents the results of our Short/Truncated-long test for prompt generation. Unlike the scenery tag test (Section 6.2), we observe that the original prompts in this test already produce decent images. Interestingly, all four methods result in worse FDD scores compared to the original prompts, suggesting that the prompt generators introduced new elements not present in the selected 10,000 entries. Given the size of the training set, this outcome is not entirely unexpected.
Examining the Aesthetic and AI Corrupt Scores, we notice that the Direct LLM approach achieves the highest aesthetic score, mirroring the results from the scenery tag test. However, its AI Corrupt Score and FDD value indicate potential issues with OOD inputs, which may occasionally hinder the model’s ability to generate coherent images from its prompts.
TIPO, on the other hand, demonstrates a more balanced performance. It achieves the best AI Corrupt Score, the second-best Aesthetic Score, and a reasonable FDD value. This balanced profile suggests that TIPO is the most suitable choice for real-world applications of modern text-to-image models, where a combination of aesthetic appeal, image coherence, and distribution alignment is crucial.
7 Conclusion
In this paper, we introduced TIPO (Text to Image with text Presampling for Prompt Optimization), a novel framework designed to enhance the quality of Text-to-Image (T2I) models through automatic prompt engineering. By leveraging Large Language Models (LLMs) to extend and refine user-provided prompts, TIPO effectively bridges the gap between simple user inputs and the detailed, content-rich prompts required for high-quality image generation.
Our comprehensive evaluation demonstrates TIPO’s effectiveness across multiple dimensions of image quality and fidelity. The framework consistently produced prompts that led to images closely aligned with original dataset distributions, while simultaneously improving aesthetic appeal and reducing the likelihood of corrupted or nonsensical outputs. This balanced performance underscores TIPO’s robustness and suitability for real-world applications of T2I models, where a combination of distribution alignment, visual appeal, and image coherence is crucial.
Notably, our experiments highlight the critical role of prompt engineering in advancing T2I technology. The improvements achieved through TIPO’s prompt modifications often rivaled or surpassed performance differences between various T2I model architectures. This finding emphasizes that effective prompt engineering can significantly enhance model performance without necessitating architectural changes, potentially offering a more efficient path to improving T2I systems.
The success of TIPO opens several promising avenues for future research:
  1. 1.
    Broader Application: Exploring TIPO’s applicability to other generative tasks, such as text generation or audio synthesis, could further demonstrate the versatility and impact of automatic prompt engineering.
  2. 2.
    Interactive Systems: Incorporating TIPO into interactive applications where users can iteratively refine prompts in real-time may enhance user experience and enable more precise control over generated content.
  3. 3.
    Advanced Alignment Techniques: Investigating more sophisticated methods for aligning user inputs with dataset distributions could further enhance the performance and reliability of generative models.
  4. 4.
    Cross-Modal Applications: Adapting TIPO’s principles to facilitate cross-modal generation tasks, such as text-to-video or image-to-text, could expand its utility in multimodal AI systems.
In conclusion, TIPO represents a significant advancement in the field of text-to-image generation, offering a versatile and efficient solution for prompt optimization. By automating the refinement and extension of user prompts, TIPO not only improves the quality and specificity of generated images but also highlights the importance of aligning user inputs with underlying data distributions. As generative models continue to evolve, frameworks like TIPO will be instrumental in unlocking their full potential, ensuring they meet diverse user needs with precision and creativity while pushing the boundaries of what’s possible in AI-generated content.
Appendix
Appendix A Dataset/Resource
A.1 Danbooru2023
The Danbooru2023 dataset [50, 49, 34] is an extensive collection of images and their corresponding tags, compiled from the Danbooru image board. This dataset includes images annotated with highly specific and detailed tags, providing a rich resource for training both the Text-to-Image (T2I) and Large Language Models (LLMs) involved in the TIPO framework. The dataset contains data up to image ID 7,349,999, encompassing a wide variety of visual content with granular annotations. These annotations allow for the creation of nuanced and precise prompts, ensuring that longer, more detailed prompts can indicate subsets of shorter prompts.
Key Characteristics:
  • Rich Annotations: Danbooru2023’s detailed tags enable the differentiation of subtle variations in image content, such as “long hair” versus “very long hair,” which is crucial for the specificity required in high-quality image generation.
  • Large Volume: The dataset’s extensive size provides a robust foundation for training models, ensuring that they can learn from a diverse array of images and annotations.
  • Tag-Based Prompting: Utilizing the detailed tags from Danbooru2023, the LLM can generate refined prompts that lead to more accurate and high-quality image generation by the T2I model.
A.2 GBC10M
The GBC10M dataset [25] is a large-scale collection of 10 million images sourced from CC12M [10], annotated using the Graph-Based Captioning (GBC) approach. Each image in GBC10M is represented by a graph where nodes correspond to object regions, compositions, and relations, and edges define the hierarchical relationships among them. Annotations are generated automatically through a pipeline that leverages pretrained multimodal large language models (MLLM) and object detection tools. The GBC structure enhances traditional image captions by providing both detailed descriptions and structural information, improving model performance in downstream tasks. All data is provided in JSON lines format, containing image URLs, bounding boxes, and caption annotations.
In TIPO, we only utilize the root node from GBC10M, since it provides both detail and short captions.
A.3 Coyo HD 11M
The Coyo HD 11M dataset [9] consists of 11.4 million high-density, high-definition images, with 22.8 million synthetic captions generated from the Coyo-700M dataset. The ”HD” refers to both the high resolution of images (filtered to have at least 512 pixels on the shortest edge) and the high concept density, addressing the issues found in previous alt-text image pair datasets, which often included low-quality or low-concept images. Synthetic captions for this dataset were produced using the LLaVA-Next-8B model [30] built upon LLaMA 3 [3] with post-processing for conciseness and clarity. This dataset provides high-quality visual and textual pairs, ideal for training advanced vision-language models.
In TIPO, we utilize the short/long captions, booru tags, and open images tags from Coyo HD 11M.
Appendix B Implementation Detail
In this appendix, we provide all the missing details about our dataset construction process that are not mentioned in Section 4.4 and 5.2.
B.1 Training Settings
TIPO-200M stage1 TIPO-200M stage2 TIPO-500M
Architecture LLaMA
Type Pretrain Finetune Pretrain
Vocab Size 32013 - 32013
Hidden Dim 768 - 1280
Attention Heads 12 - 20
MLP Dim 2304 - 3840
Hidden Layers 20 - 20
Model Parameters 203M - 508M
Max Learning Rate 2e-4 5e-5 2e-4
Optimizer AdamW AdamW AdamW
betas 0.9, 0.98 0.9, 0.98 0.9, 0.98
weight decay 0.01 0.01 0.01
Dataset GBC, Dan Coyo, GBC, Dan Coyo, GBC, Dan
max context length 512 1024 1024
global batch size 2048 2048 3584
Token Seen 22.625B 18.339B 31.274B
Hardware 4 x RTX3090 4 x RTX3090 8 x H100
Training Time (wall) 150 hour 270 hour 100 hour
Table 3: Training settings and details for TIPO-200M and TIPO-500M.
B.2 Inference Pipeline
The TIPO inference pipeline is designed to handle various input types and scenarios, combining different tasks to achieve refinement or expansion of both tag-based prompts and natural language prompts. Figure 5 illustrates this comprehensive workflow.
Figure 5: The inference workflow in TIPO, demonstrating how different task combinations are used to refine or expand tag-based and natural language prompts.
Our framework processes tags and natural language inputs separately, allowing for specialized handling of each input type. This flexible pipeline allows TIPO to adapt to various input scenarios, whether the user provides tags, natural language descriptions, or both. By leveraging different task combinations, TIPO ensures that both tag-based and natural language prompts are optimized, resulting in more detailed and effective inputs for text-to-image models.
Appendix C Evaluation Statistics
In this appendix, we provide more statistics for the result obtained in Section 6.
C.1 Scenery Tag Test
(a) The box plot for the Aesthetic Score result of scenery tag test.
(b) The box plot for the AI Corrupt Score result of scenery tag test.
Figure 6: The box plot for scenery tag test.
The box plot shown in Figure 6 presents the aesthetic scores and AI corruption scores for results from the scenery tag test described in Section 6.2. The analysis clearly demonstrates that TIPO significantly outperforms all other methods, with a substantial margin of improvement.
C.2 Short/Truncated Long Test
Figures 7 and 8 display the box plots of aesthetic scores and AI corruption scores obtained from the short/truncated long test detailed in Section E.2. While the box plots reveal that the differences in performance between various methods are subtle, the AI corruption scores provide valuable insights. Specifically, these scores indicate that implementations supported by TIPO produce more stable output images compared to other methods.
(a) The box plot for the Aesthetic Score result of short caption input.
(b) The box plot for the AI Corrupt score result of short caption input.
Figure 7: The box plot for short caption input.
(a) The box plot for the Aesthetic Score result of truncated long caption input.
(b) The box plot for the AI Corrupt score result of truncated long caption input.
Figure 8: The box plot for truncated long caption input.
Appendix D TIPO example
In this appendix, we provide some text example of TIPO’s input and output.
TIPO Format template User Input:
1girl, ciloranko, maccha (mochancc), ningen mame, ask (askzy), solo, masterpiece, absurdres, newest, safe A girl sits in a cozy cafe, cradling a cup of coffee in her hand
Formatted TIPO Input For Expand Tags:
meta: absurdres rating: safe artist: ciloranko, maccha (mochancc), ningen mame, ask (askzy) quality: masterpiece, newest aspect_ratio: 1.0 target: <|short|> <|short_to_tag|> short: A girl sits in a cozy cafe, cradling a cup of coffee in her hand tag: 1girl, solo
Formatted TIPO Output after Expand Tags and Expand Natural Prompt:
meta: absurdres rating: safe artist: ciloranko, maccha (mochancc), ningen mame, ask (askzy) quality: masterpiece, newest aspect_ratio: 1.0 target: <|short|> <|tag_to_long|> tag: 1girl, solo, sitting, closed mouth, jewelry, long hair, looking at viewer, crossed legs, plant, table, couch, bracelet, cup, smile, teacup, indoors, blue eyes, blonde hair, holding long: A girl sits in a cozy cafe, cradling a cup of coffee in her hand. The cafe has large windows with green plants on the walls and a wooden table in front of her. Behind her is a staircase leading to another room. The overall atmosphere of the image is serene and inviting.
Formatted Output For Text-to-Image
1girl, ciloranko, maccha \(mochancc\), ningen mame, ask \(askzy\), solo, sitting, closed mouth, jewelry, long hair, looking at viewer, crossed legs, plant, table, couch, bracelet, cup, smile, teacup, indoors, blue eyes, blonde hair, holding, A girl sits in a cozy cafe, cradling a cup of coffee in her hand. Behind her is a staircase leading to another room. The cafe has large windows with green plants on the walls and a wooden table in front of her. The overall atmosphere of the image is serene and inviting. masterpiece, newest, absurdres, safe
Refer to caption User InputRefer to caption TIPO Output
Figure 9: An example formatted content we used for training and inference in TIPO.
TIPO Format template User Input:
scenery, no humans, masterpiece, absurdres, newest, safe
Formatted TIPO Input For Expand Tags:
meta: absurdres rating: safe quality: masterpiece, newest aspect_ratio: 1.0 target: <|long|> tag: scenery, no humans
Formatted TIPO Output after Expand Tags and tag_to_long task:
meta: absurdres rating: safe quality: masterpiece, newest aspect_ratio: 1.0 target: <|long|> <|tag_to_long|> tag: scenery, no humans, storefront, motor vehicle, road sign, power lines, plant, railing, flower pot, vanishing point, outdoors, sign, potted plant, sidewalk, awning, tree, bicycle, window, railroad crossing, bush, building, utility pole, lamppost, shop, truck, traffic light, fence, chinese text, stairs, door, bicycle basket, town, day, streetcar (cafe), lamp, road long: A small town with a variety of buildings and houses. the sky is blue and there are trees in the background. on the left side of the image, there is an orange building with a sign that reads "chinese restaurant". on the right side, there are several other buildings with different types of shops and restaurants. in front of the buildings, there appears to be a street with cars parked along the road. in the center of the illustration, we can see a train crossing signal with two red lights and a blue sky above it. there is also a yellow building with white walls and a green roof. on top of the traffic light pole, there seems to be an air conditioning unit. the street is lined with trees and bushes, and there is graffiti on the ground.
Formatted Output For Text-to-Image
scenery, no humans, storefront, motor vehicle, road sign, power lines, plant, railing, flower pot, vanishing point, outdoors, sign, potted plant, sidewalk, awning, tree, bicycle, window, railroad crossing, bush, building, utility pole, lamppost, shop, truck, traffic light, fence, chinese text, stairs, door, bicycle basket, town, day, streetcar \(cafe\), lamp, road, A small town with a variety of buildings and houses. the sky is blue and there are trees in the background. on the left side of the image, there is an orange building with a sign that reads "chinese restaurant". on the right side, there are several other buildings with different types of shops and restaurants. in front of the buildings, there appears to be a street with cars parked along the road. in the center of the illustration, we can see a train crossing signal with two red lights and a blue sky above it. there is also a yellow building with white walls and a green roof. masterpiece, newest, absurdres, safe
Refer to caption User InputRefer to caption TIPO Output
Figure 10: An example formatted content we used for training and inference in TIPO.
Appendix E Image Examples
In this section, we present sample images from the experiments described in Section 6 to visually demonstrate the improvements achieved by TIPO.
E.1 Scenery Tag Test
(a) Scenery Tag Only
(b) TIPO + Scenery Tag
Figure 11: Comparison of generated images using simple input (left) vs. TIPO-enhanced input (right) for the scenery tag
Figure 11 demonstrates the difference in output diversity between simple input and TIPO-enhanced input for the scenery tag. As observed, TIPO significantly expands the range of generated sceneries, better reflecting the variety present in the Danbooru2023 dataset [50]. The left column shows results from simple input (scenery tag only), while the right column illustrates the enhanced diversity achieved with TIPO-enhanced input.
E.2 Short/Truncated Long Test
(a) Short Caption
(b) TIPO-Generated Caption
(c) Truncated Long Caption
(d) TIPO-Extended Caption
Figure 12: Comparison of generated images using original input (left) vs. TIPO-enhanced input (right)
Figure 12 illustrates the differences between short captions, truncated long captions, TIPO-generated captions, and TIPO-extended captions. The “short prompt” and “truncated long prompt” used in this experiment typically consist of 1-2 sentences, resulting in reasonably good quality outputs. However, the use of TIPO to refine or extend these prompts still yields noticeable improvements in aesthetics and overall quality.
Appendix F Topic Distribution Visualization
To analyze the meaningfulness and diversity of topics within text inference results, we implemented a multi-stage topic modeling and clustering methodology. Initially, we employed Latent Dirichlet Allocation (LDA) to extract varying numbers of topics—specifically 20, 30, 50, and 100—from the corpus [8]. This approach aimed to identify representative words that are significant within the dataset while filtering out common stop words, ensuring that nonsensical or irrelevant terms are excluded from topic classification.
By employing multi-level topic analysis, we empirically assessed whether the resulting topics were sufficiently large and diverse. This iterative process mitigates potential challenges such as substantial topic overlap, which can diminish distinctiveness when extracting a large number of topics (Stevens et al., 2012).
To address the potential overlap and further assess the diversity and meaningfulness of the topics, we performed a secondary clustering. We grouped the initially extracted topics into five major clusters using k-means clustering and evaluated the clustering performance by calculating the inertia (Sum of Squared Distances) [23], shown in Table 4, 5, and 6. Since the topics have already been filtered for meaningful content, a higher inertia value indicates greater diversity among the clusters, reflecting a broader range of valid and meaningful topics across the dataset. This two-tiered approach allows for a more nuanced analysis of topic diversity and ensures the robustness of the topic modeling against meaningless word groupings.
Table 4: Inertia for COYO-Dataset inference, higher is better
Size Promptdb GPT4o-mini Promptist TIPO
Run 1 Run 2 Run 1 Run 2 Run 1 Run 2 Run 1 Run 2
20 170.78 137.94 1078.38 204.71 125.90 144.59 1037.44 271.77
30 461.79 417.74 1758.40 736.74 327.48 195.13 1323.88 512.07
50 829.68 730.73 861.15 1036.33 400.90 373.74 823.51 959.18
100 1656.74 1245.60 1987.63 1628.32 877.36 657.14 1622.79 1777.61
Table 5: Inertia for GBC-Dataset inference, higher is better
Size Promptdb GPT4o-mini Promptist TIPO
Run 1 Run 2 Run 1 Run 2 Run 1 Run 2 Run 1 Run 2
20 184.53 352.16 244.79 246.15 125.47 198.94 278.56 211.98
30 452.01 566.74 505.56 441.34 204.28 328.70 372.07 471.28
50 571.77 895.47 1227.30 990.17 438.89 313.48 737.65 788.41
100 1291.60 1742.36 1675.41 1550.32 631.61 628.78 1573.47 1855.90
Table 6: Inertia for Scenery extend inference, higher is better
Size PromptDB GPT4o-mini Promptist TIPO
20 60.82 734.52 210.60 139.29
30 275.76 1141.77 415.95 355.20
50 630.50 826.29 722.75 1002.36
100 2026.39 1879.08 802.93 1883.70
We attach simple visualization of topics in scenery prompt generation, with topic n=100, cluster k=5.
Figure 13: Topic visualization for scenery prompt generation. A wider spread indicates a greater diversity of generated topics.

### v2

@@
 The need for intricate, nuanced prompts often results in a higher barrier to entry, limiting accessibility and ease of use for those unfamiliar with the complexities of prompt engineering.
-The previous works [31] tried to utilize LLMs to lenghthen prompt to solve tendency, however many ignore the root cause of why some prompts result in worse images and rely heavily on manually engineered prompts, which require substantial resources.
+Previous works have attempted to extend prompts with optimization through various methods.
+Some approaches use LLMs to expand prompts [41]; however, due to the misalignment between the LLM’s text-rich training corpus and the T2I model’s caption-focused corpus, only a limited set of manually-engineered prompts yield effective results.
+Other methods rely on curated prompt sets [5, 17], which, while potentially incorporating user preferences, fail to fully leverage the T2I model’s trained prompt distribution.
+Additionally, reinforcement learning-based techniques [26] require multi-turn inference, incurring high computational costs, and often struggle to generate diverse outputs due to the model’s inherent corpus limitations.
 To address these challenges, we introduce TIPO (Text to Image with text pre-sampling for Prompt Optimization), an innovative framework designed to enhance T2I generative models.
 TIPO introduces ‘text pre-sampling’ preprocess procedure in inference pipeline, letting text-to-image models reproduce a more accurate distribution of corresponding user prompts without additional manual engineering.
-Unlike previous approaches that focus on extending prompts with manually collected prompt sets [4, 15, 45, 21], directly use LLM to modify the prompts [26, 51] or utilize reinforcement learning like method which update model’s input or model itself from text-to-image feedback [22, 13, 31] without addressing underlying issues, TIPO introduces a method that optimizes prompts efficiently as shown Section 4.
 Our work makes the following key contributions:
@@
 Figure 1: Comparison of prompt optimization methods.
-(a) uses instructions for prompt refinement but lacks understanding of T2I-specific prompts.
-(b) improves detail but limits variety.
-(c) optimizes for model feedback but are computationally expensive.
-(d) TIPO (ours) aligns prompts with the trained distribution of T2I models, ensuring prompt detail, consistency with training data, and controlled variety within model scope.
-This paper presents TIPO’s theoretical foundations, implementation details, and experimental results demonstrating its effectiveness in improving aesthetic scores [17], reducing image corruption [32, 33], and better aligning generated images with dataset distributions [24, 35].
+(a) uses instructions for prompting but its understanding is constrained by the LLM’s knowledge base, not T2I model.
+(b) relies on a curated prompt database, enhancing detail but limiting variety by not fully leveraging the T2I model’s learned distribution.
+(c) optimizes using the scorer with RL, requiring multi-turn inference with additional cost.
+(d) aligns prompts with the T2I model’s training distribution, ensuring detailed and various topic-related prompt generation that fits the target T2I model.
+This paper presents TIPO’s theoretical foundations (Section 3), implementation details (Section 4), and experimental results (Section 5) demonstrating its effectiveness in improving aesthetic scores [19], reducing image corruption [43, 44], and better aligning generated images with dataset distributions [28, 46].
 Our findings highlight the critical role of prompt optimization in advancing T2I technology and open avenues for broader applications of automatic prompt refinement.
@@
 By contrast, TIPO offers an efficient approach to prompt optimization by directly leveraging the training dataset, removing the need for manual prompt engineering and runtime tuning, and achieving versatile, contextually aligned image generation.
-Figure 2: Comparison of prompt optimization methods for a T2I model [7].
-(a) yields a basic image with limited detail.
-(b) provides specific output but requires manual engineering.
-(c) introduce nonsensical elements (red boxes) in the teal region, where the model’s actual output distribution for a given prompt (green) and all the image distribution (blue) overlap outside the user’s expected output distribution (red).
-(d) captures user intent accurately, balancing detail and diversity.
+[…]
 3 Prompt Optimization
@@
 This manual prompt expansion can be time-consuming and labor-intensive.
-Moreover, simply creating a longer prompt does not guarantee improved results; it must be crafted in line with the model’s specific prompting conventions and interpretive nuances.
 To address these challenges, previous works have aimed to automate the prompt optimization process.
-Some methods attempt to expand prompts using the model’s familiar vocabulary and stylistic tendencies [6], while others rely on direct feedback from the model to iteratively improve prompt alignment [22].
+Some methods attempt to expand prompts using the familiar vocabulary and stylistic tendencies of the model [25], while others rely on direct feedback from the model to iteratively improve prompt alignment [26].
 However, these approaches have limitations: if the structure or language does not fully align with the model’s prompting conventions, performance may degrade, and feedback-based approaches often involve significant computational costs.
-Therefore, effective prompt optimization for text-conditioned models requires accurately capturing user intent and expanding the prompt in a way that is both precise and aligned with the model’s conventions for interpreting text inputs.
+Therefore, effective prompt optimization for text-conditioned models requires accurately capturing user intent and expanding the prompt in a way that is both precise and aligned with the model’s conventions for interpreting text input.
 This need for precision and alignment motivates a systematic approach to prompt formulation, which we define in the following section.
@@
 The simplest approach to extend a simple prompt is by adding random words.
-However, this approach has limitations, as it often introduces internally conflicting information that results in out-of-distribution (OOD) prompt inputs, leading the model to generate nonsensical images.
-In such cases, although each added word may be meaningful in isolation, these random combinations can create prompts that are internally contradictory (and are not simply creative combinations).
-Thus, while random word extension typically yields unreliable results, it may occasionally produce a usable prompt by chance.
+However, this approach has limitations, as it often introduces internally conflicting information that results in out-of-distribution (OOD) prompt inputs, leading the model to generate nonsensical images or the model to generate not user intent images.
+In such cases, although each added word may be meaningful in isolation, these random combinations can create prompts that are internally contradictory, as shown in Figure 2 (c).
+For example, the generated images may deviate from the user’s intent by not depicting an astronaut (left top), producing a horse-shaped tank (right top) instead of a horse, omitting Mars (left bottom), or even including a random snake (right bottom), which introduces nonsensical elements.
 To represent this random extension method, let $p_{r_{i}}$ denote the $i^{th}$ prompt created by adding random words to $p_{s}$.
@@
 $\bigcup_{i=0}^{n}f(p_{r_{i}})(\mathcal{N})\not\approx\mathcal{I}_{p_{s}}$
-This scenario is illustrated in Figure 2, where “Simple Prompt + Random Words” shows a scattered distribution extending beyond the bounds of the desired distribution or even the “image” distribution itself.
-4 TIPO Framework
-4.1 Overview
-TIPO introduces a dataset-driven approach to prompt optimization by training directly on the existing dataset.
-This avoids reliance on manually engineered prompts [4, 45, 15] or on computationally intensive RL techniques [22, 31].
-Methods using LLMs rely on general language knowledge, which may misalign with specific T2I model needs [6, 51].
-In contrast to previous approaches, TIPO captures the model’s training data distribution closely.
+This scenario is illustrated in Figure 2 (c), where “Simple Prompt + Random Words” shows a pointy distribution extending beyond the bounds of the desired distribution or even the “image” distribution itself.
+4 TIPO Framework and Model
 4.1 Distribution-Aligned Prompt Extension
@@
 where $p_{t_{i}}$ denotes the $i^{th}$ TIPO-generated prompt derived from $p_{s}$, and $\mathcal{I}_{p_{s}}$ represents the target distribution of image outputs associated with $p_{s}$.
-Through this systematic prompt refinement, Figure 2 “Simple Prompt + TIPO” illustrates the broader image space exploration enabled by TIPO while retaining relevance to the original prompt, effectively balancing diversity with coherence.
+Through this systematic prompt refinement, Figure 2 (d) “Simple Prompt + TIPO” illustrates the broader image space exploration enabled by TIPO while retaining relevance to the original prompt, effectively balancing diversity with coherence.
 4.2 Prompt Structure
@@
 Importantly, in certain cases, input tag or NL prompts can be treated as metadata <Content> as well, when the task involves generating tags or NL prompts conditioned on each other.
+4.4 Task Definitions
+Figure 3: TIPO’s task flow within a single generation cycle.
+Starting from any node, each arrow represents a sequential extension step, with other prompt used as metadata if provided.
+This task design enables efficient and flexible prompt extension across multiple tasks.
+TIPO defines three core tasks: extending tag sequences, extending NL prompts, and generating refined NL prompts.
+Each task is represented by an arrow in Figure 3, illustrating how TIPO can use tag and NL prompts interchangeably as input sources to generate detailed outputs within a single generation cycle.
+When an arrow cycle is established, TIPO can complete multiple extension within a single generation pass.
+For example, TIPO can start from a short NL prompt to produce a detailed tag sequence, or it can combine tasks like short_to_tag and tag_to_long to create a composite task, short_to_tag_to_long, which generates an extended NL prompt from an initial short NL input via tag generation.
+Further details on specific task sequences are provided in Appendix.
 4.5 An Example Scenario and Workflow
-Figure 3: (a) A generated image and prompts.
-(b) Prompt optimization of TIPO, starting from user input (short tag prompt $p_{s}$) to final output (detail prompt $p_{d}$).
+Figure 4: An example scenario of TIPO workflow.
+(a) A generated image and prompts.
+(b) Prompt optimization of TIPO, starting from simple user input $p_{s}$ to detailed final output $p_{d}$.
 The shading from gray to light sky blue represents an increase in context richness in the prompt.
@@
 Figure 4 provides an example of this setup, where:
-$\begin{array}[]{ll}S_{s}&=\{\textsf{A young girl with long hair...}\},\quad\verb|<meta>|=\emptyset{}\\ T_{s}&=\left\{\textsf{outdoors, scenery, water, wind, landscape, ...}\right\}\\ \end{array}$
+$\begin{array}[]{ll}S_{s}&=\{\textsf{A young girl with long hair...}\},\quad\verb|<meta>|=\emptyset{}\\ T_{s}&=\left\{\textsf{outdoors, scenery, water, wind, landscape, ...}\right\}\end{array}$
 The generation proceeds sequentially as follows:
@@
   3. 3.
-    short_tag_to_long: With both $T_{d}$ and $S_{s}$ in the metadata, TIPO generates a comprehensive long NL prompt $S_{d}$, ensuring a more detailed prompt.
+    short_to_tag_to_long: With both $T_{d}$ and $S_{s}$ in the metadata, TIPO generates a comprehensive long NL prompt $S_{d}$, ensuring a more detailed prompt.
   4. 4.
     TIPO aggregates $T_{d}$, $S_{d}$, and any additional metadata to construct a context rich prompt, $p_{d}$.
-This example highlights how TIPO leverages an autoregressive LM to perform prompt extension in a stepwise, sequential processes, through short_to_tag, tag_to_long, and short_to_tag_to_long tasks within a single generation cycle.
-More detailed task definitions for TIPO can be found in Section 5.2.
-By iteratively refining prompts in this structured way, TIPO addresses the limitations of simplistic or randomly extended prompts, enabling it to generate high-quality, diverse images that closely align with user intent.
-4.6 Versatile and Scalable Design of TIPO
-The example in Section 4.5 demonstrates how TIPO’s structured approach enables it to create coherent, intent-aligned prompts.
-Building on this foundation, TIPO’s design extends smoothly to a wide range of models and datasets, showcasing its versatility and scalability.
-A single TIPO model trained on dataset D can be applied to all text-to-image models trained on the same dataset, enabling broad applicability across different architectures and versions that share the same training data.
-Similarly, a TIPO model trained using caption method M (e.g.
-tag, NL captions) can be used with any T2I model trained with captions from method M.
-Experimental results in Section 6 further illustrate this adaptability, as TIPO successfully applies across multiple model architectures and training configurations, confirming its broad applicability and robustness.
-5 Training TIPO model
-In this section, we detail how to build the TIPO model based on the TIPO framework.
-For various T2I models [36, 6, 19, 41, 38, 39, 43, 40, 37, 42, 12, 11, 28, 20, 7], tag and NL captions are widely used captioning methods.
-To allow TIPO model to seamlessly transform captions between different formats or refine captions within the same format, we define specific tasks that enable flexible and precise prompt construction.
-5.1 Modeling
-TIPO adopts a causal autoregressive LM framework as prompt extension function, $\mathcal{E}$, training to extend and refine user-provided prompts into detailed, content-rich prompts $p_{d}$ from simpler inputs $p_{s}$, thereby effectively balancing specificity with alignment to user intent.
-This process captures a subset of outputs closely aligned with the user’s original intent, ensuring relevance.
-Through use of existing datasets for prompt optimization, along with its iterative extension function and application of LM, TIPO achieves significant coverage and diversity in generating high-quality images aligned with training distributions, all without the need for manual prompt engineering.
-5.2 Task Definitions
-TIPO encompasses three primary tasks: extending tag sequences, extending NL prompts, and generating refined NL prompts.
-To facilitate these tasks, we define several specific task types and corresponding formatting methods during training:
-  1. 1.
-    tag_to_long: Use tags as metadata to generate a new NL prompt or extend a user-provided NL prompt.
-  2. 2.
-    long_to_tag: Use an NL prompt as metadata to extend a tag sequence.
-  3. 3.
-    short_to_tag: Use the simple prompt $p_{s}$ as metadata to extend a tag sequence.
-  4. 4.
-    short_to_long: Use a user-input NL prompt as metadata to generate a refined detailed prompt $p_{d}$.
-  5. 5.
-    short_to_tag_to_long: Use a user-input NL prompt or tag sequence as metadata to generate a refined detailed prompt $p_{d}$.
-  6. 6.
-    short_to_long_to_tag: Use a user-input NL prompt or generated NL prompt as metadata to extend a tag sequence.
-  7. 7.
-    tag_to_short_to_long: Use user-input tags or NL prompts as metadata to generate a refined detailed prompt $p_{d}$.
-By defining these task types, we can achieve special cases where, for instance, inputting a short description or a short tag sequence results in a full-size tag sequence and NL prompt in a single pass (i.e., generating $T_{d}$ and $p_{d}$ simultaneously) as shown in Section 4.5.
-5.3 Model and Training Details
-We utilize the LLaMA architecture [46, 47] with models of 200 million and 500 million parameters:
-  • TIPO-200M: Pretrained on the Danbooru2023 [34, 49] and GBC10M [25] datasets for 5 epochs, then fine-tuned on the Danbooru2023, GBC10M, and CoyoHD11M [9] datasets for 3 epochs, resulting in a total of approximately 40 billion non-padding tokens seen.
-  • TIPO-500M: Pretrained on the Danbooru2023, GBC10M, and CoyoHD11M datasets for 5 epochs, resulting in a total of approximately 30 billion non-padding tokens seen.
-Note that we calculate the token count based on non-padding tokens only. Given the relatively short and variable lengths of data entries, the effective token count may be lower than anticipated.
-5.4 Training Procedure
-In our experiments, each training pass generates a new prompt of a single type.
-To cover various prompt types—such as a refined prompt $p_{d}$, an extended NL prompt $S_{d}$, and an extended tag sequence $T_{d}$—the model requires at least three passes.
-For each dataset entry, we randomly select one of the seven task types to apply and randomly decide how to split the simple prompt from the detailed prompt ($p_{s}$ from $p_{d}$) and the simple tag set from the detailed tag set ($T_{s}$ from $T_{d}$).
-This randomization in task selection and prompt splitting effectively increases the training dataset’s size, allowing the model to encounter a broader range of inputs and outputs.
-This approach enhances the model’s ability to generalize across the various tasks defined in TIPO.
+This iterative process allows TIPO to build prompts that are both detailed and contextually aligned with user inputs, through short_to_tag, tag_to_long, and short_to_tag_to_long tasks within a single generation cycle.
+4.6 Model Architecture
+TIPO utilizes a causal autoregressive language model framework as its prompt extension function, $\mathcal{E}$, to iteratively extend and refine user-provided prompts into more detailed, content-rich prompts $p_{d}$ based on simpler inputs $p_{s}$.
+Our implementation leverages the LLaMA architecture [61, 62, 3], employing models with 200M and 500M parameters.
+  • TIPO-200M: Pretrained on the Danbooru2023 [45, 65] and GBC10M [29] datasets for 5 epochs, then fine-tuned with Danbooru2023, GBC10M, and CoyoHD11M [12] for 3 epochs, covering approximately 40 billion tokens.
+  • TIPO-500M: Pretrained on the Danbooru2023, GBC10M, and CoyoHD11M datasets for 5 epochs, covering approximately 30 billion tokens.
+Throughout training, non-padding tokens were used for token count calculations, with the effective count reflecting the relatively short and variable length of data entries.
+4.7 Training Process
+In each training pass, TIPO applies dataset augmentation by selecting tasks from all possible task combinations such as short_to_tag_to_long, including extending tag sequences, NL prompt extension, and refined NL prompt generation.
+This random selection across all task combinations significantly broadens the dataset’s effective scope, enhancing model generalization without additional manual engineering or runtime tuning.
 (a) Direct LLM
@@
 (d) TIPO (Ours)
-Figure 4: Scenery test examples from 4 different methods.
+Figure 5: In-domain test regard to scenery test examples from 4 different methods.
+(a) Scenery Tag Only
+(b) Scenery Tag + TIPO
+(c) Truncated Long Prompt
+(d) Truncated Prompt + TIPO
+Figure 6: Comparison of generated images: (a) simple scenery tag, (b) scenery tag enhanced by TIPO, (c) truncated long prompt, and (d) TIPO-enhanced truncated prompt.
+TIPO adds detail and maintains variety, yielding coherent images even from minimal prompts.
 5 Evaluation
-In this section, we detail the experimental setup used to evaluate the TIPO framework.
-We outline the formatting conventions for prompts and metadata, define the various tasks employed during training, and describe the models and datasets utilized in our experiments.
 5.1 Experimental Setup
+In this section, we detail the experimental setup used to evaluate prompt optimization framework.
 To assess the quality of generated prompts and resulting images, we employed the following evaluation metrics.
-FDD (Frechet Dino Distance): Traditionally, the Frechet Inception Distance (FID) [24] has been used to measure the distribution distance between datasets and generated images.
+FDD (Frechet Dino Distance).
+Traditionally, the Frechet Inception Distance (FID) [28] has been used to measure the distribution distance between datasets and generated images.
 However, recent works have shown that FID is not always aligned with human preferences [58].
@@
 In our experiments, we used the pretrained ViT-g [20] model from DinoV2 and used the normalized embedding vector to measure the Frechet Distance result.
-Aesthetic Score: We compute the Aesthetic Score using the Aesthetic Predictor V2.5 [17].
+Aesthetic Score.
+We compute the Aesthetic Score using the Aesthetic Predictor V2.5 [19].
 Higher aesthetic scores indicate images that are perceived as more visually appealing, well-composed, and of higher quality.
 This metric aims to quantify the subjective visual appeal of generated images, with higher scores suggesting outputs that are more likely to be judged favorably by human viewers.
-AI Corrupt Score: The AI Corrupt Score is obtained from the AICorruptMetrics [32] in sdeval [33].
+AI Corrupt Score.
+The AI Corrupt Score is obtained from the AICorruptMetrics [43] in sdeval [44].
 This metric is trained on AI-generated images with human-annotated “corrupt or not” labels.
 A higher value indicates a higher likelihood of the image being “correct” or “complete”.
-All results presented in this section are tested on the TIPO-200M model.
-6.2 Scenery Tag Test
-In this test, we randomly select 32,768 entries from Danbooru2023 [50, 34] that include the “scenery” tag. We set up the following inputs:
-  • “Scenery” + Meta:
-    • Retain all metadata categories as described earlier.
-    • Only include “scenery” as the content tag.
-  • “Scenery” + Meta + TIPO:
-    • Use “scenery” + meta as input.
-    • Extend $T_{s}$ (scenery) to $T_{d}$ and generate $p_{d}$ from $T_{d}$.
-For the scenery tag test, we generate one image from each prompt using the Kohaku-XL-zeta [^2] https://huggingface.co/KBlueLeaf/Kohaku-XL-Zeta model, which is a fine-tuned SDXL [37] model on the Danbooru dataset [34].
-Figure 4 illustrates the comparative visual result by various methods.
-The results reveal significant biases in Prompt DB-based [21] and RL-based [22] approaches.
-The former consistently generates snowy mountain scenes, while the latter produces near-future fantasy scenery, often resulting in corrupted images.
-The Direct LLM method, utilizing GPT-4o [2], shows improved results but still exhibits notable bias; out of 9 generated prompts, 8 exclusively describe natural scenes only without any human-made structures.
-In contrast, TIPO demonstrates superior performance in reproducing the dataset’s distribution.
-While primarily generating nature scenes, it also incorporates architectural elements and artifacts, more accurately reflecting the diverse content of the original dataset.
-This balanced output suggests that TIPO effectively mitigates the biases observed in other methods.
+Vendi Score.
+To quantify the diversity of generated images, we employ the Vendi Score metric based on DinoV2 [46] embeddings.
+The computation involves calculating a similarity matrix $K$ using cosine similarity between image embeddings, followed by computing the von Neumann entropy of the normalized matrix $K/n$, adhering to the standard Vendi Score methodology.
+This metric provides a quantitative measure of the distributional diversity within a set of generated images.
+Higher Vendi Scores indicate greater inter-image diversity, reflecting the model’s capability to generate varied and distinct outputs.
+Extension Functions.
+We used both GPT-4o-mini [2] and LLaMA-3.2-1B [3] as Direct LLM models for prompt generation, balancing performance and efficiency.
+GPT-4o-mini, as a lightweight closed-source model, offers speed and affordability, while LLaMA-3.2-1B represents a similar scale to TIPO in open-source models.
+Also, Prompt DB [17] and Promptist [26] serve as baseline methods for prompt extension.
+5.2 In-domain test regard to scenery tag
+To evaluate the model’s capability to generate diverse and accurate outputs, we chose the scenery tag, which includes both natural landscapes and human-made structures.
+In this test, we randomly select 32,768 entries from Danbooru2023 [65, 45] that include the “scenery” tag.
+We set up the following prompt:
+<meta> + scenery + $\mathcal{E}$>
+Then, we apply extension functions to extend only the scenery, $T_{s}$, as the <Content> tag.
+We generate one image from each prompt using the Kohaku-XL-zeta [^2] https://huggingface.co/KBlueLeaf/Kohaku-XL-Zeta model, which is a fine-tuned SDXL [48] model on the Danbooru dataset [45].
 Metric Original LLM Prompt DB Promptist TIPO (Ours)
@@
 AI Corrupt Score $\uparrow$ 0.4257 0.7490 0.5024 0.5669 0.9195
-Table 1: Metric results on Scenery Tag Test.
-As shown in Table 1, applying TIPO significantly improves FDD performance and AI Corrupt score.
-The improvement in FDD score suggests that TIPO-generated prompts lead to images that more closely align with the original dataset distribution.
-Furthermore, the enhanced AI Corrupt score indicates that TIPO aids in generating more coherent images, potentially reducing corruption or distortion in the output.
-The results from the direct LLM approach (GPT4o-mini) exemplify a key concern in prompt engineering: while achieving high aesthetic scores, it performs poorly in terms of FDD.
-This discrepancy underscores the potential pitfall of using LLMs for prompt generation without considering dataset alignment.
-Such models may produce well-constructed prompts that yield visually appealing images but fail to capture the nuanced distribution of the text-to-image dataset.
-In contrast, TIPO demonstrates a balance between aesthetic quality and dataset fidelity, offering a more robust solution for prompt optimization in text-to-image generation tasks.
-6.3 Short/Truncated Long Test
-This test evaluates TIPO’s performance on enhancing natural language prompts which are already contains some useful info for model to generate reasonable images.
+Table 1: Metric results on in-domain test.
+Bold values indicate the best performance, and underlined values indicate the second-best.
+Original refers to images generated using the original text captions; LLM uses GPT-4o-mini [2]; the remaining columns represent results from Prompt DB [25], Promptist [26], and TIPO, respectively.
+As shown in Table 1, TIPO-generated prompts improve FDD scores by aligning images more closely with the original dataset distribution.
+Enhanced AI Corrupt scores show that TIPO reduces image corruption, creating more coherent outputs by replacing vague descriptions (e.g., detailed background) with specific elements (e.g., garden filled with forget-me-nots and blue butterflies).
+The direct LLM approach highlights a key issue in prompt engineering: it achieves high aesthetic scores but compromises FDD, reducing the diversity of images.
+We provide more comparative visual result in Figure 5.
+The results reveal significant biases in Prompt DB  [25] and RL-based [26] approaches.
+PromptDB and Direct-LLM method exhibits notable bias, by generating mountain/lake scene majorly.
+Meanwhile, Promptist produces diverse near-future fantasy sceneries, however often produces corrupted images due to unreasonable vocabulary usage (4K, masterpiece, instagram), supposedly introduced by user prompts.
+By contrast, TIPO consistently reproduces a balanced output distribution, incorporating both natural and architectural elements, effectively mitigating the biases observed in other methods.
+(a) Original prompt best aesthetic
+(b) GPT-4o-mini best aesthetic
+(c) TIPO best aesthetic
+Figure 7: Top-10% aesthetic score images generated by the SD-3.5-Large model [19].
+Rows show samples from Original, Direct LLM-Modified prompts, and TIPO-extended prompts.
+Direct LLM images focus primarily on natural scenery, while generated image from TIPO-extended prompt display a wider variety, including urban scenes and human figures.
+5.3 In-domain prompt generation test
+We aim to test each models’ performance on regard to reconstructing / upsampling the given prompt, in given dataset distribution.
+By comparing original ground truth and generated outputs, we can directly compare FDD, aesthetic score, AI corrupt scores.
 We randomly selected 10k entries each from [12] and [29].
 For each entry, we extracted the short caption and the truncated long caption (limited to a maximum of two sentences).
-These captions serve as user input prompts for our test.
-Using these prompts, we generated one image per prompt with the SDXL-1.0-base [^3] https://huggingface.co/stabilityai/stable-diffusion-xl-base-1.0 model [37].
+These captions serve as user input prompts $p_{s}$ for in-domain prompt generation test.
+Using these prompts, we generated single image per prompt with the SDXL-1.0-base [^3] https://huggingface.co/stabilityai/stable-diffusion-xl-base-1.0 model [48].
 Short Original LLM Prompt DB Promptist TIPO(Ours)
@@
 AI Corrupt Score $\uparrow$ 0.6868 0.6712 0.6741 0.5925 0.7130
-Table 2: Metric result in Short/Truncated-long test of prompt generation
-Table 2 presents the results of our Short/Truncated-long test for prompt generation.
-Unlike the scenery tag test (Section 6.2), we observe that the original prompts in this test already produce decent images.
-Interestingly, all four methods result in worse FDD scores compared to the original prompts, suggesting that the prompt generators introduced new elements not present in the selected 10,000 entries.
-Given the size of the training set, this outcome is not entirely unexpected.
-Examining the Aesthetic and AI Corrupt Scores, we notice that the Direct LLM approach achieves the highest aesthetic score, mirroring the results from the scenery tag test.
-However, its AI Corrupt Score and FDD value indicate potential issues with OOD inputs, which may occasionally hinder the model’s ability to generate coherent images from its prompts.
-TIPO, on the other hand, demonstrates a more balanced performance.
-It achieves the best AI Corrupt Score, the second-best Aesthetic Score, and a reasonable FDD value.
-This balanced profile suggests that TIPO is the most suitable choice for real-world applications of modern text-to-image models, where a combination of aesthetic appeal, image coherence, and distribution alignment is crucial.
+Table 2: Performance comparison in the in-domain prompt generation test.
+Bold values indicate the best performance, and underlined values indicate the second-best.
+Direct LLM achieves the highest Aesthetic Score but struggles with OOD issues.
+TIPO provides a balanced outcome, excelling in AI Corrupt Score and maintaining strong alignment, making it ideal for practical applications.
+Table 2 presents the results of our Short/Truncated caption test for prompt generation.
+Compared to in-domain test, Section 5.2, which might be considered slightly out of domain by SDXL model, we observed that prompts covered by this test produces meaningful, coherent images with sufficient detail level.
+FDD scores were degraded compared to original-ground truth prompts, not truncated long caption-which implies prompt extension functions resulted lack of diversity, or unmatched components in generation.
+We observe that Direct LLM approach achieves the highest aesthetic scores by wide vocabulary with user preference matching, however results low FDD, AI corruption score, which implies unmatched vocabularies are being used in-prompt.
+Original LLM Prompt DB Promptist TIPO(Ours)
+Top 10% $\uparrow$ 10.98 7.84 12.17 10.12 17.33
+Worst 10% $\uparrow$ 12.10 9.76 15.23 14.14 13.92
+Table 3: Vendi Scores [23] for the top and worst 10% aesthetic images across prompt generation methods.
+Higher value indicate higher diversity of samples.
+TIPO achieves high score in the top 10% (17.33), showing greater visual diversity as well as worst 10%.
+Result suggests TIPO’s balanced and robust variety in generated images.
+Meanwhile, LLM has the lowest scores in both categories, indicating limited diversity and repetition.
+5.4 Out of focus evaluation.
+In this section, we analyze TIPO’s behavior in out-of-focus model, SD-3.5-Large [^4] https://huggingface.co/stabilityai/stable-diffusion-3.5-large built upon MMDiT architecture [21].
+We compose original tag, natural language prompt created by GPT-4o-mini, tag+natural language extended prompt by TIPO (ours), then feed to SD-3.5-Large inference pipeline.
+Analysis of Vendi Scores [23] (Table 3) reveals contrasting patterns: Direct LLM-generated images show limited diversity despite high aesthetic scores, while TIPO maintains conceptual diversity across all quality levels.
+However, TIPO’s training on diverse datasets [45, 65, 12, 29] introduces domain-specific elements that may not align with SD3.5L’s training distribution.
+This mismatch demonstrates our core finding: prompt generation models perform optimally when trained on domain-matched prompt corpora.
+The balanced profile of TIPO suggests that TIPO can work as suitable solution for prompt upsampling task, which can accurately project trained prompt with few initial word input.
 6 Conclusion
 In this paper, we introduced TIPO (Text to Image with text Presampling for Prompt Optimization), a novel framework designed to enhance the quality of Text-to-Image (T2I) models through automatic prompt engineering.
-By leveraging Large Language Models (LLMs) to extend and refine user-provided prompts, TIPO effectively bridges the gap between simple user inputs and the detailed, content-rich prompts required for high-quality image generation.
-Our comprehensive evaluation demonstrates TIPO’s effectiveness across multiple dimensions of image quality and fidelity.
-The framework consistently produced prompts that led to images closely aligned with original dataset distributions, while simultaneously improving aesthetic appeal and reducing the likelihood of corrupted or nonsensical outputs.
-This balanced performance underscores TIPO’s robustness and suitability for real-world applications of T2I models, where a combination of distribution alignment, visual appeal, and image coherence is crucial.
-Notably, our experiments highlight the critical role of prompt engineering in advancing T2I technology.
-The improvements achieved through TIPO’s prompt modifications often rivaled or surpassed performance differences between various T2I model architectures.
-This finding emphasizes that effective prompt engineering can significantly enhance model performance without necessitating architectural changes, potentially offering a more efficient path to improving T2I systems.
-The success of TIPO opens several promising avenues for future research:
-  1. 1.
-    Broader Application: Exploring TIPO’s applicability to other generative tasks, such as text generation or audio synthesis, could further demonstrate the versatility and impact of automatic prompt engineering.
-  2. 2.
-    Interactive Systems: Incorporating TIPO into interactive applications where users can iteratively refine prompts in real-time may enhance user experience and enable more precise control over generated content.
-  3. 3.
-    Advanced Alignment Techniques: Investigating more sophisticated methods for aligning user inputs with dataset distributions could further enhance the performance and reliability of generative models.
-  4. 4.
-    Cross-Modal Applications: Adapting TIPO’s principles to facilitate cross-modal generation tasks, such as text-to-video or image-to-text, could expand its utility in multimodal AI systems.
-In conclusion, TIPO represents a significant advancement in the field of text-to-image generation, offering a versatile and efficient solution for prompt optimization.
-By automating the refinement and extension of user prompts, TIPO not only improves the quality and specificity of generated images but also highlights the importance of aligning user inputs with underlying data distributions.
-As generative models continue to evolve, frameworks like TIPO will be instrumental in unlocking their full potential, ensuring they meet diverse user needs with precision and creativity while pushing the boundaries of what’s possible in AI-generated content.
+TIPO represents a significant advancement in the field of T2I generation, offering a versatile and efficient solution for prompt optimization.
+By automating the refinement and extension of user prompts.
+We also highlights the importance of aligning user inputs with underlying data distributions through the comparisons between different methods for prompt engineering.
+As generative models continue to evolve, frameworks like TIPO will be instrumental in unlocking their full potential, ensuring they meet diverse user needs with precision and creativity while pushing the boundaries of what’s possible in generative models.
 Appendix
+[…]
+[…]
 Appendix A Dataset/Resource
@@
 The Coyo HD 11M dataset [12] consists of 11.4 million high-density, high-definition images, with 22.8 million synthetic captions generated from the Coyo-700M dataset.
-The ”HD” refers to both the high resolution of images (filtered to have at least 512 pixels on the shortest edge) and the high concept density, addressing the issues found in previous alt-text image pair datasets, which often included low-quality or low-concept images.
+The “HD” refers to both the high resolution of images (filtered to have at least 512 pixels on the shortest edge) and the high concept density, addressing the issues found in previous alt-text image pair datasets, which often included low-quality or low-concept images.
 Synthetic captions for this dataset were produced using the LLaVA-Next-8B model [37] built upon LLaMA 3 [3] with post-processing for conciseness and clarity.
@@
 In TIPO, we utilize the short/long captions, booru tags, and open images tags from Coyo HD 11M.
-Appendix B Implementation Detail
-In this appendix, we provide all the missing details about our dataset construction process that are not mentioned in Section 4.4 and 5.2.
-B.1 Training Settings
-TIPO-200M stage1 TIPO-200M stage2 TIPO-500M
+A.4 Stable Diffusion XL
+The Stable Diffusion XL [48] improves upon earlier versions [54] with a significantly larger UNet backbone and dual text encoders, CLIP ViT-L [49] and OpenCLIP ViT-bigG [30], to enhance text conditioning.
+It supports resolutions up to 1024×1024 and is designed for diverse image generation.
+Although explicit dataset details are not provided, figures in the paper confirm that SDXL accepts both natural language (NL) prompts and tags as input.
+A.5 Stable Diffusion 3.5
+The Stable Diffusion 3.5 (SD-3.5) incorporates a cutting-edge architecture, MMDiT [21], along with the Rectified Flow [38, 4, 35] formulation for improved text-to-image generation.
+The model utilizes triple text encoders—CLIP/ViT-L, OpenCLIP/ViT-G, and T5-XXL [50]—for robust text conditioning.
+SD-3.5 supports resolutions up to 1024×1024 and employs a 50/50 mix of original and CogVLM-generated captions for training.
+While explicit dataset details are not provided, figures in the paper confirm its capability to process both natural language prompts and tags.
+In this study, we use the SD-3.5-Large [^8] https://huggingface.co/stabilityai/stable-diffusion-3.5-large (8B parameters) with FP8 inference on RTX 3090 or RTX 4090 GPUs.
+Appendix B Implementation Details
+In this appendix, we provide all the necessary details including our dataset construction process, model configurations, inference pipeline and the model’s property which are not mentioned in Section 4.3 and 4.4.
+B.1 Training Data Construction
+This section details our methodology for constructing and preprocessing training data to ensure robust model performance across various input scenarios.
+Length Control
+To systematically control output prompt length, we implement a structured length categorization system using special length tags.
+These tags enforce specific constraints on both tag counts and natural language sentence lengths.
+For instance, the <long> tag specifies that the corresponding prompt must contain between 36 and 52 tags (inclusive), accompanied by 4 to 8 sentences of natural language description.
+We define four distinct length categories, each with strict bounds for both tag count and sentence length.
+Type Very Short Short Long Very Long
+Tags (count) 18 36 48 72
+NL (sentences) 2 4 8 18
+Table 4: Maximum length specifications for each category and caption type. For each category, the actual count/length must not exceed these values.
+Random Augmentation
+To enhance input diversity and better simulate real-world usage patterns, we implement several data augmentation strategies:
+  • Metadata Tags: For tags representing image metadata (e.g., artist, character, aspect ratio), we employ two randomization techniques:
+    • Random removal of metadata tags
+    • Random repositioning of metadata tags to the end of the prompt, after all content-related descriptions
+    This approach encourages the model to handle varying metadata positions and availability, while maintaining the ability to infer metadata relationships from the content descriptions.
+  • Content Tags: For tags describing image content (e.g., objects, actions, attributes), we implement:
+    • Random shuffling of tag order within the content section
+    • Length-based truncation to meet target length constraints while preserving key content information
+  • Natural Language: For natural language descriptions exceeding length limitations, we employ selective sentence removal, targeting middle sentences to preserve both context-setting opening sentences and concluding details.
+This maintains coherent narrative flow while meeting target length requirements.
+These augmentation strategies create a more diverse training dataset that better reflects real-world prompt variations, improving the model’s robustness and adaptability to different input styles and formats.
+B.2 Training Settings and Model Configurations
+TIPO-100M TIPO-200M stage1 TIPO-200M stage2 TIPO-500M
 Architecture LLaMA
-Type Pretrain Finetune Pretrain
-Vocab Size 32013 - 32013
-Hidden Dim 768 - 1280
-Attention Heads 12 - 20
-MLP Dim 2304 - 3840
-Hidden Layers 20 - 20
-Model Parameters 203M - 508M
-Max Learning Rate 2e-4 5e-5 2e-4
-Optimizer AdamW AdamW AdamW
-betas 0.9, 0.98 0.9, 0.98 0.9, 0.98
-weight decay 0.01 0.01 0.01
-Dataset GBC, Dan Coyo, GBC, Dan Coyo, GBC, Dan
-max context length 512 1024 1024
-global batch size 2048 2048 3584
-Token Seen 22.625B 18.339B 31.274B
-Hardware 4 x RTX3090 4 x RTX3090 8 x H100
-Training Time (wall) 150 hour 270 hour 100 hour
-Table 3: Training settings and details for TIPO-200M and TIPO-500M.
-B.2 Inference Pipeline
+Type Pretrain Pretrain Finetune Pretrain
+Vocab Size 32013
+Hidden Dim 640 768 - 1280
+Attention Heads 10 12 - 20
+MLP Dim 2240 2304 - 3840
+Hidden Layers 10 20 - 20
+Model Parameters 100M 203M - 508M
+Max Learning Rate 5e-4 2e-4 5e-5 2e-4
+Optimizer AdamW
+LR scheduler Cosine Annealing LR
+betas 0.9, 0.98
+weight decay 0.01
+Dataset Coyo, GBC, Dan GBC, Dan Coyo, GBC, Dan Coyo, GBC, Dan
+Epoch 1 5 3 5
+max context length 512 512 1024 1024
+global batch size 1024 2048 2048 3584
+Token Seen 6.0240B 22.625B 18.339B 31.274B
+Hardware 4 x RTX3090 4 x RTX3090 4 x RTX3090 8 x H100
+Training Time (wall) 22.5 hour 150 hour 270 hour 100 hour
+Table 5: Training settings and details for TIPO-100M, 200M, and 500M.
+The datasets include CoyoHD11M (Coyo), GBC10M (GBC), and Danbooru2023 (Dan).
+Stage 2 additionally incorporates Pixtral [1] to generate NL captions from Danboodu2023 dataset.
+Tokenizer and Task Tokens
+TIPO employs a vocabulary derived from LLaMA2 [62] consisting of 32,000 tokens, with additional tokens (13 tokens) specifically designated for task and length control or placeholders.
+This extended vocabulary includes task identifiers and length modifiers to ensure flexibility across different prompt types:
+  • Placeholder Token (1 token):
+        <|empty|>
+  • Task Tokens (8 tokens):
+        <|gen_meta|>, <|tag_to_long|>, <|short_to_tag|>,
+        <|long_to_tag|>, <|short_to_long|>, <|short_to_tag_to_long|>,
+        <|short_to_long_to_tag|>, <|tag_to_short_to_long|>
+  • Length Tokens (4 tokens):
+        <|very_short|>, <|short|>, <|long|>, <|very_long|>
+Optimizer and Learning Schedule
+Training is performed using the AdamW optimizer [39], with a cosine annealing learning rate scheduler [40].
+The optimizer parameters include $\beta_{1}=0.9$, $\beta_{2}=0.98$, and a weight decay of 0.01.
+Maximum learning rates are adjusted per model size, as outlined in Table 5.
+Training Configurations
+TIPO models are trained in multiple stages.
+Table 5 summarizes the configurations for pretraining and fine-tuning TIPO-100M, TIPO-200M, and TIPO-500M.
+Both pretraining and finetuning was conducted on datasets like Danbooru2023 [45], GBC10M [29], and CoyoHD11M [12].
+Augmented Task Representation
+Each dataset entry undergoes random task assignment and splitting to simulate a wide range of input-output mappings, effectively increasing the dataset size.
+For example, a single entry may contribute to tasks like short_to_tag or tag_to_long, with length modifiers dynamically controlling the output verbosity.
+This approach ensures the model can handle diverse tasks while maintaining robust generalization.
+Hardware and Time Requirements
+Training was conducted on NVIDIA RTX3090 GPUs for smaller models and H100 GPUs for TIPO-500M.
+Total wall-clock training times ranged from 22.5 hours for TIPO-100M to 270 hours for fine-tuning TIPO-200M.
+Token Seen and Effective Training
+Non-padding tokens are used to measure the effective token count during training, ensuring efficiency given the short and variable data lengths.
+Table 5 details the total tokens seen per model and training stage, illustrating the comprehensive exposure to diverse data entries.
+[…]
+Figure 8: TIPO inference workflow, with solid arrows denoting the primary generation steps and dashed arrows indicating alternative generation paths within the same cycle.
+¡TOKEN¿— represents special tokens, with all tokens detailed in Section B.2.
+B.3 Inference Pipeline
 The TIPO inference pipeline is designed to handle various input types and scenarios, combining different tasks to achieve refinement or expansion of both tag-based prompts and natural language prompts.
 Figure 8 illustrates this comprehensive workflow.
-Figure 5: The inference workflow in TIPO, demonstrating how different task combinations are used to refine or expand tag-based and natural language prompts.
 Our framework processes tags and natural language inputs separately, allowing for specialized handling of each input type.
@@
 By leveraging different task combinations, TIPO ensures that both tag-based and natural language prompts are optimized, resulting in more detailed and effective inputs for text-to-image models.
+B.4 Model Speed Comparison
+We conducted comprehensive speed tests of our 100M, 200M and 500M parameter models using the inference pipeline described in Section 5.2.
+Each prompt requires two sequential generation steps.
+Our primary metric is average tokens generated per second, which reflects real-world task performance rather than theoretical maximum throughput.
+The evaluation was performed using llama.cpp [24], an efficient C++ implementation that provides optimized support for various hardware accelerators including CUDA, HIP, and Apple Metal.
+TIPO-100M TIPO-200M TIPO-500M
+Hardware Platform tok/sec gen time tok/sec gen time tok/sec gen time
+M1 Max (32 GPU cores) 339.4 0.66 190.0 1.23 119.4 2.02
+RTX 3090 558.5 0.42 341.4 0.69 289.8 0.81
+RTX 4090 742.9 0.29 454.5 0.51 359.7 0.63
+Table 6: Model performance comparison across different hardware platforms.
+Tokens per second (tok/sec) represents the average generation speed, while generation time (gen time) shows the average time required for a complete two-step prompt optimization process.
 Appendix C Evaluation Statistics
 In this appendix, we provide more statistics for the result obtained in Section 5.
-C.1 Scenery Tag Test
+C.1 In-domain test regard to scenery tag
 (a) The box plot for the Aesthetic Score result of scenery tag test.
 (b) The box plot for the AI Corrupt Score result of scenery tag test.
-Figure 6: The box plot for scenery tag test.
-The box plot shown in Figure 6 presents the aesthetic scores and AI corruption scores for results from the scenery tag test described in Section 6.2.
+(c) The KDE plot for the Aesthetic Score result of scenery tag test.
+(d) The KDE plot for the AI Corrupt Score result of scenery tag test.
+Figure 9: The distribution of aesthetic and ai corrupt score for scenery tag test.
+The box plot and Kernel Density Estimation (KDE) plot shown in Figure 9 presents the aesthetic scores and AI corruption scores for results from the scenery tag test described in Section 5.2.
 The analysis clearly demonstrates that TIPO significantly outperforms all other methods, with a substantial margin of improvement.
-C.2 Short/Truncated Long Test
-Figures 7 and 8 display the box plots of aesthetic scores and AI corruption scores obtained from the short/truncated long test detailed in Section E.2.
+C.2 In-domain prompt generation test
+(a) The box plot for the Aesthetic Score result of short prompt input.
+(b) The box plot for the AI Corrupt score result of short prompt input.
+(c) The KDE plot for the Aesthetic Score result of short prompt input.
+(d) The KDE plot for the AI Corrupt Score result of short prompt input.
+Figure 10: The distribution of aesthetic and ai corrupt score for short prompt input.
+(a) The box plot for the Aesthetic Score result of truncated long prompt input.
+(b) The box plot for the AI Corrupt score result of truncated long prompt input.
+(c) The KDE plot for the Aesthetic Score result of truncated long prompt input.
+(d) The KDE plot for the AI Corrupt Score result of truncated long prompt input.
+Figure 11: The distribution of aesthetic and ai corrupt score for truncated long prompt input.
+Figures 10 and 11 display the box plots and KDE plots of aesthetic scores and AI corruption scores obtained from the In-domain prompt generation test detailed in Section E.2.
 While the box plots reveal that the differences in performance between various methods are subtle, the AI corruption scores provide valuable insights.
 Specifically, these scores indicate that implementations supported by TIPO produce more stable output images compared to other methods.
-(a) The box plot for the Aesthetic Score result of short caption input.
-(b) The box plot for the AI Corrupt score result of short caption input.
-Figure 7: The box plot for short caption input.
-(a) The box plot for the Aesthetic Score result of truncated long caption input.
-(b) The box plot for the AI Corrupt score result of truncated long caption input.
-Figure 8: The box plot for truncated long caption input.
+C.3 Out of focus evaluation
+(a) Original Caption
+(b) Prompt by GPT4o-mini
+(c) Prompt by Prompt-DB
+(d) Prompt by Promtist
+(e) Prompt by TIPO
+Figure 12: The similarity matrix for the 100 best aesthetic results generated in the SD3.5-Large experiments.
+Off-diagonal elements of the matrix indicate the similarity between different images.
+A lower value for an off-diagonal element indicates greater diversity among the generated images.
+(a) Original Caption
+(b) Prompt by GPT4o-mini
+(c) Prompt by Prompt-DB
+(d) Prompt by Promtist
+(e) Prompt by TIPO
+Figure 13: The similarity matrix between 100 images of worst aesthetic generated results of SD3.5-Large experiments.
+Figures 12 and 13 present similarity matrices for different prompt generation methods and their corresponding aesthetic outputs on SD3.5-Large  [22].
+A matrix with predominantly lower similarity values (overall brighter appearance) indicates high diversity across all generated images, while higher similarity values (overall darker appearance) suggest consistent but less diverse outputs.
+For quantitative diversity metrics, please refer to Table 3 in Section 5.
 Appendix D TIPO example
@@
 In this section, we present sample images from the experiments described in Section 5 to visually demonstrate the improvements achieved by TIPO.
-E.1 Scenery Tag Test
-(a) Scenery Tag Only
-(b) TIPO + Scenery Tag
+E.1 In-domain test regard to scenery tag
 Figure 16: Comparison of generated images using simple input (left) vs. TIPO-enhanced input (right) for the scenery tag
@@
 The left column shows results from simple input (scenery tag only), while the right column illustrates the enhanced diversity achieved with TIPO-enhanced input.
-E.2 Short/Truncated Long Test
+E.2 In-domain prompt generation test
 (a) Short Caption
 (b) TIPO-Generated Caption
-(c) Truncated Long Caption
-(d) TIPO-Extended Caption
 Figure 17: Comparison of generated images using original input (left) vs. TIPO-enhanced input (right)
@@
 Appendix F Topic Distribution Visualization
-To analyze the meaningfulness and diversity of topics within text inference results, we implemented a multi-stage topic modeling and clustering methodology.
-Initially, we employed Latent Dirichlet Allocation (LDA) to extract varying numbers of topics—specifically 20, 30, 50, and 100—from the corpus [8].
-This approach aimed to identify representative words that are significant within the dataset while filtering out common stop words, ensuring that nonsensical or irrelevant terms are excluded from topic classification.
+Latent Dirichlet Allocation (LDA) [9] is a generative probabilistic model for topic modeling [31], which assumes that each document is a mixture of topics, with each topic represented by a distribution over words.
+LDA uncovers hidden thematic structures by analyzing word co-occurrence patterns, while methods like TF-IDF and TextRank [42] enhance its ability to extract meaningful insights from large textual datasets.
+We implemented a multi-stage topic modeling and clustering methodology using LDA to extract varying numbers of topics (20, 30, 50, and 100) from the corpus.
+This approach focuses on identifying significant representative words while filtering out stop words and irrelevant terms to ensure meaningful topic classification.
 By employing multi-level topic analysis, we empirically assessed whether the resulting topics were sufficiently large and diverse.
-This iterative process mitigates potential challenges such as substantial topic overlap, which can diminish distinctiveness when extracting a large number of topics (Stevens et al., 2012).
-To address the potential overlap and further assess the diversity and meaningfulness of the topics, we performed a secondary clustering.
-We grouped the initially extracted topics into five major clusters using k-means clustering and evaluated the clustering performance by calculating the inertia (Sum of Squared Distances) [23], shown in Table 4, 5, and 6.
+This iterative process mitigates potential challenges such as substantial topic overlap, which can diminish distinctiveness when extracting a large number of topics [59].
+To address the potential overlap and further assess the diversity and meaningfulness of the topics, we performed a secondary clustering [66].
+We grouped the initially extracted topics into five major clusters using k-means clustering and evaluated the clustering performance by calculating the inertia (Sum of Squared Distances) [27], shown in Table 7, 8, and 9.
 Since the topics have already been filtered for meaningful content, a higher inertia value indicates greater diversity among the clusters, reflecting a broader range of valid and meaningful topics across the dataset.
 This two-tiered approach allows for a more nuanced analysis of topic diversity and ensures the robustness of the topic modeling against meaningless word groupings.
-Table 4: Inertia for COYO-Dataset inference, higher is better
 Size Promptdb GPT4o-mini Promptist TIPO
@@
 100 1656.74 1245.60 1987.63 1628.32 877.36 657.14 1622.79 1777.61
-Table 5: Inertia for GBC-Dataset inference, higher is better
+Table 7: Inertia for COYO-Dataset inference, higher is better
 Size Promptdb GPT4o-mini Promptist TIPO
@@
 100 1291.60 1742.36 1675.41 1550.32 631.61 628.78 1573.47 1855.90
-Table 6: Inertia for Scenery extend inference, higher is better
+Table 8: Inertia for GBC-Dataset inference, higher is better
 Size PromptDB GPT4o-mini Promptist TIPO
@@
 100 2026.39 1879.08 802.93 1883.70
+Table 9: Inertia for Scenery extend inference, higher is better
 We attach simple visualization of topics in scenery prompt generation, with topic n=100, cluster k=5.
 Figure 18: Topic visualization for scenery prompt generation. A wider spread indicates a greater diversity of generated topics.
+Figure 19: This visualization represents a filtered subset of posts from the Danbooru2023 dataset, centered on the ’scenery’ tag.
+The network graph is an ego network (depth 1), meaning it includes only nodes directly connected to the ’scenery’ tag.
+To refine the data and focus on meaningful associations, uncommon tags with fewer than 10 occurrences were excluded.
+The analysis, conducted using Gephi, focuses on nodes with a degree greater than 600 to highlight critical components.
+Nodes are color-coded by modularity class by Fast Unfolding Algorithm [10], revealing clusters of closely associated tags.
+Node size reflects Eigenvector Centrality[11], emphasizing tags that are both highly connected and influential within their network.

### v3

@@
 Abstract
-TIPO (Text to Image with text pre-sampling for Prompt Optimization) is an innovative framework designed to enhance text-to-image (T2I) generation by language model (LM) for automatic prompt engineering.
-By refining and extending user-provided prompts, TIPO bridges the gap between simple inputs and the detailed prompts required for high-quality image generation.
-Unlike previous approaches that rely on Large Language Models (LLMs) or reinforcement learning (RL), TIPO adjusts user input prompts with the distribution of a trained prompt dataset, eliminating the need for complex runtime cost via lightweight model.
-This pre-sampling approach enables efficient and scalable prompt optimization.
-Experimental results demonstrate TIPO’s effectiveness in improving aesthetic scores, reducing image corruption, and better aligning generated images with dataset distributions.
-These findings highlight the critical role of prompt engineering in T2I systems and open avenues for broader applications of automatic prompt refinement.
+TIPO (Text-to-Image Prompt Optimization) introduces an efficient approach for automatic prompt refinement in text-to-image (T2I) generation.
+Starting from simple user prompts, TIPO leverages a lightweight pre-trained model to expand these prompts into richer, detailed versions.
+Conceptually, TIPO samples refined prompts from a targeted sub-distribution within the broader semantic space, preserving the original intent while significantly improving visual quality, coherence, and detail.
+Unlike resource-intensive methods based on large language models (LLMs) or reinforcement learning (RL), TIPO provides computational efficiency and scalability, opening new possibilities for effective, automated prompt engineering in T2I tasks.
+We provide visual results, human preference report to investigate TIPO’s effectiveness.
+Experimental evaluations on benchmark datasets demonstrate substantial improvements in aesthetic quality, significant reduction of visual artifacts, and enhanced alignment with target distributions along with significant human preference proficiency.
+These results highlight the importance of targeted prompt engineering in text-to-image tasks and indicate broader opportunities for automated prompt refinement.
 1 Introduction
-Recent advancements in Text-to-Image (T2I) generative models have revolutionized creative applications [47, 7, 21, 55, 51, 52, 57, 53, 48, 56, 15, 14, 34, 22, 8].
-These models tend to perform better when provided with longer, more detailed prompts, which describes specific elements like style, composition, and context.
-However, this reliance on highly descriptive inputs is a significant bottleneck for generating high-quality images, especially for users who prefer or require simpler inputs.
-The need for intricate, nuanced prompts often results in a higher barrier to entry, limiting accessibility and ease of use for those unfamiliar with the complexities of prompt engineering.
-Previous works have attempted to extend prompts with optimization through various methods.
-Some approaches use LLMs to expand prompts [41]; however, due to the misalignment between the LLM’s text-rich training corpus and the T2I model’s caption-focused corpus, only a limited set of manually-engineered prompts yield effective results.
-Other methods rely on curated prompt sets [5, 17], which, while potentially incorporating user preferences, fail to fully leverage the T2I model’s trained prompt distribution.
-Additionally, reinforcement learning-based techniques [26] require multi-turn inference, incurring high computational costs, and often struggle to generate diverse outputs due to the model’s inherent corpus limitations.
-To address these challenges, we introduce TIPO (Text to Image with text pre-sampling for Prompt Optimization), an innovative framework designed to enhance T2I generative models.
-TIPO introduces ‘text pre-sampling’ preprocess procedure in inference pipeline, letting text-to-image models reproduce a more accurate distribution of corresponding user prompts without additional manual engineering.
-Our work makes the following key contributions:
-  1. 1.
-    We introduce an universal prompt optimization framework that is designed to work with any text-to-image model, providing a versatile solution for improving image generation across different architectures.
-  2. 2.
-    We provide comprehensive evaluation methodology and results including using both standard and non-standard metrics.
-Our experiments demonstrate that TIPO’s prompt engineering method can improve image quality in terms of mathematical measures and user preferences, validating its effectiveness from multiple perspectives.
-Figure 1: Comparison of prompt optimization methods.
+The rapid proliferation of Text-to-Image (T2I) generative models have revolutionized artistic creation [48, 7, 20, 57, 53, 54, 60, 55, 50, 59, 15, 14, 35, 21, 8].
+These models offer direct control over generative visual content via text prompts.
+To achieve precise control, modern T2I architectures are often trained on lengthy, detailed text descriptions, which may consist of individual, formatted tags of objects, backgrounds, styles, or complex, integrated paragraphs outlining image content and layout.
+However, the increasing complexity of prompts also forces users to iteratively refine their inputs, making high-quality T2I artwork accessible only to those with significant prompt engineering expertise.
+Recently, efforts have been made to reduce the reliance on human expertise through prompt optimization, i.e., expanding and refining a user’s primitive input into a more detailed prompt to enhance generation quality.
+A straightforward approach is to leverage pre-trained Large Language Models (LLMs) to rewrite prompts in a zero-shot manner [42].
+Yet, LLMs are primarily trained on general natural language, such as paragraphs and dialogues, which differ significantly from the structured prompts used for T2I models.
+This discrepancy often leads to additional effort in crafting LLM prompts and increased misalignment between generated images and intended prompts.
+A more effective approach is to train LLMs directly on prompt data collected from model users [5, 17].
+While promising, this method is inherently constrained by the varying levels of user expertise, often resulting in inconsistent or suboptimal outputs.
+Recent work [26] trains LLM with reinforcement learning, where aesthetic scores of generated images serve as rewards.
+However, the reinforcement learning is performed on one specific T2I model with high computational cost, hindering its application to a broader variety of models.
+Figure 1: Comparison of prompt optimization methods using LLM.
 (a) uses instructions for prompting but its understanding is constrained by the LLM’s knowledge base, not T2I model.
@@
 (d) aligns prompts with the T2I model’s training distribution, ensuring detailed and various topic-related prompt generation that fits the target T2I model.
-This paper presents TIPO’s theoretical foundations (Section 3), implementation details (Section 4), and experimental results (Section 5) demonstrating its effectiveness in improving aesthetic scores [19], reducing image corruption [43, 44], and better aligning generated images with dataset distributions [28, 46].
-Our findings highlight the critical role of prompt optimization in advancing T2I technology and open avenues for broader applications of automatic prompt refinement.
-2 Related Works
-2.1 Direct Prompt Refinement Through LLMs
-With the advent of advanced LLMs, several projects have leveraged their few-shot in-context learning capabilities to optimize prompts for Text-to-Image (T2I) models.
-For instance, CogView3 [67] and the work by Lee et al.
-[32] utilize GPT-J and Text Style Transfer (TST) techniques for prompt refinement.
-However, this approach relies heavily on the LLM’s inherent knowledge of visual content descriptions, which can be unpredictable and may not always align with specific T2I model requirements.
-2.2 Extending Prompts with Manually Collected Data
-In the open-source T2I community, a common approach involves using curated datasets of “good prompts” to fine-tune or train LLMs for prompt extension [5, 60, 17].
-These models are typically trained on image-text pairs collected from users of proprietary T2I services like Midjourney [^1] https://www.midjourney.com or Dall-E 3 [7].
-While this method is straightforward and somewhat effective, it often lacks the versatility to adapt to diverse T2I models beyond the one used in the original dataset collection.
-2.3 Image Output-Based Tuning
-Recent works have explored more sophisticated prompt optimization techniques based on image outputs.
-Reinforcement Learning Approaches: Promptist [26] introduced a two-stage framework combining supervised fine-tuning with reinforcement learning (RL).
-The RL stage incorporates aesthetic and relevance scores as environmental feedback, enabling the model to learn optimal prompt refinement strategies.
-Iterative Optimization: OPT2I [41] proposed a backpropagation-free prompt optimization framework.
-This method uses scorers and LLMs to iteratively update input prompts, aiming to improve aesthetic quality and coherence.
-These RL-based or RL-like approaches allow for prompt refinement tailored to specific model properties or user preferences.
-However, they often require extensive computational resources and may not generalize well across different T2I models or datasets.
-While these methods have shown promise, they often face challenges such as difficulty in maintaining alignment with original user intent, challenges in collecting manually engineered prompts, high computational costs for RL-based methods, and limited generalization across diverse T2I models.
-By contrast, TIPO offers an efficient approach to prompt optimization by directly leveraging the training dataset, removing the need for manual prompt engineering and runtime tuning, and achieving versatile, contextually aligned image generation.
+Deviating from existing work, we argue that the optimal prompt lies in T2I models’ training set.
+Precisely, when a prompt aligns with the data distribution used for model training, the model can better interpret the prompt, resulting in improved text alignment and higher image quality.
+Based on this insight, we introduce TIPO (Text to Image with text pre-sampling for Prompt Optimization), an innovative framework designed to enhance T2I generative models.
+At its core lies a lightweight language model trained on multiple meticulously designed pretext tasks that transform prompts into tags or sentences from coarse to fine.
+Leveraging the trained model, TIPO can progressively and seamlessly adapt to various user input types, generating accurate and robust prompts for various T2I models.
+In addition, our training data is derived from a curated text description dataset comprising over 30 million text-image pair samples and more than 20 billion tokens, ensuring comprehensive coverage of T2I training sets.
+Extensive experiments on test prompts from in- and out-of-domain demonstrate that TIPO consistently outperforms state-of-the-art methods in fidelity, aesthetics, and text alignment.
+Detailed analysis shows that TIPO has up to 62.80% win rate in user preference and 25% improvement in runtime efficiency over the second-best baseline.
+To summarize, our contributions are at least three-fold:
+  1. 1.
+    We introduce TIPO, a prompt optimization framework that leverages the large-scale text data distributions used in text-to-image (T2I) training.
+  2. 2.
+    We train a multi-task language model, which adaptively and progressively refines user inputs from either tag-based or natural language prompts into unified formats, enhancing compatibility across a broad spectrum of T2I models.
+  3. 3.
+    Extensive experiments demonstrate that TIPO achieves superior image quality, strong text alignment, higher human preference, and improved runtime efficiency across both in-domain and out-of-domain prompts, highlighting its great potential in real-world applications.
+2 Related Work
+Prompt optimization for T2I models typically leverages language models to refine user inputs.
+They can be broadly classified into two categories: (1) Model-specific strategies that tailor prompts for a particular T2I model, and (2) Universal strategies that aim to improve prompt quality across a variety of T2I models.
+2.1 Model-specific Strategies
+T2I models generate images whose quality is often measured using metrics such as fidelity, aesthetics, and user preference.
+These metrics facilitate reinforcement learning approaches that optimize prompts for a specific T2I model.
+For instance, Promptist [26] fine-tunes a pre-trained language model by using CLIP relevance scores as rewards.
+Similarly, PAE [44] extends this approach by generating dense text embeddings rather than discrete text tokens, with additional control vectors during online reinforcement learning.
+However, these methods are computationally intensive, often struggling with a larger number of training prompts.
+Moreover, a model optimized for one specific T2I system may not generalize well to others.
+In contrast, our method leverages over 30 million text descriptions to cover a wide range of high-quality prompts, ensuring compatibility with a broad spectrum of T2I models.
+2.2 Universal Strategies
+To reduce the dependency on specific T2I models, some researchers have focused on refining prompts solely using language models.
+For example, CogView3 [71] employ GLM-4 [24], and Lee et al.
+[33] employ GPT-J and Text Style Transfer (TST) techniques, respectively, for prompt enhancement.
+However, both of them rely heavily on the LLM’s inherent understanding of visual content descriptions, which may result in a misalignment with the diverse requirements of various T2I models.
+Alternatively, other approaches collect high-quality prompts from T2I model users to fine-tune or train LLMs [5, 63, 17].
+Such methods, however, are limited by the inconsistent expertise of users.
+Conversely, our approach constructs both tag-based and natural language prompts using a large-scale dataset of image-text descriptions, thereby aligning closely with the text distributions underlying T2I models.
+3 Preliminaries
+Notations. We present the notations in our work:
+  • $\mathcal{P}$: The space of all possible prompts.
+  • $\mathcal{I}$: The space of all possible images.
+  • $p$: User-provided prompt.
+  • $p_{o}$: Optimized prompt.
+  • $p_{ps}$: Sampled modified prompt derived from p.
+  • $\mathcal{M}$: Metadata (e.g., style tags) aiding prompt transformation.
+  • $\mathcal{T}$: Transformation function for pre-sampling.(e.g., Random augmentation, TIPO or LLM)
+  • $T_{p},T_{ps}$: Sets of tags for original and pre-sampled prompts.
+  • $S_{p},S_{ps}$: Ordered lists of natural language descriptions for original and pre-sampled prompts.
+Text-to-Image Model. A text-to-image (T2I) model is a mapping function
+$f(p)\;\;\to\;\;\mathcal{I}_{p}\subseteq\mathcal{I},$
+which maps a user-provided text prompt $p$ to a subset of images $\mathcal{I}_{p}$ corresponding to the prompt.
+Problem statement.
+Let $\mathcal{I}_{u}$ denote the user’s intended distribution over images.
+The task of Prompt optimization is to generate a prompt $p_{o}$ that minimizes a distance $d$ between the T2I output distribution $\mathcal{I}_{p}$ and $\mathcal{I}_{u}$:
+$p_{o}\;=\;\arg\min_{p\in\mathcal{P}}\;d\bigl{(}\mathcal{I}_{p},\,\mathcal{I}_{u}\bigr{)},$
+where $d$ is a suitable metric (e.g., Frechet Inception Distance) comparing two image distributions.
 […]
-3 Prompt Optimization
-3.1 Motivation and Challenges
-Nowadays text-conditioned models [8, 48, 21] generate outputs based on user-provided text inputs.
-However, when users seek high-quality outputs or wish to explore diverse outputs while preserving key aspects of their input, they often need to manually expand their prompts.
-This manual prompt expansion can be time-consuming and labor-intensive.
-To address these challenges, previous works have aimed to automate the prompt optimization process.
-Some methods attempt to expand prompts using the familiar vocabulary and stylistic tendencies of the model [25], while others rely on direct feedback from the model to iteratively improve prompt alignment [26].
-However, these approaches have limitations: if the structure or language does not fully align with the model’s prompting conventions, performance may degrade, and feedback-based approaches often involve significant computational costs.
-Therefore, effective prompt optimization for text-conditioned models requires accurately capturing user intent and expanding the prompt in a way that is both precise and aligned with the model’s conventions for interpreting text input.
-This need for precision and alignment motivates a systematic approach to prompt formulation, which we define in the following section.
-3.2 Formulation
-To formalize prompt extension, we define $\mathcal{P}$ as the set of all possible prompts, $\mathcal{N}$ as the Gaussian noise vector space, and $\mathcal{I}$ as the set of all possible images.
-A T2I model then acts as a mapping function:
-$f(p):\mathcal{N}\rightarrow\mathcal{I}_{p}$
-where, for any prompt $p\in\mathcal{P}$, the model maps noise vectors from $\mathcal{N}$ to a subset of images $\mathcal{I}_{p}\subseteq\mathcal{I}$ associated with the prompt $p$.
-Now we define a pair of prompts: simple prompt $p_{s}$ and detailed prompt $p_{d}$.
-Formally, for a simple prompt $p_{s}$ and a detailed prompt $p_{d}$ that extends $p_{s}$, we have the following relationship:
-$f(p_{d})(\mathcal{N})\subseteq f(p_{s})(\mathcal{N})$
-This formulation implies that the image subset generated by the detailed prompt $p_{d}$ is contained within the subset generated by the simple prompt $p_{s}$, thereby preserving the relevance of the initial prompt while adding specificity.
-To further encapsulate the prompt generation process, we define the prompt extension function, $\mathcal{E}$.
-This function takes a simple prompt $p_{s}$ and additional metadata $\mathcal{M}$ (such as tags) as inputs, and produces a detailed prompt $p_{d}$ as output:
-$\mathcal{E}:\mathcal{P}_{s}\times\mathcal{M}\rightarrow\mathcal{P}_{d}$
-where $\mathcal{P}_{s}$ is the space of simple prompts and $\mathcal{P}_{d}$ is the space of detailed prompts.
-Here, $\mathcal{M}$ provides auxiliary information that guides the extension of $p_{s}$ into a more refined prompt $p_{d}$.
-For each instance of prompt extension, we apply the function as follows:
-$p_{d}=\mathcal{E}(p_{s},M)$
-where $M$ could represent a specific metadata component, such as $T_{s}$, or other forms of contextual information depending on the input type.
-The simplest approach to extend a simple prompt is by adding random words.
-However, this approach has limitations, as it often introduces internally conflicting information that results in out-of-distribution (OOD) prompt inputs, leading the model to generate nonsensical images or the model to generate not user intent images.
-In such cases, although each added word may be meaningful in isolation, these random combinations can create prompts that are internally contradictory, as shown in Figure 2 (c).
-For example, the generated images may deviate from the user’s intent by not depicting an astronaut (left top), producing a horse-shaped tank (right top) instead of a horse, omitting Mars (left bottom), or even including a random snake (right bottom), which introduces nonsensical elements.
-To represent this random extension method, let $p_{r_{i}}$ denote the $i^{th}$ prompt created by adding random words to $p_{s}$.
-The resulting union of distributions fails to approximate the desired distribution.
-$\bigcup_{i=0}^{n}f(p_{r_{i}})(\mathcal{N})\supseteq\mathcal{I}\supseteq\mathcal{I}_{p_{s}}$
-Which means:
-$\bigcup_{i=0}^{n}f(p_{r_{i}})(\mathcal{N})\not\approx\mathcal{I}_{p_{s}}$
-This scenario is illustrated in Figure 2 (c), where “Simple Prompt + Random Words” shows a pointy distribution extending beyond the bounds of the desired distribution or even the “image” distribution itself.
-4 TIPO Framework and Model
-4.1 Distribution-Aligned Prompt Extension
-Central to TIPO’s method is an extension function $\mathcal{E}$, which iteratively refines prompts to enhance both the diversity and quality of generated images while remaining aligned with the intent of the original user-provided prompt and the training data distribution.
-By applying $\mathcal{E}$ to a simple prompt $p_{s}$ to generate a detailed prompt $p_{d}$, TIPO aims to approximate the target distribution:
-$\bigcup_{i=0}^{n}f(p_{t_{i}})(\mathcal{N})\approx\mathcal{I}_{p_{s}}$
-where $p_{t_{i}}$ denotes the $i^{th}$ TIPO-generated prompt derived from $p_{s}$, and $\mathcal{I}_{p_{s}}$ represents the target distribution of image outputs associated with $p_{s}$.
-Through this systematic prompt refinement, Figure 2 (d) “Simple Prompt + TIPO” illustrates the broader image space exploration enabled by TIPO while retaining relevance to the original prompt, effectively balancing diversity with coherence.
-4.2 Prompt Structure
-The construction of simple prompts, $p_{s}$, and their corresponding detailed prompts, $p_{d}$, varies depending on the dataset and input modality.
-We consider two primary scenarios: tag-based captions and natural language (NL) captions.
-Tags are used to describe the content of images and are commonly formatted as a comma-separated string when input into T2I models.
-For a given image-tag pair, let $T_{n}=\{t_{1},t_{2},t_{3},\ldots,t_{n}\}$ represent the complete set of tags for an image.
-We derive the simple tag set $T_{s}=\{t_{1},t_{2},\ldots,t_{m}\}$ where $m<n$, and define the detailed tag set as $T_{d}=T_{n}$.
-The simple and detailed tag prompts are then constructed as:
-$p_{s}=\text{concat}(T_{s}),\quad p_{d}=\text{concat}(T_{d})$
-NL captions are typically generated by LMs with vision modality [36, 1, 2, 33, 6, 16, 63, 18] and may contain multiple sentences.
-Each image is associated with either two types of NL captions—a short NL caption $S_{s}$ providing a brief, general description, and a long NL caption $S_{d}$ offering an extended, detailed description—or a single long caption that can be segmented.
-For images with a single long caption, let
-$S_{d}=\{\text{sentence}_{1},\text{sentence}_{2},\text{sentence}_{3},\ldots,\text{sentence}_{n}\}$
-represent the complete set of sentences in $p_{d}$. We derive the simple sentence set
-$S_{s}=\{\text{sentence}_{1},\text{sentence}_{2},\ldots,\text{sentence}_{m}\}$
-where $m<n$, and define the detailed sentence set as $S_{d}=S_{n}$. The simple and detailed NL prompts are then constructed as:
-$p_{s}=\text{concat}(S_{s}),\quad p_{d}=\text{concat}(S_{d})$
-These ensure that $p_{d}$ is a superset of $p_{s}$, aligning with the concept that detailed prompts provide a more specific description.
-To generate $p_{d}$ from $p_{s}$, TIPO employs two distinct formats based on the relationship between $p_{s}$ and $p_{d}$:
-  1. 1.
-    When $p_{s}$ is a substring of $p_{d}$: The format <meta> <p_d> is used.
-As TIPO model is an autoregressive LM, $p_{d}$ is directly used after the <meta> tokens, allowing the model to generate subsequent tokens based on any substring of $p_{d}$.
-  2. 2.
-    When $p_{s}$ is not a substring of $p_{d}$: The format <meta> <p_s> <p_d> is employed.
-Since $p_{s}$ is not contained within $p_{d}$, both are included separately after the <meta> tokens to guide the generation process effectively.
-4.3 Prompt Formatting
-In TIPO, we define a simple and consistent format for the content of the <meta> token or the simple prompt $p_{s}$.
-The format is as follows: <Category>: <Content>.
-Metadata categories might include artist, copyright, aspect ratio, quality, and year.
-For instance,
-quality: masterpiece, year: newest
-Importantly, in certain cases, input tag or NL prompts can be treated as metadata <Content> as well, when the task involves generating tags or NL prompts conditioned on each other.
-4.4 Task Definitions
-Figure 3: TIPO’s task flow within a single generation cycle.
-Starting from any node, each arrow represents a sequential extension step, with other prompt used as metadata if provided.
+4 Methodology
+We aim to optimize user prompts to enhance image generation quality.
+Instead of end-to-end optimizations tailored to a single T2I model, our focus is on prompt rewriting that generalizes across a broad spectrum of models.
+Our core intuition is that an ideal prompt should align with the texts used in T2I model training.
+However, rapid advances in image captioning have rendered these texts increasingly diverse and complex.
+To address this, we propose to (1) design a clearly structured prompt schema compatible with most text descriptions, and (2) implement a pre-sampling algorithm that progressively refines arbitrary, coarse user input into organized, fine-grained prompts.
+4.1 Text Set Preparation
+Although the text descriptions used for T2I model training are notably diverse, most are image captions that fall into two broad categories: tag-based and natural language (NL)-based captions.
+Tag-based captions, such as those in the Danbooru2023 dataset [46, 68], use comma-separated, succinct terms to describe image content.
+In contrast, NL-based captions, typically generated by language models with visual capabilities [37, 1, 2, 34, 6, 16, 67, 18], may comprise multiple sentences.
+We represent both types using a unified text set $T=\{t_{1},t_{2},t_{3},\ldots,t_{n}\}$, where each element is an individual tag or sentence.
+While the original image captions are fine-grained and detailed, which can yield high-quality images when all elements are used, they often result in prompts that are excessively lengthy or overloaded with information.
+Such prompts diverge from typical user input and pose alignment challenges.
+To mitigate this, we construct a simpler subset $T_{s}\subset T$ by removing some tags and sentences from the original set [^1] For long sentences, off-the-shelf text splitters (e.g., FastText) are used to segment them into multiple sentences., as detailed in Section 4.2.
+4.2 Formatted Prompt Construction
+We aim to construct prompts in a unified format compatible with existing image captions.
+First, we incorporate the common metadata present in these image captions, typically represented as <Category>: <Content>.
+These metadata categories primarily include artist, copyright, aspect ratio, quality, and year (e.g., quality: masterpiece, artist: Picasso).
+This structured metadata is intuitive for users to read and edit, while also providing strong guidance to downstream T2I models on the generation scope.
+Next, we construct both tag-based and NL-based prompts using text sets $T$.
+Our design generates both simple (incomplete) and complete prompts for each image, and we train an auto-regressive language model to extend the simple prompts into complete versions.
+For tag-based prompts, since the tags are largely order-insensitive (i.e., the order has minimal impact on T2I outcomes), we propose a prefix-based dropout strategy.
+We first randomly shuffle the complete set of tags $T=\{t_{1},t_{2},t_{3},\ldots,t_{n}\}$ from a given image caption.
+Then, we construct a simple tag set $T_{s}=\{t_{1},t_{2},\ldots,t_{m}\}$ by randomly selecting $m<n$ tags.
+The prompts are constructed as:
+$p_{s}=\text{concat}(T_{s}),\quad p_{o}=\text{concat}(T)$
+Because $p_{s}$ is always a prefix of $p_{o}$, the language model can readily expand the simple tag-based prompt into its complete version.
+For NL-based prompts, however, this strategy cannot be applied directly because the first sentence often contains crucial information [25] and the order of sentences significantly influences the caption’s semantics.
+Therefore, we preserve the first sentence and randomly drop some of the subsequent sentences without changing their order.
+Let:
+$S=[\text{sentence}_{1},\text{sentence}_{2},\text{sentence}_{3},\ldots,\text{sentence}_{n}]$
+represent the ordered sequence of sentences in an image caption.
+We derive a simple subsequence $S_{s}$ by randomly selecting $m$ sentences from $S$ while ensuring that the first sentence is always included and that the original order is maintained.
+In other words,
+$S_{s}=[\text{sentence}_{1},\text{sentence}_{i_{2}},\ldots,\text{sentence}_{i_{m}}],$
+with $1<i_{2}<\ldots<i_{m}\leq n$ and $m<n$. The simple and complete NL-based prompts are then constructed as:
+$p_{s}=\text{concat}(S_{s}),\quad p_{o}=\text{concat}(S_{s},S)$
+This ensures that $p_{s}$ remains a prefix of $p_{o}$.
+Although some sentences may be repeated in $p_{o}$, selecting a smaller $m$ effectively mitigates this, and it does not empirically affect the generation quality.
+4.3 Text Pre-sampling
+We aim to align user input with the constructed high-quality training prompts $p_{o}$ via pre-sampling, which stands for “text sampling before image sampling” - a process that first constrains the target distribution in text space before proceeding to image generation.
+Conceptually, while pre-sampling strategies generally add more text tokens to user input, the outcomes of different strategies may still vary (as illustrated in Figure 2).
+To enrich the image content while mitigating undesired output, we perform pre-sampling based on a language model trained on $p_{o}$.
+A straightforward pre-sampling strategy is to do text completion, i.e., directly append output tokens to the end of user input.
+However, we posit that the user input is unlikely to be a part of a good prompt.
+The na’́ive text-completion strategy prioritizes the user’s original input, resulting in worse outcomes as evidenced by Section 5.6.
+Figure 3: TIPO’s task flow within a single generation.
+Starting from any node, each arrow represents a sequential extension step, with other prompts used as metadata if provided.
 This task design enables efficient and flexible prompt extension across multiple tasks.
-TIPO defines three core tasks: extending tag sequences, extending NL prompts, and generating refined NL prompts.
-Each task is represented by an arrow in Figure 3, illustrating how TIPO can use tag and NL prompts interchangeably as input sources to generate detailed outputs within a single generation cycle.
-When an arrow cycle is established, TIPO can complete multiple extension within a single generation pass.
-For example, TIPO can start from a short NL prompt to produce a detailed tag sequence, or it can combine tasks like short_to_tag and tag_to_long to create a composite task, short_to_tag_to_long, which generates an extended NL prompt from an initial short NL input via tag generation.
-Further details on specific task sequences are provided in Appendix.
-4.5 An Example Scenario and Workflow
-Figure 4: An example scenario of TIPO workflow.
-(a) A generated image and prompts.
-(b) Prompt optimization of TIPO, starting from simple user input $p_{s}$ to detailed final output $p_{d}$.
-The shading from gray to light sky blue represents an increase in context richness in the prompt.
+We propose the core technique of TIPO, a flexible pre-sampling mechanism that divides the prompt optimization process into three interconnected subtasks: enriching tag sequences, extending NL prompts, and refining NL prompts.
+As shown in Figure 3, TIPO uses both tag and NL prompts interchangeably as input sources to generate detailed outputs in various forms.
+For instance, TIPO can begin with a short NL prompt to produce a detailed tag sequence, denoted as short_to_tag.
+Training data for the language model is constructed by concatenating different types of $p_{o}$ with task-specific metadata embedded into the text.
+The individual tasks are defined as follows:
+  • tag_to_long: Use tags as metadata to generate a new NL prompt.
+  • long_to_tag: Use an NL prompt as metadata to extend a tag sequence.
+  • short_to_tag: Use the simple prompt $p_{s}$ as metadata to extend a tag sequence.
+  • short_to_long: Use a user-provided NL prompt as metadata to generate a refined, detailed prompt.
+We also consider composite tasks that can be performed within a single forward pass. These tasks offer the language model a holistic view of the objective while reducing computational overhead:
+  • short_to_tag_to_long: Use a user-provided NL prompt or tag sequence as metadata to generate a refined, detailed prompt.
+  • short_to_long_to_tag: Use a user-provided NL prompt or generated NL prompt as metadata to extend a tag sequence.
+  • tag_to_short_to_long: Use user-provided tags or NL prompts as metadata to generate a refined NL prompt.
+We randomly select from the aforementioned tasks during training to enhance model generalization.
+By extensively training on these tasks, TIPO can seamlessly adapt to various input types, flexibly refining user input whether it consists of tags, short sentences, or long sentences.
 Figure 4 illustrates a scenario where both tag captions $T_{s}$ and short NL captions $S_{s}$ are available.
-In such cases, TIPO processes each input type separately to maintain clarity and coherence.
-During each generation cycle, one input type is treated as the primary prompt which is target prompt for extending or refining, while the other types are used as <Content> of metadata to enrich the context.
-Figure 4 provides an example of this setup, where:
-$\begin{array}[]{ll}S_{s}&=\{\textsf{A young girl with long hair...}\},\quad\verb|<meta>|=\emptyset{}\\ T_{s}&=\left\{\textsf{outdoors, scenery, water, wind, landscape, ...}\right\}\end{array}$
+In such cases, TIPO processes each input type separately to maintain clarity and coherence:
+$\begin{array}[]{ll}S_{s}&=\{\textsf{A young girl with long hair...}\},\quad\texttt{metadata}=\emptyset\\ T_{s}&=\left\{\textsf{outdoors, scenery, water, wind, landscape, \ldots}\right\}\end{array}$
 The generation proceeds sequentially as follows:
   1. 1.
-    short_to_tag: TIPO uses $T_{s}$ as the primary prompt to generate a detailed tag sequence $T_{d}$
+    short_to_tag: TIPO uses $T_{s}$ as the primary prompt to generate a detailed tag sequence $T_{d}$.
   2. 2.
@@
   3. 3.
-    short_to_tag_to_long: With both $T_{d}$ and $S_{s}$ in the metadata, TIPO generates a comprehensive long NL prompt $S_{d}$, ensuring a more detailed prompt.
+    short_to_tag_to_long: With both $T_{d}$ and $S_{s}$ in the metadata, TIPO generates a comprehensive long NL prompt $S_{d}$, ensuring a more detailed output.
   4. 4.
-    TIPO aggregates $T_{d}$, $S_{d}$, and any additional metadata to construct a context rich prompt, $p_{d}$.
-This iterative process allows TIPO to build prompts that are both detailed and contextually aligned with user inputs, through short_to_tag, tag_to_long, and short_to_tag_to_long tasks within a single generation cycle.
-4.6 Model Architecture
-TIPO utilizes a causal autoregressive language model framework as its prompt extension function, $\mathcal{E}$, to iteratively extend and refine user-provided prompts into more detailed, content-rich prompts $p_{d}$ based on simpler inputs $p_{s}$.
-Our implementation leverages the LLaMA architecture [61, 62, 3], employing models with 200M and 500M parameters.
+    TIPO aggregates $T_{d}$, $S_{d}$, and any additional metadata to construct a context-rich prompt $p_{d}$.
+This progressive process enables TIPO to build prompts that are both detailed and contextually aligned with the user’s input.
+Figure 4: An example scenario of TIPO workflow.
+(a) A generated image and prompts.
+(b) Prompt optimization of TIPO, from simple user input $p_{s}$ to detailed final output $p_{d}$.
+The shading from gray to light sky blue represents an increase in context richness in the prompt.
+4.4 Implementation Details
+TIPO can utilize various types of causal autoregressive language models by design. In practice, we leverage the LLaMA architecture [64, 65, 3] for simplicity, with 200M and 500M parameters.
   • TIPO-200M: Pretrained on the Danbooru2023 [46, 69] and GBC10M [29] datasets for 5 epochs, then fine-tuned with Danbooru2023, GBC10M, and CoyoHD11M [12] for 3 epochs, covering approximately 40 billion tokens.
   • TIPO-500M: Pretrained on the Danbooru2023, GBC10M, and CoyoHD11M datasets for 5 epochs, covering approximately 30 billion tokens.
-Throughout training, non-padding tokens were used for token count calculations, with the effective count reflecting the relatively short and variable length of data entries.
-4.7 Training Process
-In each training pass, TIPO applies dataset augmentation by selecting tasks from all possible task combinations such as short_to_tag_to_long, including extending tag sequences, NL prompt extension, and refined NL prompt generation.
-This random selection across all task combinations significantly broadens the dataset’s effective scope, enhancing model generalization without additional manual engineering or runtime tuning.
-(a) Direct LLM
-(b) Prompt DB
-(c) RL (Promptist)
-(d) TIPO (Ours)
-Figure 5: In-domain test regard to scenery test examples from 4 different methods.
+5 Experiments
+5.1 Experimental Settings
+Baselines. We compare with state-of-the-art prompt optimization methods as follows:
+  • GPT [2] series are powerful large language models capable of zero-shot prompt optimization. We use GPT-4o-mini for lower cost and better efficiency.
+  • MagicPrompt [17] trains GPT-2 on high-quality prompts collected from stable-diffusion model users.
+  • Promptist [26] optimizes user input into model-preferred prompts via reinforcement learning.
+Evaluation Metrics.
+We employ four latest metrics FDD [61], Aesthetic Score [19], AI Corrupt Score [45], and Vendi Score [22] to measure the quality of generated images.
+Specifically, FDD (Frechet Dino Distance) measures the image fidelity using the similarity between generated and dataset images via DinoV2 features [47], which better aligns with human perception than traditional FID [28].
+Aesthetic Score is computed via Aesthetic Predictor V2.5 [19], quantifying visual appeal, composition quality, and artistic merit.
+AI Corrupt Score detects technical flaws in generated images by identifying visual artifacts.
+Vendi Score quantifies image diversity by calculating the von Neumann entropy from a normalized cosine similarity matrix using DinoV2 embeddings.
+Evaluation Protocols.
+Our proposed TIPO leverages large-scale image caption datasets for training, which overlap with the training text distributions of many T2I models.
+Following Promptist [26], we divide our experiments into two settings: (1) In-domain, where the T2I model’s training texts overlap with those used by TIPO, and (2) Out-of-domain, where no overlap exists.
+5.2 In-domain Tag-based Prompt Optimization
+To assess prompt optimization performance on tag-based prompts, we generate scenery images, as they contain abundant descriptive tags for objects and backgrounds.
+We randomly sample 32,768 tag-based prompts from Danbooru2023 [69, 46], shuffle and concatenate the scenery tags into new prompts (thereby preventing data leakage of the original captions), and generate one image per prompt using Kohaku-XL-Zeta, an SDXL [50, 46] variant fine-tuned on Danbooru2023.
+This evaluation tests the in-domain capabilities, as Danbooru2023 is also used during TIPO training.
+Metric Original GPT MagicPrompt Promptist TIPO (Ours)
+FDD $\downarrow$ 0.3558 0.5414 0.3247 0.2350 0.2282
+Aesthetic Score $\uparrow$ 5.0569 6.3676 6.1609 5.9468 6.2571
+AI Corrupt Score $\uparrow$ 0.4257 0.7490 0.5024 0.5669 0.9195
+Vendi Score $\uparrow$ 16.8135 8.6632 11.9014 14.3273 13.3070
+Table 1: Performance on in-domain tag-based prompts, with best results bolded and the second-best underlined.
+The results in Table 1 reveal two key insights.
+First, MagicPrompt and Promptist, which rely on user prompts or reinforcement learning, underperform in Aesthetic and AI Corrupted Scores due to the quality or quantity limitations of their collected samples (e.g., Promptist uses 90K samples, limited by the high reinforcement learning cost).
+In contrast, GPT and TIPO benefit from large-scale training corpora ($>$30M samples), yielding higher-quality outputs.
+Second, TIPO achieves the best FDD by a substantial margin over GPT, which can be attributed to its superior distribution alignment with T2I models.
+5.3 In-domain NL-based Prompt Optimization
+We evaluate the prompt optimization performance on NL-based prompts by selecting 10,000 short prompts and 10,000 long prompts from CaptionEmporium [12] and GBC [29] as test prompts.
+In particular, since the long prompts are much longer than typical user input, we truncate them to two sentences ($<$ 40 words) to simulate real-world applications.
+We use SDXL-1.0-base [50] as the T2I model, whose training text data largely overlap with TIPO.
+Short Original GPT MagicPrompt Promptist TIPO(Ours)
+FDD $\downarrow$ 0.0957 0.1668 0.0980 0.1783 0.1168
+Aesthetic Score $\uparrow$ 5.8370 6.0589 5.8213 5.7963 5.8531
+AI Corrupt Score $\uparrow$ 0.7113 0.6985 0.7064 0.6314 0.7131
+Vendi Score $\uparrow$ 38.1715 34.7139 38.1553 34.1270 37.0649
+Truncated Long Original GPT MagicPrompt Promptist TIPO(Ours)
+FDD $\downarrow$ 0.0955 0.1683 0.1247 0.2096 0.1210
+Aesthetic Score $\uparrow$ 5.7497 6.0168 5.8191 5.7759 5.8364
+AI Corrupt Score $\uparrow$ 0.6868 0.6712 0.6741 0.5925 0.7130
+Vendi Score $\uparrow$ 38.2533 34.8108 37.8412 33.5266 37.0900
+Table 2: Performance on in-domain NL-based prompts.
 (a) Scenery Tag Only
@@
 (d) Truncated Prompt + TIPO
-Figure 6: Comparison of generated images: (a) simple scenery tag, (b) scenery tag enhanced by TIPO, (c) truncated long prompt, and (d) TIPO-enhanced truncated prompt.
-TIPO adds detail and maintains variety, yielding coherent images even from minimal prompts.
-5 Evaluation
-5.1 Experimental Setup
-In this section, we detail the experimental setup used to evaluate prompt optimization framework.
-To assess the quality of generated prompts and resulting images, we employed the following evaluation metrics.
-FDD (Frechet Dino Distance).
-Traditionally, the Frechet Inception Distance (FID) [28] has been used to measure the distribution distance between datasets and generated images.
-However, recent works have shown that FID is not always aligned with human preferences [58].
-Therefore, we measure the Frechet Distance using DinoV2 [46] outputs.
-In our experiments, we used the pretrained ViT-g [20] model from DinoV2 and used the normalized embedding vector to measure the Frechet Distance result.
-Aesthetic Score.
-We compute the Aesthetic Score using the Aesthetic Predictor V2.5 [19].
-Higher aesthetic scores indicate images that are perceived as more visually appealing, well-composed, and of higher quality.
-This metric aims to quantify the subjective visual appeal of generated images, with higher scores suggesting outputs that are more likely to be judged favorably by human viewers.
-AI Corrupt Score.
-The AI Corrupt Score is obtained from the AICorruptMetrics [43] in sdeval [44].
-This metric is trained on AI-generated images with human-annotated “corrupt or not” labels.
-A higher value indicates a higher likelihood of the image being “correct” or “complete”.
-Vendi Score.
-To quantify the diversity of generated images, we employ the Vendi Score metric based on DinoV2 [46] embeddings.
-The computation involves calculating a similarity matrix $K$ using cosine similarity between image embeddings, followed by computing the von Neumann entropy of the normalized matrix $K/n$, adhering to the standard Vendi Score methodology.
-This metric provides a quantitative measure of the distributional diversity within a set of generated images.
-Higher Vendi Scores indicate greater inter-image diversity, reflecting the model’s capability to generate varied and distinct outputs.
-Extension Functions.
-We used both GPT-4o-mini [2] and LLaMA-3.2-1B [3] as Direct LLM models for prompt generation, balancing performance and efficiency.
-GPT-4o-mini, as a lightweight closed-source model, offers speed and affordability, while LLaMA-3.2-1B represents a similar scale to TIPO in open-source models.
-Also, Prompt DB [17] and Promptist [26] serve as baseline methods for prompt extension.
-5.2 In-domain test regard to scenery tag
-To evaluate the model’s capability to generate diverse and accurate outputs, we chose the scenery tag, which includes both natural landscapes and human-made structures.
-In this test, we randomly select 32,768 entries from Danbooru2023 [65, 45] that include the “scenery” tag.
-We set up the following prompt:
-<meta> + scenery + $\mathcal{E}$>
-Then, we apply extension functions to extend only the scenery, $T_{s}$, as the <Content> tag.
-We generate one image from each prompt using the Kohaku-XL-zeta [^2] https://huggingface.co/KBlueLeaf/Kohaku-XL-Zeta model, which is a fine-tuned SDXL [48] model on the Danbooru dataset [45].
-Metric Original LLM Prompt DB Promptist TIPO (Ours)
-FDD $\downarrow$ 0.3558 0.5414 0.3247 0.2350 0.2282
-Aesthetic Score $\uparrow$ 5.0569 6.3676 6.1609 5.9468 6.2571
-AI Corrupt Score $\uparrow$ 0.4257 0.7490 0.5024 0.5669 0.9195
-Table 1: Metric results on in-domain test.
-Bold values indicate the best performance, and underlined values indicate the second-best.
-Original refers to images generated using the original text captions; LLM uses GPT-4o-mini [2]; the remaining columns represent results from Prompt DB [25], Promptist [26], and TIPO, respectively.
-As shown in Table 1, TIPO-generated prompts improve FDD scores by aligning images more closely with the original dataset distribution.
-Enhanced AI Corrupt scores show that TIPO reduces image corruption, creating more coherent outputs by replacing vague descriptions (e.g., detailed background) with specific elements (e.g., garden filled with forget-me-nots and blue butterflies).
-The direct LLM approach highlights a key issue in prompt engineering: it achieves high aesthetic scores but compromises FDD, reducing the diversity of images.
-We provide more comparative visual result in Figure 5.
-The results reveal significant biases in Prompt DB  [25] and RL-based [26] approaches.
-PromptDB and Direct-LLM method exhibits notable bias, by generating mountain/lake scene majorly.
-Meanwhile, Promptist produces diverse near-future fantasy sceneries, however often produces corrupted images due to unreasonable vocabulary usage (4K, masterpiece, instagram), supposedly introduced by user prompts.
-By contrast, TIPO consistently reproduces a balanced output distribution, incorporating both natural and architectural elements, effectively mitigating the biases observed in other methods.
-(a) Original prompt best aesthetic
-(b) GPT-4o-mini best aesthetic
-(c) TIPO best aesthetic
-Figure 7: Top-10% aesthetic score images generated by the SD-3.5-Large model [19].
-Rows show samples from Original, Direct LLM-Modified prompts, and TIPO-extended prompts.
-Direct LLM images focus primarily on natural scenery, while generated image from TIPO-extended prompt display a wider variety, including urban scenes and human figures.
-5.3 In-domain prompt generation test
-We aim to test each models’ performance on regard to reconstructing / upsampling the given prompt, in given dataset distribution.
-By comparing original ground truth and generated outputs, we can directly compare FDD, aesthetic score, AI corrupt scores.
-We randomly selected 10k entries each from [12] and [29].
-For each entry, we extracted the short caption and the truncated long caption (limited to a maximum of two sentences).
-These captions serve as user input prompts $p_{s}$ for in-domain prompt generation test.
-Using these prompts, we generated single image per prompt with the SDXL-1.0-base [^3] https://huggingface.co/stabilityai/stable-diffusion-xl-base-1.0 model [48].
-Short Original LLM Prompt DB Promptist TIPO(Ours)
-FDD $\downarrow$ 0.0957 0.1668 0.0980 0.1783 0.1168
-Aesthetic Score $\uparrow$ 5.8370 6.0589 5.8213 5.7963 5.8531
-AI Corrupt Score $\uparrow$ 0.7113 0.6985 0.7064 0.6314 0.7131
-Truncated Long Original LLM Prompt DB Promptist TIPO(Ours)
-FDD $\downarrow$ 0.0955 0.1683 0.1247 0.2096 0.1210
-Aesthetic Score $\uparrow$ 5.7497 6.0168 5.8191 5.7759 5.8364
-AI Corrupt Score $\uparrow$ 0.6868 0.6712 0.6741 0.5925 0.7130
-Table 2: Performance comparison in the in-domain prompt generation test.
-Bold values indicate the best performance, and underlined values indicate the second-best.
-Direct LLM achieves the highest Aesthetic Score but struggles with OOD issues.
-TIPO provides a balanced outcome, excelling in AI Corrupt Score and maintaining strong alignment, making it ideal for practical applications.
-Table 2 presents the results of our Short/Truncated caption test for prompt generation.
-Compared to in-domain test, Section 5.2, which might be considered slightly out of domain by SDXL model, we observed that prompts covered by this test produces meaningful, coherent images with sufficient detail level.
-FDD scores were degraded compared to original-ground truth prompts, not truncated long caption-which implies prompt extension functions resulted lack of diversity, or unmatched components in generation.
-We observe that Direct LLM approach achieves the highest aesthetic scores by wide vocabulary with user preference matching, however results low FDD, AI corruption score, which implies unmatched vocabularies are being used in-prompt.
-Original LLM Prompt DB Promptist TIPO(Ours)
-Top 10% $\uparrow$ 10.98 7.84 12.17 10.12 17.33
-Worst 10% $\uparrow$ 12.10 9.76 15.23 14.14 13.92
-Table 3: Vendi Scores [23] for the top and worst 10% aesthetic images across prompt generation methods.
-Higher value indicate higher diversity of samples.
-TIPO achieves high score in the top 10% (17.33), showing greater visual diversity as well as worst 10%.
-Result suggests TIPO’s balanced and robust variety in generated images.
-Meanwhile, LLM has the lowest scores in both categories, indicating limited diversity and repetition.
-5.4 Out of focus evaluation.
-In this section, we analyze TIPO’s behavior in out-of-focus model, SD-3.5-Large [^4] https://huggingface.co/stabilityai/stable-diffusion-3.5-large built upon MMDiT architecture [21].
-We compose original tag, natural language prompt created by GPT-4o-mini, tag+natural language extended prompt by TIPO (ours), then feed to SD-3.5-Large inference pipeline.
-Analysis of Vendi Scores [23] (Table 3) reveals contrasting patterns: Direct LLM-generated images show limited diversity despite high aesthetic scores, while TIPO maintains conceptual diversity across all quality levels.
-However, TIPO’s training on diverse datasets [45, 65, 12, 29] introduces domain-specific elements that may not align with SD3.5L’s training distribution.
-This mismatch demonstrates our core finding: prompt generation models perform optimally when trained on domain-matched prompt corpora.
-The balanced profile of TIPO suggests that TIPO can work as suitable solution for prompt upsampling task, which can accurately project trained prompt with few initial word input.
+Figure 5: Generated images from 4 types of prompts: (a) simple scenery tag, (b) scenery tag enhanced by TIPO, (c) truncated ($<$ 40 words) long prompt, (d) TIPO-enhanced truncated prompt.
+TIPO adds detail and maintains variety, yielding coherent images from simple prompts.
+Table 2 demonstrates that TIPO achieves either the best or second-best scores in Aesthetic and AI Corrupt Score by effectively enriching the original prompt with appropriate textual elements while rarely introducing extraneous noise.
+While all methods compromise fidelity and diversity, as reflected in the FDD and Vendi Score, TIPO remains competitive because it maintains small semantic deviation from the original sentences via progressive refinement.
+(a) Oringial Prompt
+(b) Text Completion
+(c) TIPO
+Figure 6: Qualitative comparison of prompt completion versus TIPO (ours).
+(a) shows images generated by original prompts.
+(b) depicts images generated by text completion using TIPO-500M.
+(c) shows images generated by TIPO.
+5.4 Out-of-domain Performance
+Some recent T2I models are trained on proprietary images and captions.
+As a representative example, SD-3.5-Large[20] is trained on private images captioned with CogView[71], which differ markedly from the texts used to train TIPO.
+To evaluate model performance in this out-of-domain scenario, we generate 8,192 original tag- and NL-based prompts using the baseline GPT-4o-mini rather than relying on existing prompt datasets.
+We apply the remaining methods to these prompts and assess their performance.
+GPT MagicPrompt Promptist TIPO(Ours)
+Aesthetic Score $\uparrow$ 6.7125 6.4507 6.3924 6.0536
+AI Corrupt Score $\uparrow$ 0.9482 0.8577 0.9053 0.9280
+Vendi Score $\uparrow$ 8.9718 15.8721 16.4891 21.5706
+Table 3: Performance on the out-of-domain T2I model
+As shown in Table 3, SD-3.5-Large faithfully generates images that align well with GPT-produced prompts.
+Consequently, additional optimizations tend to reduce fidelity and introduce more artifacts.
+Nevertheless, GPT-generated prompts are accurate and lack diversity.
+TIPO optimization enriches the prompts with additional details that harmonize with the original themes, significantly enhancing the diversity of the generated images.
+5.5 Human Preference Evaluation
+Quantitative metrics may not fully align with human preference.
+Therefore, we conducted a user study based on pairwise image comparisons between the original prompt, MagicPrompt, Promptist, and TIPO on over 1,400 images, gathering preferences from 221 volunteers.
+As illustrated in Figure 7, TIPO achieved the highest overall win rate at 51.3%, significantly outperforming competitors.
+In out-of-domain scenarios, TIPO’s win rate increased to 52.5%, demonstrating consistently strong user preference across different contexts.
+For further results and statistics, please refer to Appendix F.
+Figure 7: The win-tie-lose rate on overall preference on all data and out-of-domain situation where images are generated by SD-3.5-Medium.
+Figure 8: ELO rating on overall preference on all data
+5.6 Ablation Study
+Original Completion TIPO
+Aesthetic Score $\uparrow$ 5.9507 5.9631 5.9960
+Corrupt Score $\uparrow$ 0.6799 0.8131 0.8971
+Vendi Score $\uparrow$ 19.7208 18.2896 13.9864
+Table 4: Ablation results
+We attribute TIPO’s success primarily to its multi-task pre-sampling design detailed in Section 4.3.
+However, a natural question arises: can a simple text completion approach achieve similar or better results?
+To validate our design, we compare TIPO’s pre-sampling strategy with a direct text completion using the same language model, TIPO-500M.
+As shown in Table 4, TIPO significantly improves aesthetic scores and reduces artifacts compared to text completion.
+For visual examples, please refer to Figure 6.
+For more ablation details and visualization, please refer to Appendix 5.6.
 6 Conclusion
-In this paper, we introduced TIPO (Text to Image with text Presampling for Prompt Optimization), a novel framework designed to enhance the quality of Text-to-Image (T2I) models through automatic prompt engineering.
-TIPO represents a significant advancement in the field of T2I generation, offering a versatile and efficient solution for prompt optimization.
-By automating the refinement and extension of user prompts.
-We also highlights the importance of aligning user inputs with underlying data distributions through the comparisons between different methods for prompt engineering.
-As generative models continue to evolve, frameworks like TIPO will be instrumental in unlocking their full potential, ensuring they meet diverse user needs with precision and creativity while pushing the boundaries of what’s possible in generative models.
+We introduced TIPO, a lightweight prompt pre-sampling framework designed for efficient real-world Text-to-Image (T2I) applications.
+By aligning user prompts with the intrinsic distributions of T2I training datasets, TIPO enhances semantic coherence, image fidelity, and diversity with minimal inference overhead.
+TIPO consistently outperforms existing prompt optimization methods, achieving superior performance across multiple evaluation metrics.
+Its ability to preserve user intent while enhancing prompt diversity and image quality makes it well-suited for both in-domain and out-of-domain scenarios, demonstrating robustness across various prompt formats and model architectures.
+Extensive user studies also confirmed TIPO’s strong alignment with human preferences, further highlighting its potential for practical applications.
+To encourage wider adoption and facilitate reproducibility, we release our trained models and source code.
+We hope TIPO will inspire further advancements in efficient, scalable, and robust generative frameworks for creative systems.
 Appendix
+[…]
+[…]
+[…]
+[…]
 […]
@@
 The Danbooru2023 dataset [69, 68, 46] is an extensive collection of images and their corresponding tags, compiled from the Danbooru image board.
-This dataset includes images annotated with highly specific and detailed tags, providing a rich resource for training both the Text-to-Image (T2I) and Large Language Models (LLMs) involved in the TIPO framework.
-The dataset contains data up to image ID 7,349,999, encompassing a wide variety of visual content with granular annotations.
-These annotations allow for the creation of nuanced and precise prompts, ensuring that longer, more detailed prompts can indicate subsets of shorter prompts.
+This dataset includes images annotated with particular and detailed tags, providing a rich resource for training both the Text-to-Image (T2I) and Large Language Models (LLMs) involved in the TIPO framework.
+The dataset contains data up to image ID 7,349,999, encompassing various visual content with granular annotations.
+These annotations allow for creating nuanced and precise prompts, ensuring that longer, more detailed prompts can indicate subsets of shorter prompts.
 Key Characteristics:
-  • Rich Annotations: Danbooru2023’s detailed tags enable the differentiation of subtle variations in image content, such as “long hair” versus “very long hair,” which is crucial for the specificity required in high-quality image generation.
-  • Large Volume: The dataset’s extensive size provides a robust foundation for training models, ensuring that they can learn from a diverse array of images and annotations.
-  • Tag-Based Prompting: Utilizing the detailed tags from Danbooru2023, the LLM can generate refined prompts that lead to more accurate and high-quality image generation by the T2I model.
+  • Rich Annotations: Detailed tags differentiate subtle variations, crucial for specific image generation.
+  • Large Volume: Extensive dataset size ensures diverse training examples.
+  • Tag-Based Prompting: Refined prompts from detailed tags enhance image generation accuracy.
 A.2 GBC10M
 The GBC10M dataset [29] is a large-scale collection of 10 million images sourced from CC12M [13], annotated using the Graph-Based Captioning (GBC) approach.
-Each image in GBC10M is represented by a graph where nodes correspond to object regions, compositions, and relations, and edges define the hierarchical relationships among them.
-Annotations are generated automatically through a pipeline that leverages pretrained multimodal large language models (MLLM) and object detection tools.
-The GBC structure enhances traditional image captions by providing both detailed descriptions and structural information, improving model performance in downstream tasks.
-All data is provided in JSON lines format, containing image URLs, bounding boxes, and caption annotations.
-In TIPO, we only utilize the root node from GBC10M, since it provides both detail and short captions.
+Each image is represented by a graph where nodes correspond to object regions, compositions, and relations, and edges define their hierarchical relationships.
+Annotations are generated automatically through a pipeline leveraging pretrained multimodal large language models (MLLM) and object detection tools.
+The GBC structure enhances traditional image captions by providing detailed descriptions and structural information.
+Data is provided in JSON lines format, including image URLs, bounding boxes, and captions.
+In TIPO, only the root node captions from GBC10M are utilized for concise yet descriptive prompts.
 A.3 Coyo HD 11M
-The Coyo HD 11M dataset [12] consists of 11.4 million high-density, high-definition images, with 22.8 million synthetic captions generated from the Coyo-700M dataset.
-The “HD” refers to both the high resolution of images (filtered to have at least 512 pixels on the shortest edge) and the high concept density, addressing the issues found in previous alt-text image pair datasets, which often included low-quality or low-concept images.
-Synthetic captions for this dataset were produced using the LLaVA-Next-8B model [37] built upon LLaMA 3 [3] with post-processing for conciseness and clarity.
-This dataset provides high-quality visual and textual pairs, ideal for training advanced vision-language models.
-In TIPO, we utilize the short/long captions, booru tags, and open images tags from Coyo HD 11M.
+The Coyo HD 11M dataset [12] consists of 11.4 million high-resolution, high-concept-density images paired with 22.8 million synthetic captions generated from the Coyo-700M dataset.
+Images maintain a minimum of 512 pixels on the shortest edge to ensure high visual quality.
+Captions, generated with the LLaVA-Next-8B model [38] based on LLaMA 3 [3], undergo post-processing for conciseness and clarity.
+TIPO uses short and long captions, booru tags, and open image tags from this dataset.
 A.4 Stable Diffusion XL
-The Stable Diffusion XL [48] improves upon earlier versions [54] with a significantly larger UNet backbone and dual text encoders, CLIP ViT-L [49] and OpenCLIP ViT-bigG [30], to enhance text conditioning.
-It supports resolutions up to 1024×1024 and is designed for diverse image generation.
-Although explicit dataset details are not provided, figures in the paper confirm that SDXL accepts both natural language (NL) prompts and tags as input.
-A.5 Stable Diffusion 3.5
-The Stable Diffusion 3.5 (SD-3.5) incorporates a cutting-edge architecture, MMDiT [21], along with the Rectified Flow [38, 4, 35] formulation for improved text-to-image generation.
-The model utilizes triple text encoders—CLIP/ViT-L, OpenCLIP/ViT-G, and T5-XXL [50]—for robust text conditioning.
-SD-3.5 supports resolutions up to 1024×1024 and employs a 50/50 mix of original and CogVLM-generated captions for training.
-While explicit dataset details are not provided, figures in the paper confirm its capability to process both natural language prompts and tags.
-In this study, we use the SD-3.5-Large [^8] https://huggingface.co/stabilityai/stable-diffusion-3.5-large (8B parameters) with FP8 inference on RTX 3090 or RTX 4090 GPUs.
+Stable Diffusion XL (SDXL) [50] improves upon earlier models [56] with a more considerable UNet backbone and dual text encoders (CLIP ViT-L [51] and OpenCLIP ViT-bigG [31]), enhancing text conditioning.
+Supporting resolutions up to 1024×1024, SDXL accepts natural language prompts and tags, suitable for diverse image generation.
+A.5 Illustrious
+Illustrious is a series of fine-tuned Stable Diffusion XL models primarily trained on the Danbooru2023 dataset.
+In this study, we specifically employ the v3.5 version variant with v-parameterization [58], which is notable for its extensive incorporation of natural language prompts.
+The inclusion of both tag-based and natural language formats allows Illustrious to leverage a broad range of semantic knowledge for image generation.
+Within TIPO, we perform an ablation study to analyze the effectiveness of different prompting strategies—namely extended tags versus natural language prompts—to identify which approach contributes most significantly to enhanced image generation performance.
+A.6 Stable Diffusion 3.5
+Stable Diffusion 3.5 (SD-3.5) incorporates the MMDiT architecture [20] and the Rectified Flow formulation [39, 4, 36] for improved text-to-image generation.
+Utilizing triple text encoders (CLIP/ViT-L, OpenCLIP/ViT-G, T5-XXL [52]), SD-3.5 supports resolutions up to 1024×1024 and uses a 50/50 mix of original and CogVLM-generated captions.
+Figures confirm the capability to process both natural language prompts and tags.
+This study employs SD-3.5-Large [^6] https://huggingface.co/stabilityai/stable-diffusion-3.5-large (8B parameters) with FP8 inference on RTX 3090 or RTX 4090 GPUs.
 Appendix B Implementation Details
-In this appendix, we provide all the necessary details including our dataset construction process, model configurations, inference pipeline and the model’s property which are not mentioned in Section 4.3 and 4.4.
+In this appendix, we provide all the necessary details including our dataset construction process, model configurations, inference pipeline and the model’s properties not mentioned in Section 4.2 and 4.3.
 B.1 Training Data Construction
@@
 Length Control
-To systematically control output prompt length, we implement a structured length categorization system using special length tags.
-These tags enforce specific constraints on both tag counts and natural language sentence lengths.
+To systematically control output prompt length, we implement a structured length categorization system using unique length tags.
+These tags enforce specific constraints on tag counts and natural language sentence lengths.
 For instance, the <long> tag specifies that the corresponding prompt must contain between 36 and 52 tags (inclusive), accompanied by 4 to 8 sentences of natural language description.
-We define four distinct length categories, each with strict bounds for both tag count and sentence length.
+We define four distinct length categories, each with strict bounds for tag count and sentence length.
 Type Very Short Short Long Very Long
@@
     • Random repositioning of metadata tags to the end of the prompt, after all content-related descriptions
-    This approach encourages the model to handle varying metadata positions and availability, while maintaining the ability to infer metadata relationships from the content descriptions.
+    This approach encourages the model to handle varying metadata positions and availability, while maintaining the ability to infer metadata relationships from content descriptions.
   • Content Tags: For tags describing image content (e.g., objects, actions, attributes), we implement:
@@
     • Length-based truncation to meet target length constraints while preserving key content information
-  • Natural Language: For natural language descriptions exceeding length limitations, we employ selective sentence removal, targeting middle sentences to preserve both context-setting opening sentences and concluding details.
+  • Natural Language: For natural language descriptions exceeding length limitations, we employ selective sentence removal, targeting middle sentences to preserve context-setting opening sentences and concluding details.
 This maintains coherent narrative flow while meeting target length requirements.
@@
 Token Seen 6.0240B 22.625B 18.339B 31.274B
-Hardware 4 x RTX3090 4 x RTX3090 4 x RTX3090 8 x H100
+Hardware 4 × RTX3090 4 × RTX3090 4 × RTX3090 8 × H100
 Training Time (wall) 22.5 hour 150 hour 270 hour 100 hour
-Table 5: Training settings and details for TIPO-100M, 200M, and 500M.
+Table 6: Training settings for TIPO models.
 The datasets include CoyoHD11M (Coyo), GBC10M (GBC), and Danbooru2023 (Dan).
@@
 Table 6 summarizes the configurations for pretraining and fine-tuning TIPO-100M, TIPO-200M, and TIPO-500M.
-Both pretraining and finetuning was conducted on datasets like Danbooru2023 [45], GBC10M [29], and CoyoHD11M [12].
+Both pretraining and fine-tuning was conducted on datasets like Danbooru2023 [46], GBC10M [29], and CoyoHD11M [12].
 Augmented Task Representation
@@
 B.3 Inference Pipeline
-The TIPO inference pipeline is designed to handle various input types and scenarios, combining different tasks to achieve refinement or expansion of both tag-based prompts and natural language prompts.
+The TIPO inference pipeline is designed to handle various input types and scenarios, combining different tasks to refine or expand both tag-based and natural language prompts.
 Figure 9 illustrates this comprehensive workflow.
@@
 This flexible pipeline allows TIPO to adapt to various input scenarios, whether the user provides tags, natural language descriptions, or both.
-By leveraging different task combinations, TIPO ensures that both tag-based and natural language prompts are optimized, resulting in more detailed and effective inputs for text-to-image models.
+By leveraging different task combinations, TIPO ensures that tag-based and natural language prompts are optimized, resulting in more detailed and effective input for text-to-image models.
 B.4 Model Speed Comparison
@@
 Table 7: Model performance comparison across different hardware platforms.
-Tokens per second (tok/sec) represents the average generation speed, while generation time (gen time) shows the average time required for a complete two-step prompt optimization process.
+Tokens per second (tok/sec) represents the average generation speed, while generation time (gen time) shows the average time in seconds required for a complete two-step prompt optimization process.
 Appendix C Evaluation Statistics
 In this appendix, we provide more statistics for the result obtained in Section 5.
-C.1 In-domain test regard to scenery tag
+C.1 In-domain test regarding scenery tag
 (a) The box plot for the Aesthetic Score result of scenery tag test.
@@
 Figure 10: The distribution of aesthetic and AI corrupt score for scenery tag test.
-The box plot and Kernel Density Estimation (KDE) plot shown in Figure 9 presents the aesthetic scores and AI corruption scores for results from the scenery tag test described in Section 5.2.
-The analysis clearly demonstrates that TIPO significantly outperforms all other methods, with a substantial margin of improvement.
+The box plot and Kernel Density Estimation (KDE) plot displayed in Figure 10 illustrate the aesthetic scores and AI corruption scores from the scenery tag test described in Section 5.2.
+The analysis shows that TIPO significantly outperforms all other methods, demonstrating a considerable margin of improvement.
 C.2 In-domain prompt generation test
@@
 Figure 12: The distribution of aesthetic and AI corrupt score for truncated long prompt input.
-Figures 10 and 11 display the box plots and KDE plots of aesthetic scores and AI corruption scores obtained from the In-domain prompt generation test detailed in Section E.2.
-While the box plots reveal that the differences in performance between various methods are subtle, the AI corruption scores provide valuable insights.
-Specifically, these scores indicate that implementations supported by TIPO produce more stable output images compared to other methods.
-C.3 Out of focus evaluation
+Figures 11 and 12 display the box plots and KDE plots of aesthetic scores and AI corruption scores obtained from the In-domain prompt generation test detailed in Section E.2.
+While the box plots reveal subtle differences in performance between various methods, the AI corruption scores provide valuable insights.
+Specifically, these scores indicate that implementations supported by TIPO produce more stable output images than other methods.
+C.3 Out-of-domain evaluation
+(a) The box plot for the Aesthetic Score result of out-of-focus test.
+(b) The box plot for the AI Corrupt Score result of out-of-focus test.
+(c) The KDE plot for the Aesthetic Score result of out-of-focus test.
+(d) The KDE plot for the AI Corrupt Score result of out-of-focus test.
+Figure 13: The distribution of aesthetic and AI corrupt score for out-of-focus test.
 (a) Original Caption
 (b) Prompt by GPT4o-mini
-(c) Prompt by Prompt-DB
+(c) Prompt by MagicPrompt
 (d) Prompt by Promtist
@@
 (b) Prompt by GPT4o-mini
-(c) Prompt by Prompt-DB
+(c) Prompt by MagicPrompt
 (d) Prompt by Promtist
@@
 Figure 15: The similarity matrix between 100 images of worst aesthetic generated results of SD3.5-Large experiments.
-Figures 12 and 13 present similarity matrices for different prompt generation methods and their corresponding aesthetic outputs on SD3.5-Large  [22].
-A matrix with predominantly lower similarity values (overall brighter appearance) indicates high diversity across all generated images, while higher similarity values (overall darker appearance) suggest consistent but less diverse outputs.
-For quantitative diversity metrics, please refer to Table 3 in Section 5.
+Figures 14 and 15 present similarity matrices for different prompt generation methods and their corresponding aesthetic outputs on SD3.5-Large  [21].
+A matrix with predominantly lower similarity values (brighter appearance) indicates high diversity among generated images, while higher values (darker appearance) suggest consistent but less diverse outputs.
+Please refer to Table 3 in Section 5.
+C.4 Ablation Test
+(a) The box plot for the Aesthetic Score result of NL ablation test.
+(b) The box plot for the AI Corrupt Score result of NL ablation test.
+(c) The KDE plot for the Aesthetic Score result of NL ablation test.
+(d) The KDE plot for the AI Corrupt Score result of NL ablation test.
+Figure 16: The distribution of aesthetic and AI corrupt score for NL ablation test.
+(a) The box plot for the Aesthetic Score result of tags ablation test.
+(b) The box plot for the AI Corrupt Score result of tags ablation test.
+(c) The KDE plot for the Aesthetic Score result of tags ablation test.
+(d) The KDE plot for the AI Corrupt Score result of tags ablation test.
+Figure 17: The distribution of aesthetic and AI corrupt score for tags ablation test.
+Figures 17 present the tag ablation test in the TIPO effect on the aesthetic score and AI Corrupt Score among the original tag, tag-extend and the tags TIPO.
+The box plot reveals that the tag TIPO is better than the original tag and the tag extend is the best.
+In detail, KDE plot reveals that the tag TIPO has a similar performance compared with the tag extend.
+Both of them are better than the original tag, which indicates that the tag TIPO aspect helps control corruption and promotes the aesthetic score.
 Appendix D TIPO example
-In this appendix, we provide some text example of TIPO’s input and output.
-TIPO Format template User Input:
-1girl, ciloranko, maccha (mochancc), ningen mame, ask (askzy), solo, masterpiece, absurdres, newest, safe A girl sits in a cozy cafe, cradling a cup of coffee in her hand
-Formatted TIPO Input For Expand Tags:
-meta: absurdres rating: safe artist: ciloranko, maccha (mochancc), ningen mame, ask (askzy) quality: masterpiece, newest aspect_ratio: 1.0 target: <|short|> <|short_to_tag|> short: A girl sits in a cozy cafe, cradling a cup of coffee in her hand tag: 1girl, solo
-Formatted TIPO Output after Expand Tags and Expand Natural Prompt:
-meta: absurdres rating: safe artist: ciloranko, maccha (mochancc), ningen mame, ask (askzy) quality: masterpiece, newest aspect_ratio: 1.0 target: <|short|> <|tag_to_long|> tag: 1girl, solo, sitting, closed mouth, jewelry, long hair, looking at viewer, crossed legs, plant, table, couch, bracelet, cup, smile, teacup, indoors, blue eyes, blonde hair, holding long: A girl sits in a cozy cafe, cradling a cup of coffee in her hand.
+In this section, we provide some text example of TIPO’s input and output.
+TIPO Format Template User Input: 1girl, ciloranko, maccha (mochancc), ningen mame, ask (askzy), solo, masterpiece, absurdres, newest, safe A girl sits in a cozy cafe, cradling a cup of coffee in her hand Formatted TIPO Input for Expand Tags: meta: absurdres rating: safe artist: ciloranko, maccha (mochancc), ningen mame, ask (askzy) quality: masterpiece, newest aspect_ratio: 1.0 target: <|short|> <|short_to_tag|> short: A girl sits in a cozy cafe, cradling a cup of coffee in her hand tag: 1girl, solo Formatted TIPO Output after Expand Tags and Expand Natural Prompt: meta: absurdres rating: safe artist: ciloranko, maccha (mochancc), ningen mame, ask (askzy) quality: masterpiece, newest aspect_ratio: 1.0 target: <|short|> <|tag_to_long|> tag: 1girl, solo, sitting, closed mouth, jewelry, long hair, looking at viewer, crossed legs, plant, table, couch, bracelet, cup, smile, teacup, indoors, blue eyes, blonde hair, holding long: A girl sits in a cozy cafe, cradling a cup of coffee in her hand.
 The cafe has large windows with green plants on the walls and a wooden table in front of her.
@@
 The overall atmosphere of the image is serene and inviting.
-Formatted Output For Text-to-Image
-1girl, ciloranko, maccha \(mochancc\), ningen mame, ask \(askzy\), solo, sitting, closed mouth, jewelry, long hair, looking at viewer, crossed legs, plant, table, couch, bracelet, cup, smile, teacup, indoors, blue eyes, blonde hair, holding, A girl sits in a cozy cafe, cradling a cup of coffee in her hand.
+Formatted Output for Text-to-Image: 1girl, ciloranko, maccha (mochancc), ningen mame, ask (askzy), solo, sitting, closed mouth, jewelry, long hair, looking at viewer, crossed legs, plant, table, couch, bracelet, cup, smile, teacup, indoors, blue eyes, blonde hair, holding, A girl sits in a cozy cafe, cradling a cup of coffee in her hand.
 Behind her is a staircase leading to another room.
@@
 The overall atmosphere of the image is serene and inviting.
-masterpiece, newest, absurdres, safe
-Refer to caption User Input Refer to caption TIPO Output
-Figure 14: An example formatted content we used for training and inference in TIPO.
+masterpiece, newest, absurdres, safe Refer to caption User Input Refer to caption TIPO Output
+Figure 18: An example of formatted content used for training and inference in TIPO.
 TIPO Format template User Input:
@@
 However, the use of TIPO to refine or extend these prompts still yields noticeable improvements in aesthetics and overall quality.
-Appendix F Topic Distribution Visualization
+Appendix F Human Preference
+(a) ELO rating on Illustrious
+(b) Win rate matrix on Illustrious
+(c) ELO rating on SD3.5-medium
+(d) Win rate matrix on SD3.5-medium
+Figure 22: ELO ratings and win rate matrices across different experimental settings comparing five prompting methods (TIPO, Promptist, Promptextend, MagicPrompt, and Original) on three evaluation dimensions
+(a) Full Win-Tie-Lose plot
+(b) Illustrious Win-Tie-Lose plot
+(c) SD3.5-medium Win-Tie-Lose
+Figure 23: Win-Tie-Lose comparison across different experimental settings showing the relative performance of five prompting methods on prompt adherence, image quality, and aesthetic appeal metrics
+(a) The UI of survey system before submitting the choices.
+(b) The UI of survey system after submitting the choices.
+Figure 24: Survey interface for human evaluation of image pairs, showing the evaluation process before submission (a) where users compare two images based on four metrics, and after submission (b) where the generated prompts for each image are revealed
+Comparison Win Ratio (A:B) Proportion A $p$-value Significant?
+Original vs. PromptExtend 45:66 0.405 0.0572 Marginally
+Original vs. Tipo 28:80 0.259 $<0.0001$ Yes***
+Original vs. Promptist 35:64 0.354 0.0046 Yes**
+PromptExtend vs. Promptist 41:80 0.339 0.0005 Yes***
+Promptist vs. Tipo 30:101 0.229 $<0.0001$ Yes***
+PromptExtend vs. Tipo 30:90 0.250 $<0.0001$ Yes***
+MagicPrompt vs. Promptist 14:34 0.292 0.0055 Yes**
+MagicPrompt vs. Tipo 11:48 0.186 $<0.0001$ Yes***
+MagicPrompt vs. Original 15:28 0.349 0.0660 No
+MagicPrompt vs. PromptExtend 19:35 0.352 0.0402 Yes*
+Table 8: Pairwise win rates and statistical significance (Overall Dimension). Significance levels: * $p<0.05$, ** $p<0.01$, *** $p<0.001$
+We conducted a series of A/B tests to compare five prompt transformations, (TIPO, Promptist, Promptext, MagicPrompt, and Original(unmodified)), for two models, Illustrious, SD3.5-medium, which is known for both core word/natural language understanding.
+In total, we collected responses for $\sim$1,500 pairwise comparisons, from more than 20 anonymous evaluators.
+Each evaluation asked participants to compare two generated images—labeled A and B—and select which they preferred (or a tie) according to specific criteria (e.g., prompt adherence, image quality, or aesthetic appeal).
+F.1 User Interface for Human Preference Evaluation
+We developed a specialized survey interface to facilitate efficient and unbiased human evaluation of generated images.
+As illustrated in Figure 24, the interface presents evaluators with an original prompt and two corresponding images (labeled A and B) generated using different prompting methods.
+Before submission, users can see the original prompt in the center panel while the processed prompts used to generate each image remain hidden to prevent bias.
+The evaluation framework requires participants to compare the image pairs across four distinct metrics: prompt adherence (how well the image follows the original prompt), image quality (detail and correctness), aesthetic appeal (color, composition, and style), and overall personal preference.
+For each metric, users can select one of three options: “A is better,” “A and B are equal,” or “B is better.”
+When evaluators encounter image pairs that appear to be from different prompts or settings, they are instructed to click “Refresh” to obtain a new comparison.
+After submitting their evaluations, the interface reveals the transformed prompts used to generate each image, providing transparency about how the original prompt was modified by each method.
+F.2 Extended Human Evaluation.
+Participants assessed each image’s performance on prompt adherence, image quality, and aesthetic appeal, with visually shown unmodified and image pairs.
+TIPO exhibited superior outcomes in all comparison settings.
+Notably, it attained a 64.4% peak win rate (against MagicPrompt) under the Full scenario and 57.5% (also against MagicPrompt) under SD35-medium, emphasizing TIPO’s proficiency in generating images that closely follow prompt specifications while maintaining visual coherence.
+F.3 ELO Ratings.
+We computed theoretical ELO ratings from the aggregated pairwise comparisons to quantify overall performance differences among the five methods.
+The rating update rules were based on each pair’s binary outcome (win or lose), ignoring tie cases.
+The result is depicted in Figure 22, TIPO has secured the highest ELO rating over other models.
+F.4 Human Preference ELO Method
+We computed theoretical ELO ratings from human-judged pairwise preference data to quantitatively evaluate the relative performance of each prompting method.
+The ELO rating system, initially designed for ranking chess players, aggregates binary outcomes into numerical ratings representing comparative performance.
+Pairwise Outcomes.
+Human evaluators assessed comparisons between methods, resulting in one of three outcomes:
+  • Method $i$ wins: assigned a score of $1$ for method $i$, and 00 for method $j$.
+  • Method $j$ wins: assigned score $1$ for method $j$, and 00 for method $i$.
+  • Tie: assigned score $0.5$ to both methods.
+Conversion to ELO Differences.
+Win and tie rates were converted to ELO rating differences using:
+Adjusted Win Rate=Win Rate+Tie $\text{Adjusted Win Rate}=\text{Win Rate}+\frac{\text{Tie Rate}}{2}$
+ELO $\text{ELO Difference}=400\times\log_{10}\left(\frac{\text{Adjusted Win Rate}}{1-\text{Adjusted Win Rate}}\right)$
+To ensure numerical stability, extreme adjusted win rates were constrained as follows:
+ELO $\text{ELO Difference}=\begin{cases}-800,&\text{Adjusted Win Rate}\leq 0.001\\ +800,&\text{Adjusted Win Rate}\geq 0.999\\ \end{cases}$
+Calculating Method ELO Ratings.
+Final ELO ratings were determined by averaging each method’s pairwise ELO differences and centering these averages around a baseline rating (e.g., $1000$):
+$\text{ELO}_{\text{method}_{i}}=\text{Base Rating}+\left(\text{Average ELO Difference for method }i-\text{Overall Mean ELO Difference}\right)$
+Interpretation of ELO Scores.
+Methods with higher ELO scores consistently outperform lower-scored methods. A rating difference of 400 points corresponds to a 90% expected win probability for the superior method.
+F.5 Statistical Significance
+As summarized in Table 8, we conducted two-sided binomial and McNemar’s tests (p <0.05) to assess the statistical significance of observed differences.
+The result confirms that TIPO’s advantages are unlikely to be explained by random variation, which also supports a consistent performance hierarchy: TIPO ranks highest, followed by Promptist, PromptExtend, Original, and MagicPrompt.
+Collectively, these findings illustrate TIPO’s robust, model-agnostic effectiveness and underscore the model-sensitivity of alternative methods, particularly Promptext and MagicPrompt.
+F.6 Survey Response Examples
+In this section we provided some responses of our human preference survey as reference.
+Figure 25: Some survey responses on illustrious-3.5-vpred generated image with different prompt optimization method
+Figure 26: Some survey responses on illustrious-3.5-vpred generated image with different prompt optimization method
+Figure 27: Some survey responses on SD3.5-medium generated image with different prompt optimization method
+Figure 28: Some survey responses on SD3.5-medium generated image with different prompt optimization method
+F.7 Conclusion
+The extended evaluations presented here reinforce TIPO’s standing as a reliable and effective prompt-optimization strategy.
+Its consistent performance gains under diverse model conditions highlight its potential for broad applicability, with its strong alignment with user-specified prompts, high image quality, and favorable aesthetic outcomes.
+Appendix G Ablation Study on TIPO
+In this section, we investigate the effect of incorporating TIPO (Tags + Inferred Prompt Objects) across various generation settings.
+Our primary goal is to validate whether additional structured information (e.g., core tags and minimal spatial/contextual cues) can improve image quality, reduce artifacts.
+G.1 Experimental Setup
+Prompt Variants.
+To systematically analyze TIPO’s contribution, we consider four types of input prompts improvement task:
+  1. 1.
+    Tag $\to$ More core words: Given an initial set of core words, generate more refined or expanded core words.
+  2. 2.
+    NL $\to$ More NL: Given a short natural language (NL) description, elaborate into a richer NL prompt.
+  3. 3.
+    Tag $\to$ (More core words + NL): Combine expanded tags with a corresponding NL description derived from them.
+  4. 4.
+    NL $\to$ (More NL + core words): Use the NL prompt to add relevant tags, forming a mixed prompt of NL plus core words.
+In each case, we compare the baseline prompts (without TIPO cues) against prompts incorporating TIPO’s structured, tag-based critical information and minimal spatial hints.
+Data Preparation.
+We start by randomly sampling core words from a word table to represent a diverse range of topics (e.g., objects, environments, descriptors).
+Additionally, for each word set, we generate a corresponding short NL sentence using a compact language model (GPT4o-mini).
+Overall, the six prompt variants are tested on 4,000 images, ensuring a balanced comparison.
+Inference Procedure.
+Prompts are fed into our image-generation pipeline under identical model settings (classifier free guidance, sampler, steps, etc.), using the v-parameterized variant of Illustrious v3.5[49].
+We focus on how TIPO modifications alter the generation outcomes and whether they introduce additional computational overhead.
+G.2 Evaluation Metrics
+(a) Corrupt Score Distributions
+(b) Aesthetic Score Distributions
+Figure 29: Side-by-side comparison of Corrupt (left) and Aesthetic (right) score distributions across prompt types.
+Aesthetic Score.
+We employ an off-the-shelf aesthetic predictor to estimate image quality. In the following paragraph, we discuss the model’s bias.
+AI Corruption Score.
+Using an automated ’ AI corruption ’ detection model, we measure generation artifacts, such as distorted objects and unnatural shapes. Higher scores imply cleaner, more coherent outputs.
+G.3 Results & Discussion
+Impact on Aesthetics.
+Figure 29(b) shows that TIPO-enhanced prompts generally achieve higher aesthetic scores than their non-TIPO counterparts, albeit with some variance.
+Notably, we observe a correlation between wider color ranges and higher aesthetic scores discussed in Figure 30, suggesting a bias toward more colorful or varied compositions.
+Improvements in Corruption Score.
+As shown in Figure 29(a), TIPO-based prompts yield significantly lower corruption scores, indicating fewer artifacts.
+We hypothesize that the additional spatial and contextual details encoded via TIPO help the model place objects more consistently.
+(a) Saturation vs. Aesthetic Score
+(b) Binned Saturation vs. Aesthetic
+(c) Saturation vs. Corrupt Score
+Figure 30: Scatter plots (left and right) and binned analysis (middle) showing the relationship between saturation and image metrics.
+We find a moderate positive correlation between saturation and aesthetic score (Pearson $r=0.2821$), particularly at lower saturation ranges, based on 24k samples.
+However, saturation shows no notable correlation with corrupt score (Pearson $r=0.0125$).
+G.4 Speed Test and Overhead Analysis
+A key concern for production pipelines is whether TIPO generation imposes a substantial time overhead.
+We benchmarked prompt-generation inference on four smaller models, excluding any large proprietary LLMs.
+As illustrated in Table 9, the additional TIPO-related computation remains well below the image-generation time.
+Hence, even in a synchronous pipeline, TIPO prompt expansion does not constitute a bottleneck.
+Table 9: Speed Test Results for TIPO and Other Prompt Methods
+Method Model/Config # Runs Avg. Time (s) Std. Dev. (s)
+TIPO LLaMA-500M 500 1.4207 1.0730
+LLaMA-200M 200 1.0306 0.8982
+LLaMA-100M 200 1.0078 0.9394
+PROMPTIST GPT2-125M 1000 1.4593 0.2857
+PROMPTEXTEND GPT2-125M 1000 1.3849 0.2151
+MAGICPROMPT GPT2-125M 1000 1.1398 0.4043
+Memory Footprint.
+We also confirm that TIPO’s overhead in terms of VRAM usage is minimal. The practical adoption of TIPO in pipelines has shown no critical memory concerns, which aligns with our measurements.
+G.5 Conclusion of Ablation
+Our experiments suggest that TIPO (1) consistently lowers AI corruption artifacts, (2) can boost aesthetic scores, and (3) remains computationally inexpensive.
+The improvements in metrics support the viability of TIPO prompts for real-world image-generation tasks.
+In short, a concise natural language prompt with core tag-based critical information appears to be an effective, suggested form for most use cases.
+Appendix H Topic Distribution Visualization
 Latent Dirichlet Allocation (LDA) [9] is a generative probabilistic model for topic modeling [32], which assumes that each document is a mixture of topics, with each topic represented by a distribution over words.
@@
 This approach focuses on identifying significant representative words while filtering out stop words and irrelevant terms to ensure meaningful topic classification.
-By employing multi-level topic analysis, we empirically assessed whether the resulting topics were sufficiently large and diverse.
+We empirically assessed whether the resulting topics were sufficiently large and diverse by employing multi-level topic analysis.
 This iterative process mitigates potential challenges such as substantial topic overlap, which can diminish distinctiveness when extracting a large number of topics [62].
 To address the potential overlap and further assess the diversity and meaningfulness of the topics, we performed a secondary clustering [70].
-We grouped the initially extracted topics into five major clusters using k-means clustering and evaluated the clustering performance by calculating the inertia (Sum of Squared Distances) [27], shown in Table 7, 8, and 9.
+We grouped the initially extracted topics into five major clusters using k-means clustering.
+We evaluated the clustering performance by calculating the inertia (Sum of Squared Distances) [27], shown in Table 10, 11, and 12.
 Since the topics have already been filtered for meaningful content, a higher inertia value indicates greater diversity among the clusters, reflecting a broader range of valid and meaningful topics across the dataset.
 This two-tiered approach allows for a more nuanced analysis of topic diversity and ensures the robustness of the topic modeling against meaningless word groupings.
-Size Promptdb GPT4o-mini Promptist TIPO
+Size MagicPrompt GPT4o-mini Promptist TIPO
 Run 1 Run 2 Run 1 Run 2 Run 1 Run 2 Run 1 Run 2
@@
 Table 10: Inertia for COYO-Dataset inference, higher is better
-Size Promptdb GPT4o-mini Promptist TIPO
+Size MagicPrompt GPT4o-mini Promptist TIPO
 Run 1 Run 2 Run 1 Run 2 Run 1 Run 2 Run 1 Run 2
@@
 Table 11: Inertia for GBC-Dataset inference, higher is better
-Size PromptDB GPT4o-mini Promptist TIPO
+Size MagicPrompt GPT4o-mini Promptist TIPO
 20 60.82 734.52 210.60 139.29
@@
 Table 12: Inertia for Scenery extend inference, higher is better
-We attach simple visualization of topics in scenery prompt generation, with topic n=100, cluster k=5.
+We attach a simple visualization of topics in scenery prompt generation, with topic n=100, cluster k=5.
 Figure 31: Topic visualization for scenery prompt generation. A wider spread indicates a greater diversity of generated topics.
 Figure 32: This visualization represents a filtered subset of posts from the Danbooru2023 dataset, centered on the ’scenery’ tag.
-The network graph is an ego network (depth 1), meaning it includes only nodes directly connected to the ’scenery’ tag.
+The network graph is an ego network (depth 1), which includes only nodes directly connected to the ’scenery’ tag.
 To refine the data and focus on meaningful associations, uncommon tags with fewer than 10 occurrences were excluded.
@@
 Nodes are color-coded by modularity class by Fast Unfolding Algorithm [10], revealing clusters of closely associated tags.
-Node size reflects Eigenvector Centrality[11], emphasizing tags that are both highly connected and influential within their network.
+Node size reflects Eigenvector Centrality[11], emphasizing highly connected and influential tags within their network.
+Appendix I Discussion and Future Work
+Despite the promising performance demonstrated by TIPO, several areas merit further investigation:
+Personalization and User Adaptation.
+TIPO currently does not directly incorporate personalization methods (e.g.
+LoRA [30]).
+Incorporating such modules could further align generated outputs to individual user preferences, enabling more tailored creative assistance.
+Investigating joint training strategies that integrate personalized fine-tuning represents a valuable future direction.
+Generalization to Uncommon Compositions.
+While TIPO performs effectively on typical and broadly represented concepts, it may still struggle with highly unconventional combinations (e.g., “avocado chair”).
+Future research should systematically evaluate TIPO’s performance on rare or unusual prompts (but considered creative prompts), potentially exploring auxiliary training or adaptation methods to ensure broader compositional robustness.
+Capacity and Backbone Scalability.
+Given TIPO model’s architecture [66, 65], its capacity for handling highly complex or lengthy prompts could be inherently limited.
+Future work could investigate more extensive or specialized architectures enhance compositional and semantic capacity, ideally without compromising real-time inference efficiency, are crucial for practical deployment.

### v4

@@
 Unlike resource-intensive methods based on large language models (LLMs) or reinforcement learning (RL), TIPO provides computational efficiency and scalability, opening new possibilities for effective, automated prompt engineering in T2I tasks.
-We provide visual results, human preference report to investigate TIPO’s effectiveness.
-Experimental evaluations on benchmark datasets demonstrate substantial improvements in aesthetic quality, significant reduction of visual artifacts, and enhanced alignment with target distributions along with significant human preference proficiency.
-These results highlight the importance of targeted prompt engineering in text-to-image tasks and indicate broader opportunities for automated prompt refinement.
+Extensive experiments across multiple domains demonstrate that TIPO delivers stronger text alignment, reduced visual artifacts, and consistently higher human preference rates, while maintaining competitive aesthetic quality.
+These results highlight the effectiveness of distribution-aligned prompt engineering and point toward broader opportunities for scalable, automated refinement in text-to-image generation.
+National Tsing Hua University  Nanyang Technological University  Yonsei University
+Onoma AI  Anhui Medical University  Seoul National University  Comfy Org Research
+Corresponding Author  Equal Contribution
 1 Introduction
-The rapid proliferation of Text-to-Image (T2I) generative models have revolutionized artistic creation [48, 7, 20, 57, 53, 54, 60, 55, 50, 59, 15, 14, 35, 21, 8].
+The rapid proliferation of Text-to-Image (T2I) generative models has revolutionized artistic creation (Ossa et al., 2024; Betker et al., 2023; Esser et al., 2024a; Saharia et al., 2022; Ramesh et al., 2021; 2022; Shi et al., 2020; Rombach et al., 2022a; Podell et al., 2024; Sauer et al., 2024; Chen et al., 2024b; a; Li et al., 2024b; Esser et al., 2024b; black-forest-labs, 2024).
 These models offer direct control over generative visual content via text prompts.
 To achieve precise control, modern T2I architectures are often trained on lengthy, detailed text descriptions, which may consist of individual, formatted tags of objects, backgrounds, styles, or complex, integrated paragraphs outlining image content and layout.
-However, the increasing complexity of prompts also forces users to iteratively refine their inputs, making high-quality T2I artwork accessible only to those with significant prompt engineering expertise.
-Recently, efforts have been made to reduce the reliance on human expertise through prompt optimization, i.e., expanding and refining a user’s primitive input into a more detailed prompt to enhance generation quality.
-A straightforward approach is to leverage pre-trained Large Language Models (LLMs) to rewrite prompts in a zero-shot manner [42].
+However, the increasing complexity of prompts often forces users to iteratively refine them to convey intent.
+Moreover, most state-of-the-art T2I models are aesthetically fine-tuned (e.g., on LAION-aesthetics (Podell et al., 2024)) to favor nuanced artistic and stylistic cues, making high-quality T2I artwork mainly accessible to those with significant artistic expertise.
+Extensive efforts have been made to reduce the reliance on human expertise through prompt optimization, i.e., expanding and refining a user’s primitive input into a more detailed prompt to enhance generation quality.
+As shown in Figure 1(a), a straightforward approach is to leverage pre-trained Large Language Models (LLMs) to rewrite prompts in a zero-shot manner (mañas2024improvingtexttoimageconsistencyautomatic).
 Yet, LLMs are primarily trained on general natural language, such as paragraphs and dialogues, which differ significantly from the structured prompts used for T2I models.
 This discrepancy often leads to additional effort in crafting LLM prompts and increased misalignment between generated images and intended prompts.
-A more effective approach is to train LLMs directly on prompt data collected from model users [5, 17].
+Figure 1(b) shows a more effective approach: training LLMs directly on prompt data collected from model users (AUTOMATIC, 2022; daspartho, 2022).
 While promising, this method is inherently constrained by the varying levels of user expertise, often resulting in inconsistent or suboptimal outputs.
-Recent work [26] trains LLM with reinforcement learning, where aesthetic scores of generated images serve as rewards.
-However, the reinforcement learning is performed on one specific T2I model with high computational cost, hindering its application to a broader variety of models.
+Recent work (Hao et al., 2023) trains LLMs with reinforcement learning, where aesthetic scores of generated images serve as rewards, as depicted in Figure 1(c).
+However, reinforcement learning is performed on one specific T2I model with high computational cost, hindering its application to a broader variety of models.
 Figure 1: Comparison of prompt optimization methods using LLM.
-(a) uses instructions for prompting but its understanding is constrained by the LLM’s knowledge base, not T2I model.
+(a) uses instructions for prompting but its understanding is constrained by the LLM’s knowledge base, not the T2I model.
 (b) relies on a curated prompt database, enhancing detail but limiting variety by not fully leveraging the T2I model’s learned distribution.
 (c) optimizes using the scorer with RL, requiring multi-turn inference with additional cost.
-(d) aligns prompts with the T2I model’s training distribution, ensuring detailed and various topic-related prompt generation that fits the target T2I model.
-Deviating from existing work, we argue that the optimal prompt lies in T2I models’ training set.
-Precisely, when a prompt aligns with the data distribution used for model training, the model can better interpret the prompt, resulting in improved text alignment and higher image quality.
-Based on this insight, we introduce TIPO (Text to Image with text pre-sampling for Prompt Optimization), an innovative framework designed to enhance T2I generative models.
-At its core lies a lightweight language model trained on multiple meticulously designed pretext tasks that transform prompts into tags or sentences from coarse to fine.
-Leveraging the trained model, TIPO can progressively and seamlessly adapt to various user input types, generating accurate and robust prompts for various T2I models.
-In addition, our training data is derived from a curated text description dataset comprising over 30 million text-image pair samples and more than 20 billion tokens, ensuring comprehensive coverage of T2I training sets.
-Extensive experiments on test prompts from in- and out-of-domain demonstrate that TIPO consistently outperforms state-of-the-art methods in fidelity, aesthetics, and text alignment.
-Detailed analysis shows that TIPO has up to 62.80% win rate in user preference and 25% improvement in runtime efficiency over the second-best baseline.
-To summarize, our contributions are at least three-fold:
+(d) aligns prompts with the T2I model’s training distribution, ensuring detailed and diverse prompt generation that fits the target T2I model.
+In contrast to prior work, we argue that good prompts should align with the large-scale text distributions of T2I models’ training, including those emphasized in aesthetic fine-tuning.
+Such alignment allows models to better interpret user intent and leverage their learned priors, resulting in stronger text alignment and overall image quality.
+Based on this insight, we introduce TIPO (Text to Image with text pre-sampling for Prompt Optimization), a framework that brings prompt optimization into the domain of large-scale multi-task pre-training.
+TIPO is supported by a curated 30M-pair, 40B-token caption corpus, which is filtered and balanced to maximize compatibility with leading T2I models and to preserve aesthetic quality.
+On top of this corpus, we design a suite of pretext tasks that reformulate raw user inputs, including both concise natural sentences and tag-based prompts, into enriched and distribution-consistent forms.
+Through this multi-task sampling pipeline, a lightweight language model expands (rather than fully rewrites) user prompts, preserving original semantics while enriching details that are diverse, edit-friendly, and aligned with T2I training distributions.
+Extensive experiments on both in-domain and out-of-domain prompts show that TIPO consistently outperforms strong baselines, achieving a 62.8% win rate in human preference (validated by more than 1,400 pairwise comparisons from 221 volunteers), and providing up to a 29.4% runtime efficiency improvement.
+Figure 1 illustrates the conceptual differences between existing methods and TIPO.
+To summarize, our contributions are at least threefold:
   1. 1.
-    We introduce TIPO, a prompt optimization framework that leverages the large-scale text data distributions used in text-to-image (T2I) training.
+    We introduce TIPO, a prompt optimization framework that leverages the large-scale text distributions used in text-to-image (T2I) training.
   2. 2.
-    We train a multi-task language model, which adaptively and progressively refines user inputs from either tag-based or natural language prompts into unified formats, enhancing compatibility across a broad spectrum of T2I models.
+    We train a lightweight multi-task language model that progressively refines both tag-based and natural language inputs into unified prompts, enhancing compatibility across a broad spectrum of T2I models.
   3. 3.
-    Extensive experiments demonstrate that TIPO achieves superior image quality, strong text alignment, higher human preference, and improved runtime efficiency across both in-domain and out-of-domain prompts, highlighting its great potential in real-world applications.
+    Extensive experiments demonstrate that TIPO achieves superior image quality, stronger text alignment, higher human preference, competitive aesthetic quality, and improved runtime efficiency against strong baselines with SOTA T2I models, highlighting its practical value.
 2 Related Work
-Prompt optimization for T2I models typically leverages language models to refine user inputs.
-They can be broadly classified into two categories: (1) Model-specific strategies that tailor prompts for a particular T2I model, and (2) Universal strategies that aim to improve prompt quality across a variety of T2I models.
-2.1 Model-specific Strategies
+Prompt optimization for T2I models typically leverages language models, which can be broadly classified into two categories: (1) Model-specific strategies that tailor prompts for a particular T2I model, and (2) Universal strategies that improve prompt quality across a variety of T2I models.
+Model-specific Strategies
 T2I models generate images whose quality is often measured using metrics such as fidelity, aesthetics, and user preference.
 These metrics facilitate reinforcement learning approaches that optimize prompts for a specific T2I model.
-For instance, Promptist [26] fine-tunes a pre-trained language model by using CLIP relevance scores as rewards.
-Similarly, PAE [44] extends this approach by generating dense text embeddings rather than discrete text tokens, with additional control vectors during online reinforcement learning.
+For instance, Promptist (Hao et al., 2023) fine-tunes a pre-trained language model by using CLIP relevance scores as rewards.
+Similarly, PAE (Mo et al., 2024) extends this approach by generating dense text embeddings rather than discrete text tokens, with additional control vectors during online reinforcement learning.
 However, these methods are computationally intensive, often struggling with a larger number of training prompts.
@@
 In contrast, our method leverages over 30 million text descriptions to cover a wide range of high-quality prompts, ensuring compatibility with a broad spectrum of T2I models.
-2.2 Universal Strategies
+Universal Strategies
 To reduce the dependency on specific T2I models, some researchers have focused on refining prompts solely using language models.
-For example, CogView3 [71] employ GLM-4 [24], and Lee et al.
-[33] employ GPT-J and Text Style Transfer (TST) techniques, respectively, for prompt enhancement.
+For example, CogView3 (Zheng et al., 2024) employ GLM-4 (GLM et al., 2024), and Lee et al.
+(2024) employ GPT-J and Text Style Transfer (TST) techniques, respectively, for prompt enhancement.
 However, both of them rely heavily on the LLM’s inherent understanding of visual content descriptions, which may result in a misalignment with the diverse requirements of various T2I models.
-Alternatively, other approaches collect high-quality prompts from T2I model users to fine-tune or train LLMs [5, 63, 17].
+Alternatively, other approaches collect high-quality prompts from T2I model users to fine-tune or train LLMs (AUTOMATIC, 2022; succintly, 2022; daspartho, 2022).
 Such methods, however, are limited by the inconsistent expertise of users.
+More recently,  He et al.
+(2025); Liu et al.
+(2024b) leverage vision language models to optimize prompts in an iterative loop: a user prompt generates images via a T2I model, the prompt–image pairs are evaluated by the VLM to suggest refinements, and the revised prompts are reapplied to T2I generation for continual improvement.
+While such iterative refinement can improve quality, the distribution mismatch between VLMs and T2I training data persists.
 Conversely, our approach constructs both tag-based and natural language prompts using a large-scale dataset of image-text descriptions, thereby aligning closely with the text distributions underlying T2I models.
 3 Preliminaries
-Notations. We present the notations in our work:
-  • $\mathcal{P}$: The space of all possible prompts.
-  • $\mathcal{I}$: The space of all possible images.
-  • $p$: User-provided prompt.
-  • $p_{o}$: Optimized prompt.
-  • $p_{ps}$: Sampled modified prompt derived from p.
-  • $\mathcal{M}$: Metadata (e.g., style tags) aiding prompt transformation.
-  • $\mathcal{T}$: Transformation function for pre-sampling.(e.g., Random augmentation, TIPO or LLM)
-  • $T_{p},T_{ps}$: Sets of tags for original and pre-sampled prompts.
-  • $S_{p},S_{ps}$: Ordered lists of natural language descriptions for original and pre-sampled prompts.
-Text-to-Image Model. A text-to-image (T2I) model is a mapping function
-$f(p)\;\;\to\;\;\mathcal{I}_{p}\subseteq\mathcal{I},$
-which maps a user-provided text prompt $p$ to a subset of images $\mathcal{I}_{p}$ corresponding to the prompt.
+We present the formal definition of T2I models and the problem statement of this work.
+Text-to-Image Model. A text-to-image (T2I) model defines a conditional distribution
+$\mathcal{I}_{p}=P(x\mid p),\quad x\in\mathcal{X},$
+which maps a prompt $p$ to a distribution of images over the space of all possible images $\mathcal{X}$.
 Problem statement.
-Let $\mathcal{I}_{u}$ denote the user’s intended distribution over images.
-The task of Prompt optimization is to generate a prompt $p_{o}$ that minimizes a distance $d$ between the T2I output distribution $\mathcal{I}_{p}$ and $\mathcal{I}_{u}$:
+Let $\mathcal{I}_{u}$ denote the user’s intended distribution over $\mathcal{X}$.
+The task of prompt optimization is to find an optimized prompt $p_{o}$ from the space of all possible prompts $\mathcal{P}$ to minimize a distance $d$ between the T2I output distribution $\mathcal{I}_{p}$ and the intended distribution $\mathcal{I}_{u}$:
 $p_{o}\;=\;\arg\min_{p\in\mathcal{P}}\;d\bigl(\mathcal{I}_{p},\,\mathcal{I}_{u}\bigr),$
-where $d$ is a suitable metric (e.g., Frechet Inception Distance) comparing two image distributions.
+where $d(\cdot,\cdot)$ is a distance between image distributions (e.g., Fréchet Inception Distance).
+4 Methodology
 […]
-4 Methodology
 We aim to optimize user prompts to enhance image generation quality.
@@
 Although the text descriptions used for T2I model training are notably diverse, most are image captions that fall into two broad categories: tag-based and natural language (NL)-based captions.
-Tag-based captions, such as those in the Danbooru2023 dataset [46, 68], use comma-separated, succinct terms to describe image content.
-In contrast, NL-based captions, typically generated by language models with visual capabilities [37, 1, 2, 34, 6, 16, 67, 18], may comprise multiple sentences.
+Tag-based captions, such as those in the Danbooru2023 dataset (nyanko202, 2023; Yeh, 2024b), use comma-separated, succinct terms to describe image content.
+In contrast, NL-based captions, typically generated by language models with visual capabilities (Liu et al., 2023; Agrawal et al., 2024; OpenAI, 2024; Li et al., 2024a; Bai et al., 2023; Dai et al., 2024; Xiao et al., 2024; Deitke et al., 2024), may comprise multiple sentences.
 We represent both types using a unified text set $T=\{t_{1},t_{2},t_{3},\ldots,t_{n}\}$, where each element is an individual tag or sentence.
@@
 Such prompts diverge from typical user input and pose alignment challenges.
-To mitigate this, we construct a simpler subset $T_{s}\subset T$ by removing some tags and sentences from the original set [^1] For long sentences, off-the-shelf text splitters (e.g., FastText) are used to segment them into multiple sentences., as detailed in Section 4.2.
+To mitigate this, we construct a simpler subset by removing some tags and sentences from the original prompt 1set, as detailed in Section 4.2.
 4.2 Formatted Prompt Construction
@@
 First, we incorporate the common metadata present in these image captions, typically represented as <Category>: <Content>.
-These metadata categories primarily include artist, copyright, aspect ratio, quality, and year (e.g., quality: masterpiece, artist: Picasso).
+These metadata categories primarily include style, aspect ratio, quality, and year (e.g., quality: masterpiece, style: Impressionist).
 This structured metadata is intuitive for users to read and edit, while also providing strong guidance to downstream T2I models on the generation scope.
@@
 We first randomly shuffle the complete set of tags $T=\{t_{1},t_{2},t_{3},\ldots,t_{n}\}$ from a given image caption.
-Then, we construct a simple tag set $T_{s}=\{t_{1},t_{2},\ldots,t_{m}\}$ by randomly selecting $m<n$ tags.
+Then, we construct a simpler tag set $T_{s}=\{t_{1},t_{2},\ldots,t_{m}\}$ by randomly selecting $m<n$ tags.
 The prompts are constructed as:
 $p_{s}=\text{concat}(T_{s}),\quad p_{o}=\text{concat}(T)$
-Because $p_{s}$ is always a prefix of $p_{o}$, the language model can readily expand the simple tag-based prompt into its complete version.
-For NL-based prompts, however, this strategy cannot be applied directly because the first sentence often contains crucial information [25] and the order of sentences significantly influences the caption’s semantics.
+Here, $p_{s}$ and $p_{o}$ denote the simple and original prompts, respectively.
+By ensuring that $p_{s}$ is always a prefix of $p_{o}$, the language model can readily expand simple tag-based prompts into complete versions.
+For NL-based prompts, however, this strategy cannot be applied directly because the first sentence often contains crucial information (Godbole et al., 2024) and the order of sentences significantly influences the caption’s semantics.
 Therefore, we preserve the first sentence and randomly drop some of the subsequent sentences without changing their order.
@@
 4.3 Text Pre-sampling
-We aim to align user input with the constructed high-quality training prompts $p_{o}$ via pre-sampling, which stands for “text sampling before image sampling” - a process that first constrains the target distribution in text space before proceeding to image generation.
-Conceptually, while pre-sampling strategies generally add more text tokens to user input, the outcomes of different strategies may still vary (as illustrated in Figure 2).
-To enrich the image content while mitigating undesired output, we perform pre-sampling based on a language model trained on $p_{o}$.
-A straightforward pre-sampling strategy is to do text completion, i.e., directly append output tokens to the end of user input.
-However, we posit that the user input is unlikely to be a part of a good prompt.
-The na’́ive text-completion strategy prioritizes the user’s original input, resulting in worse outcomes as evidenced by Section 5.6.
-Figure 3: TIPO’s task flow within a single generation.
-Starting from any node, each arrow represents a sequential extension step, with other prompts used as metadata if provided.
-This task design enables efficient and flexible prompt extension across multiple tasks.
-We propose the core technique of TIPO, a flexible pre-sampling mechanism that divides the prompt optimization process into three interconnected subtasks: enriching tag sequences, extending NL prompts, and refining NL prompts.
-As shown in Figure 3, TIPO uses both tag and NL prompts interchangeably as input sources to generate detailed outputs in various forms.
-For instance, TIPO can begin with a short NL prompt to produce a detailed tag sequence, denoted as short_to_tag.
-Training data for the language model is constructed by concatenating different types of $p_{o}$ with task-specific metadata embedded into the text.
-The individual tasks are defined as follows:
-  • tag_to_long: Use tags as metadata to generate a new NL prompt.
-  • long_to_tag: Use an NL prompt as metadata to extend a tag sequence.
-  • short_to_tag: Use the simple prompt $p_{s}$ as metadata to extend a tag sequence.
-  • short_to_long: Use a user-provided NL prompt as metadata to generate a refined, detailed prompt.
-We also consider composite tasks that can be performed within a single forward pass. These tasks offer the language model a holistic view of the objective while reducing computational overhead:
-  • short_to_tag_to_long: Use a user-provided NL prompt or tag sequence as metadata to generate a refined, detailed prompt.
-  • short_to_long_to_tag: Use a user-provided NL prompt or generated NL prompt as metadata to extend a tag sequence.
-  • tag_to_short_to_long: Use user-provided tags or NL prompts as metadata to generate a refined NL prompt.
+We aim to reformulate user input into forms that better align with the high-quality training text distribution $p_{o}$ via pre-sampling, which stands for “text sampling before image sampling.” A naïve strategy is plain text completion, where tokens are directly appended to the user input in an unstructured manner.
+Such completion often mirrors the low-quality phrasing of the input, deviates from the distribution of high-quality prompts, and produces inferior generations.
+On the other hand, a full rewrite risks deviating from the user’s original intention.
+To balance these issues, TIPO preserves the original input and appends a structured, distribution-consistent expansion.
+This appended segment is typically paragraph-like or tag-like, making it both informative and easy to edit or remove (see Figure 2 for an illustration and Appendix E for concrete examples).
+We propose the core technique of TIPO, a flexible pre-sampling mechanism that decomposes prompt optimization into three subtasks: enriching tag sequences, extending natural language (NL) prompts, and refining NL prompts.
+For example, a short NL prompt can be expanded into a detailed tag sequence (short_to_tag), as illustrated in Figure 3.
+We further distinguish between basic tasks, which perform a single transformation, and composite tasks, which chain multiple transformations within a single forward pass.
+The latter expose the model to more holistic training signals while reducing computational overhead.
+Table 1 summarizes all tasks and their input–output forms.
+Task Description
+Basic tasks
+tag_to_long Generate a new NL prompt given tags.
+long_to_tag Extend a tag sequence given an NL prompt.
+short_to_tag Extend a tag sequence given a short/simple prompt $p_{s}$.
+short_to_long Generate a refined, detailed NL prompt given a user-provided NL prompt.
+Composite tasks
+short_to_tag_to_long From a short NL prompt or tags, produce a refined NL prompt.
+short_to_long_to_tag From a short or generated NL prompt, extend a tag sequence.
+tag_to_short_to_long From tags or NL prompts, generate a refined NL prompt.
+Table 1: Pre-sampling tasks in TIPO. Basic tasks focus on one-step transformations, while composite tasks combine two basic tasks within a single forward pass.
 We randomly select from the aforementioned tasks during training to enhance model generalization.
 By extensively training on these tasks, TIPO can seamlessly adapt to various input types, flexibly refining user input whether it consists of tags, short sentences, or long sentences.
-Figure 4 illustrates a scenario where both tag captions $T_{s}$ and short NL captions $S_{s}$ are available.
+Figure 3 (b) illustrates a scenario where both tag captions $T_{s}$ and short NL captions $S_{s}$ are available.
 In such cases, TIPO processes each input type separately to maintain clarity and coherence:
-$\begin{array}[]{ll}S_{s}&=\{\textsf{A young girl with long hair...}\},\quad\texttt{metadata}=\emptyset\\ T_{s}&=\left\{\textsf{outdoors, scenery, water, wind, landscape, \ldots}\right\}\end{array}$
+[…]
 The generation proceeds sequentially as follows:
@@
 This progressive process enables TIPO to build prompts that are both detailed and contextually aligned with the user’s input.
-Figure 4: An example scenario of TIPO workflow.
-(a) A generated image and prompts.
-(b) Prompt optimization of TIPO, from simple user input $p_{s}$ to detailed final output $p_{d}$.
-The shading from gray to light sky blue represents an increase in context richness in the prompt.
-4.4 Implementation Details
-TIPO can utilize various types of causal autoregressive language models by design. In practice, we leverage the LLaMA architecture [64, 65, 3] for simplicity, with 200M and 500M parameters.
-  • TIPO-200M: Pretrained on the Danbooru2023 [46, 69] and GBC10M [29] datasets for 5 epochs, then fine-tuned with Danbooru2023, GBC10M, and CoyoHD11M [12] for 3 epochs, covering approximately 40 billion tokens.
-  • TIPO-500M: Pretrained on the Danbooru2023, GBC10M, and CoyoHD11M datasets for 5 epochs, covering approximately 30 billion tokens.
+(a) Scenery tag only input
+(b) Tag + short NL input
+Figure 3: Example prompt optimization paths in TIPO.
+(a) shows generation from a single tag input, while (b) uses both tag and natural language input.
+These illustrate representative pipelines, not the full range of use cases.
+Blue shading indicates increasing prompt richness.
+Implementation Details
+In implementation, TIPO adopt the LLaMA architecture (Touvron et al., 2023a; b; AI@Meta, 2024) [^1] The multi-task design of TIPO is compatible with many autoregressive language models (e.g., GPT, LLaMA, Qwen, etc.).
+Intuitively, adopting more advanced backbones could further enhance efficiency and effectiveness, but such exploration would require extra training cost, which we defer to future work., with all experiments are conducted with a 200M-parameter model  [^2] We also trained TIPO-100M and 500M variants to analyze the impact of model scales.
+See Appendix C..
+Our training dataset is about 40 billion tokens curated from Danbooru2023 (nyanko202, 2023; Yeh, 2024a), GBC10M (Hsieh et al., 2024), and CoyoHD11M (CaptionEmporium, 2024).
 5 Experiments
 5.1 Experimental Settings
-Baselines. We compare with state-of-the-art prompt optimization methods as follows:
-  • GPT [2] series are powerful large language models capable of zero-shot prompt optimization. We use GPT-4o-mini for lower cost and better efficiency.
-  • MagicPrompt [17] trains GPT-2 on high-quality prompts collected from stable-diffusion model users.
-  • Promptist [26] optimizes user input into model-preferred prompts via reinforcement learning.
-Evaluation Metrics.
-We employ four latest metrics FDD [61], Aesthetic Score [19], AI Corrupt Score [45], and Vendi Score [22] to measure the quality of generated images.
-Specifically, FDD (Frechet Dino Distance) measures the image fidelity using the similarity between generated and dataset images via DinoV2 features [47], which better aligns with human perception than traditional FID [28].
-Aesthetic Score is computed via Aesthetic Predictor V2.5 [19], quantifying visual appeal, composition quality, and artistic merit.
+Baselines and T2I Models
+We compare against representative prompt-optimization baselines: GPT-4o-mini for zero-shot rewriting (OpenAI, 2024), MagicPrompt, which fine-tunes GPT-2 on community-collected prompts (daspartho, 2022), Promptist, which applies reinforcement learning for model-preferred prompt optimization (Hao et al., 2023), and Gemini-2.0-flash-image  [^3] https://developers.googleblog.com/experiment-with-gemini-20-flash-native-image-generation/, which uniquely serves both as a prompt-refinement baseline and a T2I generator.
+For image generation backbones, our main experiments adopt SDXL-base-1.0 (Podell et al., 2024), Illustrious v3.5 (v-pred), Kohaku-XL-Zeta, and Stable Diffusion 3.5 Large(Esser et al., 2024a).
+To further assess generalization, we additionally evaluate on four diverse backbones with undisclosed training data: FLUX.1-dev (black-forest-labs, 2024), Omnigen2 (Wu et al., 2025), Lumina-2 (Qin et al., 2025), and HiDream-I1 (Cai et al., 2025).
+The details of T2I models are provided in Appendix B.
+Evaluation Metrics
+We employ four latest metrics FDD (Stein et al., 2023), Aesthetic Score (discus0434, 2024), AI Corrupt Score (narugo1992, 2023), and Vendi Score (Friedman and Dieng, 2022) to measure the quality of generated images.
+Specifically, FDD (Fréchet DINO Distance) quantifies fidelity by comparing the distribution of DINOv2 features (Oquab et al., 2023) between reference images in the evaluation dataset and images generated from the corresponding captions, which better aligns with human perception than traditional FID (Heusel et al., 2017).
+Aesthetic Score is computed via Aesthetic Predictor V2.5 (discus0434, 2024), quantifying visual appeal, composition quality, and artistic merit.
 AI Corrupt Score detects technical flaws in generated images by identifying visual artifacts.
 Vendi Score quantifies image diversity by calculating the von Neumann entropy from a normalized cosine similarity matrix using DinoV2 embeddings.
-Evaluation Protocols.
+Notably, Vendi is defined directly on feature-space dispersion and is sensitive to non-semantic variations such as low-level noise or artifacts.
+As a result, it “should be used alongside a quality metric” (Friedman and Dieng, 2022) and a higher value alone does not always correspond to more meaningful semantic diversity.
+Evaluation Protocols
 Our proposed TIPO leverages large-scale image caption datasets for training, which overlap with the training text distributions of many T2I models.
-Following Promptist [26], we divide our experiments into two settings: (1) In-domain, where the T2I model’s training texts overlap with those used by TIPO, and (2) Out-of-domain, where no overlap exists.
-5.2 In-domain Tag-based Prompt Optimization
-To assess prompt optimization performance on tag-based prompts, we generate scenery images, as they contain abundant descriptive tags for objects and backgrounds.
-We randomly sample 32,768 tag-based prompts from Danbooru2023 [69, 46], shuffle and concatenate the scenery tags into new prompts (thereby preventing data leakage of the original captions), and generate one image per prompt using Kohaku-XL-Zeta, an SDXL [50, 46] variant fine-tuned on Danbooru2023.
-This evaluation tests the in-domain capabilities, as Danbooru2023 is also used during TIPO training.
-Metric Original GPT MagicPrompt Promptist TIPO (Ours)
-FDD $\downarrow$ 0.3558 0.5414 0.3247 0.2350 0.2282
-Aesthetic Score $\uparrow$ 5.0569 6.3676 6.1609 5.9468 6.2571
-AI Corrupt Score $\uparrow$ 0.4257 0.7490 0.5024 0.5669 0.9195
-Vendi Score $\uparrow$ 16.8135 8.6632 11.9014 14.3273 13.3070
-Table 1: Performance on in-domain tag-based prompts, with best results bolded and the second-best underlined.
-The results in Table 1 reveal two key insights.
-First, MagicPrompt and Promptist, which rely on user prompts or reinforcement learning, underperform in Aesthetic and AI Corrupted Scores due to the quality or quantity limitations of their collected samples (e.g., Promptist uses 90K samples, limited by the high reinforcement learning cost).
-In contrast, GPT and TIPO benefit from large-scale training corpora ($>$30M samples), yielding higher-quality outputs.
-Second, TIPO achieves the best FDD by a substantial margin over GPT, which can be attributed to its superior distribution alignment with T2I models.
-5.3 In-domain NL-based Prompt Optimization
-We evaluate the prompt optimization performance on NL-based prompts by selecting 10,000 short prompts and 10,000 long prompts from CaptionEmporium [12] and GBC [29] as test prompts.
-In particular, since the long prompts are much longer than typical user input, we truncate them to two sentences ($<$ 40 words) to simulate real-world applications.
-We use SDXL-1.0-base [50] as the T2I model, whose training text data largely overlap with TIPO.
-Short Original GPT MagicPrompt Promptist TIPO(Ours)
-FDD $\downarrow$ 0.0957 0.1668 0.0980 0.1783 0.1168
-Aesthetic Score $\uparrow$ 5.8370 6.0589 5.8213 5.7963 5.8531
-AI Corrupt Score $\uparrow$ 0.7113 0.6985 0.7064 0.6314 0.7131
-Vendi Score $\uparrow$ 38.1715 34.7139 38.1553 34.1270 37.0649
-Truncated Long Original GPT MagicPrompt Promptist TIPO(Ours)
-FDD $\downarrow$ 0.0955 0.1683 0.1247 0.2096 0.1210
-Aesthetic Score $\uparrow$ 5.7497 6.0168 5.8191 5.7759 5.8364
-AI Corrupt Score $\uparrow$ 0.6868 0.6712 0.6741 0.5925 0.7130
-Vendi Score $\uparrow$ 38.2533 34.8108 37.8412 33.5266 37.0900
-Table 2: Performance on in-domain NL-based prompts.
+Following Promptist (Hao et al., 2023), we divide our experiments into two settings: (1) In-domain, where the T2I model’s training texts overlap with those used by TIPO, and (2) Out-of-domain, where no overlap exists.
 (a) Scenery Tag Only
@@
 (c) Truncated Long Prompt
-(d) Truncated Prompt + TIPO
-Figure 5: Generated images from 4 types of prompts: (a) simple scenery tag, (b) scenery tag enhanced by TIPO, (c) truncated ($<$ 40 words) long prompt, (d) TIPO-enhanced truncated prompt.
+(d) Truncated + TIPO
+Figure 4: Generated images from 4 types of prompts: (a) simple scenery tag, (b) scenery tag enhanced by TIPO, (c) truncated (<< 40 words) long prompt, (d) TIPO-enhanced truncated prompt.
 TIPO adds detail and maintains variety, yielding coherent images from simple prompts.
+5.2 Experimental Results
+In-domain Tag-based Prompt Optimization
+To assess prompt optimization performance on tag-based prompts, we generate scenery images, as they contain abundant descriptive tags for objects and backgrounds.
+We randomly sample 32,768 tag-based prompts from Danbooru2023 (Yeh, 2024a; nyanko202, 2023), shuffle and concatenate the scenery tags into new prompts (thereby preventing data leakage of the original captions), and generate one image per prompt using Kohaku-XL-Zeta.
+This evaluation tests the in-domain capabilities, as Danbooru2023 is used during both TIPO and Kohaku-XL-Zeta training.
+The results in Table 2 reveal two key insights.
+First, MagicPrompt and Promptist, which rely on user prompts or reinforcement learning, underperform in Aesthetic and AI Corrupted Scores due to the quality or quantity limitations of their collected samples (e.g., Promptist uses 90K samples, limited by the high reinforcement learning cost).
+In contrast, GPT and TIPO benefit from large-scale training corpora (>>30M samples), yielding higher-quality outputs.
+Second, TIPO achieves the best FDD by a substantial margin over GPT, which can be attributed to its superior distribution alignment with T2I models.
+Table 2: Comprehensive performance comparison of TIPO against baselines across different in-domain prompt types.
+Metrics are marked with $\uparrow$ (higher is better) or $\downarrow$ (lower is better).
+For OOD tests, the ’Original’ baseline and the FDD metric was not applicable.
+In the table, Aesthetic refers to Aesthetic Score, Corrupt to AI Corrupt Score, and Vendi to Vendi Score.
+TIPO demonstrates significant improvements, achieving the highest average rank among all baselines.
+Best results are in bold and second-best are underlined.
+Prompt Type / Task Metric Original GPT MagicPrompt Promptist TIPO
+In-domain Tag-based Prompts FDD $\downarrow$ 0.3558 0.5414 0.3247 0.2350 0.2282
+Aesthetic $\uparrow$ 5.0569 6.3676 6.1609 5.9468 6.2571
+Corrupt $\downarrow$ 0.5743 0.2510 0.4976 0.4331 0.0805
+Vendi $\uparrow$ 16.814 8.663 11.901 14.327 13.307
+In-domain NL-based (Short) FDD $\downarrow$ 0.0957 0.1668 0.0980 0.1783 0.1168
+Aesthetic $\uparrow$ 5.8370 6.0589 5.8213 5.7963 5.8531
+Corrupt $\downarrow$ 0.2887 0.3015 0.2936 0.3686 0.2870
+Vendi $\uparrow$ 38.172 34.714 38.155 34.127 37.065
+In-domain NL-based (Truncated Long) FDD $\downarrow$ 0.0955 0.1683 0.1247 0.2096 0.1210
+Aesthetic $\uparrow$ 5.7497 6.0168 5.8191 5.7759 5.8364
+Corrupt $\downarrow$ 0.3132 0.3288 0.3259 0.4075 0.2870
+Vendi $\uparrow$ 38.253 34.811 37.841 33.527 37.090
+Out-of-Domain (OOD) Test Aesthetic $\uparrow$ N/A 6.7125 6.4507 6.3924 6.0536
+Corrupt $\downarrow$ N/A 0.0518 0.1423 0.0947 0.0720
+Vendi $\uparrow$ N/A 8.9718 15.872 16.489 21.571
+Overall Average Rank $\downarrow$ 2.58 3.00 3.00 3.87 2.07
+Table 3: TIPO on T2I models with undisclosed training data. Significant improvements are in bold.
+Model Variant Aesthetic ↑ AI Corrupt ↓ FDD ↓ Vendi ↑
+FLUX.1-dev Original 5.2029 0.1202 0.1185 36.2597
+TIPO 5.2746 0.0938 0.1202 35.6489
+Omnigen2 Original 5.2629 0.1110 0.1373 33.7724
+TIPO 5.0661 0.1187 0.1253 34.2681
+Lumina-2 Original 5.2588 0.0981 0.1318 34.7599
+TIPO 5.4105 0.0916 0.1286 34.4655
+HiDream-I1 Original 5.7768 0.1055 0.1444 34.6318
+TIPO 5.8143 0.1019 0.1379 34.3544
+Gemini-2.0-Flash-Image Self-refined 5.2584 0.0928 0.8084 32.7422
+TIPO 5.2649 0.0700 0.7964 32.1136
+In-domain NL-based Prompt Optimization
+We evaluate the prompt optimization performance on NL-based prompts by selecting 10,000 short prompts and 10,000 long prompts from CaptionEmporium (CaptionEmporium, 2024) and GBC (Hsieh et al., 2024) as test prompts.
+In particular, since the long prompts are much longer than typical user input, we truncate them to two sentences (<< 40 words) to simulate real-world applications.
+We use SDXL-1.0-base as the T2I model, whose training text data largely overlap with TIPO.
 Table 2 demonstrates that TIPO achieves either the best or second-best scores in Aesthetic and AI Corrupt Score by effectively enriching the original prompt with appropriate textual elements while rarely introducing extraneous noise.
 While all methods compromise fidelity and diversity, as reflected in the FDD and Vendi Score, TIPO remains competitive because it maintains small semantic deviation from the original sentences via progressive refinement.
-(a) Oringial Prompt
-(b) Text Completion
-(c) TIPO
-Figure 6: Qualitative comparison of prompt completion versus TIPO (ours).
-(a) shows images generated by original prompts.
-(b) depicts images generated by text completion using TIPO-500M.
-(c) shows images generated by TIPO.
-5.4 Out-of-domain Performance
+Out-of-domain Performance
 Some recent T2I models are trained on proprietary images and captions.
-As a representative example, SD-3.5-Large[20] is trained on private images captioned with CogView[71], which differ markedly from the texts used to train TIPO.
+As a representative example, SD-3.5-Large is trained on private images captioned with CogView (Zheng et al., 2024), which differ markedly from the texts used to train TIPO.
 To evaluate model performance in this out-of-domain scenario, we generate 8,192 original tag- and NL-based prompts using the baseline GPT-4o-mini rather than relying on existing prompt datasets.
 We apply the remaining methods to these prompts and assess their performance.
-GPT MagicPrompt Promptist TIPO(Ours)
-Aesthetic Score $\uparrow$ 6.7125 6.4507 6.3924 6.0536
-AI Corrupt Score $\uparrow$ 0.9482 0.8577 0.9053 0.9280
-Vendi Score $\uparrow$ 8.9718 15.8721 16.4891 21.5706
-Table 3: Performance on the out-of-domain T2I model
 As shown in Table 2 , SD-3.5-Large faithfully generates images that align well with GPT-produced prompts.
@@
 TIPO optimization enriches the prompts with additional details that harmonize with the original themes, significantly enhancing the diversity of the generated images.
-5.5 Human Preference Evaluation
+Compatibility with other T2I Models
+In addition to the above baselines with publicly known training data, we further evaluate TIPO on recent models with undisclosed training sources: FLUX.1-dev (black-forest-labs, 2024), Omnigen2 (Wu et al., 2025), Lumina-2 (Qin et al., 2025), HiDream-I1 (Cai et al., 2025), and Gemini-2.0-flash-image (Google, ), using 1,000 prompts sampled from the COYO/GBC datasets.
+Notably, Gemini-2.0-flash-image can perform its own prompt optimization, making it a strong closed-source baseline.
+As shown in Table 3, TIPO generally improves image quality and alignment relative to the baselines.
+For example, Lumina-2 and HiDream-I1 achieve higher aesthetic scores and lower corruption rates after TIPO optimization.
+Despite Gemini’s integrated optimization, TIPO still yields measurable improvements, highlighting the practical value of our lightweight, task-specific strategy.
+Overall, these results demonstrate that TIPO maintains robust compatibility with heterogeneous T2I models, even when their training data are unavailable and potentially mismatched with TIPO’s optimization corpus.
+Efficiency
+A key concern is whether TIPO’s iterative pre-sampling strategy introduces noticeable latency.
+Hence, we benchmark prompt-generation latency in Table 4, where TIPO reduces per-prompt latency with an improvement up to 29.4%.
+Details of training and inference are in Appendix H.4.
+Table 4: Prompt generation latency and relative speedup
+TIPO Promptist PromptExtend MagicPrompt
+Avg. Time (s) 1.03 1.46 1.38 1.14
+TIPO speedup vs. each +29.4% +25.6% +9.6%
+Human Preference Evaluation
 Quantitative metrics may not fully align with human preference.
 Therefore, we conducted a user study based on pairwise image comparisons between the original prompt, MagicPrompt, Promptist, and TIPO on over 1,400 images, gathering preferences from 221 volunteers.
-As illustrated in Figure 7, TIPO achieved the highest overall win rate at 51.3%, significantly outperforming competitors.
+As illustrated in Figure 5(a), TIPO achieved the highest overall win rate at 51.3%, significantly outperforming competitors.
 In out-of-domain scenarios, TIPO’s win rate increased to 52.5%, demonstrating consistently strong user preference across different contexts.
-For further results and statistics, please refer to Appendix F.
-Figure 7: The win-tie-lose rate on overall preference on all data and out-of-domain situation where images are generated by SD-3.5-Medium.
-Figure 8: ELO rating on overall preference on all data
-5.6 Ablation Study
-Original Completion TIPO
-Aesthetic Score $\uparrow$ 5.9507 5.9631 5.9960
-Corrupt Score $\uparrow$ 0.6799 0.8131 0.8971
-Vendi Score $\uparrow$ 19.7208 18.2896 13.9864
-Table 4: Ablation results
-We attribute TIPO’s success primarily to its multi-task pre-sampling design detailed in Section 4.3.
-However, a natural question arises: can a simple text completion approach achieve similar or better results?
-To validate our design, we compare TIPO’s pre-sampling strategy with a direct text completion using the same language model, TIPO-500M.
-As shown in Table 4, TIPO significantly improves aesthetic scores and reduces artifacts compared to text completion.
-For visual examples, please refer to Figure 6.
-For more ablation details and visualization, please refer to Appendix 5.6.
+For further results and statistics, please refer to Appendix G.
+(a) Pairwise comparison results showing win-tie-lose percentages for overall user preference. Evaluations cover ‘All Data’ and ‘Out-of-Domain’ sets for TIPO and baseline methods.
+(b) ELO ratings comparing TIPO and baseline methods across four criteria: Prompt Adherence, Image Quality, Aesthetic, and Overall. The dashed line indicates the initial ELO rating.
+Figure 5: Human preference evaluation demonstrates that TIPO consistently achieves higher user preference compared to baseline prompt optimization methods (Promptist, prompttext, magicprompt, and original prompts).
+All evaluated images were generated using SD-3.5-Medium.
+Prompt Distribution Alignment
+While TIPO consistently achieves the best or competitive results in image quality, it remains unclear whether such gains arise from a closer distributional alignment between its optimized prompts and the T2I models’ training text corpora.
+To verify this hypothesis, we sample from two representative T2I training sets, obtaining 1,000 natural-language captions from COYO and 1,000 tag-based captions from Danbooru2023.
+We then encode ground-truth captions and the outputs of all compared optimization methods using two widely adopted text encoders—T5-XXL (Raffel et al., 2020a) (used in SD3, Flux, and PixArt) and jina-embeddings-v3 (Sturua et al., 2024) (a popular recent text-embedding model).
+Finally, we measure the embedding-space alignment between the ground-truth captions and optimized prompts with Fréchet Distance (FD) and Maximum Mean Discrepancy (MMD with RBF kernel).
+As shown in Table 5, baseline methods show varying extents of alignment across different prompt types, text encoders, and metrics, while TIPO maintains consistently better alignment under all settings.
+This indicates that TIPO aligns well with the T2I training-text distribution in a prompt-compatible and encoder-insensitive manner.
+Table 5: Embedding-space distances (FD and MMD) between optimized prompts and T2I training corpora. Lower is better. Best results are in bold and second-best are underlined.
+Prompt Encoder Metric TIPO Promptist MagicPrompt GPT-4o-mini
+NL-short Jina FD 0.0322 0.1003 0.0385 0.1064
+MMD 0.0320 0.1501 0.0624 0.1699
+T5 FD 0.0704 0.2072 0.1441 0.1252
+MMD 0.1438 0.2914 0.1972 0.2297
+NL-trunc Jina FD 0.0309 0.1192 0.0493 0.0963
+MMD 0.0359 0.1700 0.0772 0.1642
+T5 FD 0.0674 0.2312 0.1884 0.1276
+MMD 0.1404 0.3147 0.2270 0.2323
+Tag-based Jina FD 0.1094 0.1891 0.1958 0.2479
+MMD 0.1539 0.2473 0.2415 0.3050
+T5 FD 0.0524 0.2080 0.2578 0.0728
+MMD 0.1846 0.3573 0.3948 0.2194
+Prompt–Image Alignment
+To further assess the semantic consistency between optimized prompts and their generated images, we compute CLIPScore between each prompt and its corresponding SD1.5-generated image using openai/clip-vit-large-patch14-336, on the same prompt sets as in the Prompt Distribution Alignment experiment.
+As shown in Table 6, TIPO achieves the strongest alignment for tag-based prompts.
+While MagicPrompt shows clear advantages on NL-based prompts, this is likely due to its 1.8M+ SD1.5 community training corpus, where prompts are typically explicit and stylistically strong, effectively eliciting the CLIP encoder to produce high alignment scores.
+However, such stylistic bias is often less preferred by mainstream users.
+In contrast, TIPO attains competitive performance on NL-based prompts without such bias, as reflected by the win rate in Table 11, where MagicPrompt vs.
+TIPO = 11:48.
+Table 6: CLIPScore between optimized prompts and T2I generated images. Higher is better. Best results are in bold and second-best are underlined.
+Prompt Type TIPO MagicPrompt GPT-4o-mini Promptist
+Tag-based 0.2217 0.1782 0.1774 0.1642
+NL-short 0.2413 0.2834 0.2378 0.2347
+NL-trunc 0.2310 0.2517 0.2275 0.2063
 6 Conclusion
@@
 By aligning user prompts with the intrinsic distributions of T2I training datasets, TIPO enhances semantic coherence, image fidelity, and diversity with minimal inference overhead.
-TIPO consistently outperforms existing prompt optimization methods, achieving superior performance across multiple evaluation metrics.
-Its ability to preserve user intent while enhancing prompt diversity and image quality makes it well-suited for both in-domain and out-of-domain scenarios, demonstrating robustness across various prompt formats and model architectures.
-Extensive user studies also confirmed TIPO’s strong alignment with human preferences, further highlighting its potential for practical applications.
+Experimental results show that TIPO consistently outperforms existing prompt optimization methods across multiple evaluation metrics, while extensive user studies confirm its strong alignment with human preferences.
+Despite these promising results, several aspects remain open for future exploration, such as generalization to out-of-distribution user input, model-specific adaptation, personalization, image-feedback-aware refinement, and large-scale model scaling.
+We discuss these limitations and potential extensions in Appendix J.
 To encourage wider adoption and facilitate reproducibility, we release our trained models and source code.
 We hope TIPO will inspire further advancements in efficient, scalable, and robust generative frameworks for creative systems.
+Ethics Statement
+Our work democratizes prompt optimization for text-to-image models, but automatic prompt enrichment can inherit biases from training data or be misused to produce misleading or harmful content.
+We therefore emphasize responsible use, including bias mitigation, transparency, and appropriate user controls.
+No human subjects or sensitive personal data were involved, and we avoided including identifiers or metadata that could raise copyright- or attribution-related concerns.
+Reproducibility Statement
+An anonymized repository with code and configuration files is provided in the supplementary materials, enabling reproduction of all reported results.
+The main paper and appendix reference the exact experimental settings, hyperparameters, random seeds, and evaluation scripts; data preprocessing steps and any additional resources needed to re-create the experiments are also documented there.
+For hardware, we report that all experiments were conducted on NVIDIA RTX 3090 and/or A6000 GPUs, with full details described in the appendix and repository.
 Appendix
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
 Appendix A Dataset/Resource
 A.1 Danbooru2023
-The Danbooru2023 dataset [69, 68, 46] is an extensive collection of images and their corresponding tags, compiled from the Danbooru image board.
+The Danbooru2023 dataset (Yeh, 2024a; b; nyanko202, 2023) is an extensive collection of images and their corresponding tags, compiled from the Danbooru image board.
 This dataset includes images annotated with particular and detailed tags, providing a rich resource for training both the Text-to-Image (T2I) and Large Language Models (LLMs) involved in the TIPO framework.
@@
 A.2 GBC10M
-The GBC10M dataset [29] is a large-scale collection of 10 million images sourced from CC12M [13], annotated using the Graph-Based Captioning (GBC) approach.
+The GBC10M dataset (Hsieh et al., 2024) is a large-scale collection of 10 million images sourced from CC12M (Changpinyo et al., 2021), annotated using the Graph-Based Captioning (GBC) approach.
 Each image is represented by a graph where nodes correspond to object regions, compositions, and relations, and edges define their hierarchical relationships.
@@
 A.3 Coyo HD 11M
-The Coyo HD 11M dataset [12] consists of 11.4 million high-resolution, high-concept-density images paired with 22.8 million synthetic captions generated from the Coyo-700M dataset.
+The Coyo HD 11M dataset (CaptionEmporium, 2024) consists of 11.4 million high-resolution, high-concept-density images paired with 22.8 million synthetic captions generated from the Coyo-700M dataset.
 Images maintain a minimum of 512 pixels on the shortest edge to ensure high visual quality.
-Captions, generated with the LLaVA-Next-8B model [38] based on LLaMA 3 [3], undergo post-processing for conciseness and clarity.
+Captions, generated with the LLaVA-Next-8B model (Liu et al., 2024a) based on LLaMA 3 (AI@Meta, 2024), undergo post-processing for conciseness and clarity.
 TIPO uses short and long captions, booru tags, and open image tags from this dataset.
-A.4 Stable Diffusion XL
-Stable Diffusion XL (SDXL) [50] improves upon earlier models [56] with a more considerable UNet backbone and dual text encoders (CLIP ViT-L [51] and OpenCLIP ViT-bigG [31]), enhancing text conditioning.
+Appendix B Baselines/T2I models
+B.1 Prompt-Optimization Baselines
+GPT-4o-mini.
+GPT-4o-mini(OpenAI, 2024) is a multimodal model introduced by OpenAI in July 2024 as a cost-efficient variant of GPT-4o.
+It supports a 128k-token context window and up to 16k output tokens, trained on text and vision data.
+In this paper, it serves as a zero-shot rewriting baseline for prompt refinement.
+MagicPrompt.
+MagicPrompt(daspartho, 2022) fine-tunes GPT-2 models on large-scale, community-collected prompts (e.g., from Lexica.art and the Stable Diffusion Prompts dataset).
+It learns to generate extended prompts with richer descriptive content, originally designed to improve prompt quality for Stable Diffusion.
+Promptist.
+Promptist(Hao et al., 2023) is a reinforcement learning framework for prompt optimization.
+Starting from a seed prompt, it generates refined prompts using supervised pre-training and reinforcement learning, guided by aesthetic and semantic alignment rewards.
+It consistently improves model-preferred prompt quality when paired with diffusion backbones.
+Gemini-2.0-Flash-Image.
+Gemini-2.0-Flash-Image [^4] https://developers.googleblog.com/experiment-with-gemini-20-flash-native-image-generation/ is a variant of Google’s Gemini 2.0 Flash family with native image generation support.
+It provides both prompt-refinement and text-to-image generation within a single model, supporting high-fidelity text rendering, compositional control, and iterative editing.
+Unlike other baselines, it functions as both optimizer and generator.
+B.2 T2I Models
+SDXL
+Stable Diffusion XL (SDXL) (Podell et al., 2024) improves upon earlier models (Rombach et al., 2022b) with a more considerable UNet backbone and dual text encoders (CLIP ViT-L (Radford et al., 2021) and OpenCLIP ViT-bigG (Ilharco et al., 2021)), enhancing text conditioning.
 Supporting resolutions up to 1024×1024, SDXL accepts natural language prompts and tags, suitable for diverse image generation.
-A.5 Illustrious
+Illustrious
 Illustrious is a series of fine-tuned Stable Diffusion XL models primarily trained on the Danbooru2023 dataset.
-In this study, we specifically employ the v3.5 version variant with v-parameterization [58], which is notable for its extensive incorporation of natural language prompts.
+In this study, we specifically employ the v3.5 version variant with v-parameterization (Salimans and Ho, 2022), which is notable for its extensive incorporation of natural language prompts.
 The inclusion of both tag-based and natural language formats allows Illustrious to leverage a broad range of semantic knowledge for image generation.
 Within TIPO, we perform an ablation study to analyze the effectiveness of different prompting strategies—namely extended tags versus natural language prompts—to identify which approach contributes most significantly to enhanced image generation performance.
-A.6 Stable Diffusion 3.5
-Stable Diffusion 3.5 (SD-3.5) incorporates the MMDiT architecture [20] and the Rectified Flow formulation [39, 4, 36] for improved text-to-image generation.
-Utilizing triple text encoders (CLIP/ViT-L, OpenCLIP/ViT-G, T5-XXL [52]), SD-3.5 supports resolutions up to 1024×1024 and uses a 50/50 mix of original and CogVLM-generated captions.
+Stable Diffusion 3.5
+Stable Diffusion 3.5 (SD-3.5) incorporates the MMDiT architecture (Esser et al., 2024a) and the Rectified Flow formulation (Liu et al., 2022; Albergo and Vanden-Eijnden, 2022; Lipman et al., 2022) for improved text-to-image generation.
+Utilizing triple text encoders (CLIP/ViT-L, OpenCLIP/ViT-G, T5-XXL (Raffel et al., 2020b)), SD-3.5 supports resolutions up to 1024×1024 and uses a 50/50 mix of original and CogVLM-generated captions.
 Figures confirm the capability to process both natural language prompts and tags.
 This study employs SD-3.5-Large [^9] https://huggingface.co/stabilityai/stable-diffusion-3.5-large (8B parameters) with FP8 inference on RTX 3090 or RTX 4090 GPUs.
-Appendix B Implementation Details
-In this appendix, we provide all the necessary details including our dataset construction process, model configurations, inference pipeline and the model’s properties not mentioned in Section 4.2 and 4.3.
-B.1 Training Data Construction
+FLUX.1-dev
+FLUX.1-dev is an open-weights text-to-image model released by Black Forest Labs as a guidance-distilled variant of their FLUX.1 family, targeting research and non-commercial use (black-forest-labs, 2024).
+Architecturally, FLUX adopts a transformer-based rectified-flow formulation (Lipman et al., 2022) with a hybrid stack of multimodal/parallel diffusion transformer blocks and rotary position encodings, scaled to $\sim$12B parameters, emphasizing prompt adherence, typography, and aspect-ratio flexibility (black-forest-labs, 2024).
+We use the FLUX.1-dev checkpoint for T2I without additional refiners; it accepts both natural-language prompts and tag-like inputs and supports resolutions in the 0.10.12.02.0 MP range.
+[^10] https://huggingface.co/black-forest-labs/FLUX.1-dev
+OmniGen2
+OmniGen2 is a unified multimodal generator that decouples autoregressive text modeling from diffusion-based image generation via two distinct pathways with unshared parameters (Wu et al., 2025).
+The diffusion side conditions on hidden states from the MLLM while exclusively feeding VAE features into the diffusion decoder to preserve low-level fidelity; a 3D rotary scheme (Omni-RoPE) disentangles sequence ID and 2D spatial coordinates to stabilize editing and in-context generation (Wu et al., 2025).
+Beyond standard T2I, OmniGen2 natively supports image editing and subject-driven in-context generation, and introduces a reflection mechanism/dataset to iteratively refine outputs.
+Lumina-Image 2.0
+[…]
+HiDream-I1
+HiDream-I1 is a 17B-parameter image foundation model based on a sparse Diffusion Transformer with dynamic Mixture-of-Experts (MoE) (Cai et al., 2025).
+It employs a dual-stream (text/image) sparse DiT for separate encoding followed by a single-stream sparse DiT to fuse modalities efficiently; hybrid text encoders (e.g., CLIP-L/14, CLIP-G/14, T5-XXL) and an LLM aggregator provide robust conditioning (Cai et al., 2025).
+The suite includes I1-Full (50+50{+} steps), I1-Dev (guidance-distilled, 28 steps), and I1-Fast (14 steps), with a GAN-powered diffusion distillation to retain sharpness at low step counts.
+Appendix C TIPO Implementation Details
+In this appendix, we provide all the necessary details including our dataset construction process, model configurations, inference pipeline, and the model’s properties not mentioned in Section 4.2 and 4.3.
+C.1 TIPO Training Data Construction
 This section details our methodology for constructing and preprocessing training data to ensure robust model performance across various input scenarios.
@@
 These augmentation strategies create a more diverse training dataset that better reflects real-world prompt variations, improving the model’s robustness and adaptability to different input styles and formats.
-B.2 Training Settings and Model Configurations
+C.2 TIPO Training Settings and Model Configurations
 TIPO-100M TIPO-200M stage1 TIPO-200M stage2 TIPO-500M
@@
 The datasets include CoyoHD11M (Coyo), GBC10M (GBC), and Danbooru2023 (Dan).
-Stage 2 additionally incorporates Pixtral [1] to generate NL captions from Danboodu2023 dataset.
+Stage 2 additionally incorporates Pixtral (Agrawal et al., 2024) to generate NL captions from Danbooru2023 dataset.
 Tokenizer and Task Tokens
-TIPO employs a vocabulary derived from LLaMA2 [65] consisting of 32,000 tokens, with additional tokens (13 tokens) specifically designated for task and length control or placeholders.
+TIPO employs a vocabulary derived from LLaMA2 (Touvron et al., 2023b) consisting of 32,000 tokens, with additional tokens (13 tokens) specifically designated for task and length control or placeholders.
 This extended vocabulary includes task identifiers and length modifiers to ensure flexibility across different prompt types:
@@
 Optimizer and Learning Schedule
-Training is performed using the AdamW optimizer [40], with a cosine annealing learning rate scheduler [41].
+Training is performed using the AdamW optimizer (Loshchilov, 2017), with a cosine annealing learning rate scheduler (Loshchilov and Hutter, 2017).
 The optimizer parameters include $\beta_{1}=0.9$, $\beta_{2}=0.98$, and a weight decay of 0.01.
@@
 Table 8 summarizes the configurations for pretraining and fine-tuning TIPO-100M, TIPO-200M, and TIPO-500M.
-Both pretraining and fine-tuning was conducted on datasets like Danbooru2023 [46], GBC10M [29], and CoyoHD11M [12].
+Both pretraining and fine-tuning was conducted on datasets like Danbooru2023 (nyanko202, 2023), GBC10M (Hsieh et al., 2024), and CoyoHD11M (CaptionEmporium, 2024).
 Augmented Task Representation
@@
 Figure 6: TIPO inference workflow, with solid arrows denoting the primary generation steps and dashed arrows indicating alternative generation paths within the same cycle.
-¡TOKEN¿— represents special tokens, with all tokens detailed in Section B.2.
-B.3 Inference Pipeline
+¡TOKEN¿— represents special tokens, with all tokens detailed in Section C.2.
+C.3 TIPO Inference Settings
+Sampling Strategy
+We employ a hybrid stochastic decoding strategy combining nucleus sampling (top-$p$ = 0.95) and top-$k$ = 60 filtering.
+This follows standard practice in open-ended text generation, as adopted in the official Hugging Face generation examples  [^11] Hugging Face.
+“Usage — transformers 2.11.0 documentation.” Example of text generation with XLNet uses top-$p$ = 0.95 and top-$k$ = 60.
+https://huggingface.co/transformers/v2.11.0/usage.html.
+This hybrid approach maintains diversity while preserving coherence, preventing both overly deterministic and excessively noisy generations.
+Inference Pipeline
 The TIPO inference pipeline is designed to handle various input types and scenarios, combining different tasks to refine or expand both tag-based and natural language prompts.
@@
 By leveraging different task combinations, TIPO ensures that tag-based and natural language prompts are optimized, resulting in more detailed and effective input for text-to-image models.
-B.4 Model Speed Comparison
+C.4 Impact of Model Size on Performance
+To analyze the impact of model scales on prompt-optimization performance, we compare TIPO-200M and TIPO-500M using a 1,000-image subsample from the COYO and GBC datasets. Results are shown in Table 9.
+Table 9: Prompt optimization performance of TIPO-200M and TIPO-500M on a 1k subsample.
+Metric Task TIPO-200M TIPO-500M
+FDD (↓) NL-short 0.1529 0.1356
+NL-long 0.1650 0.1398
+Aesthetic (↑) NL-short 5.8531 $\pm$ 0.7501 5.8943 $\pm$ 0.7064
+NL-long 5.8364 $\pm$ 0.7501 5.9030 $\pm$ 0.7015
+AI Corrupt (↓) NL-short 0.2870 $\pm$ 0.4167 0.2891 $\pm$ 0.4189
+NL-long 0.2870 $\pm$ 0.4150 0.2862 $\pm$ 0.4151
+Overall, TIPO-500M shows consistent gains in FDD and Aesthetic scores, while performance on AI Corrupt remains comparable.
+However, the larger 500M variant entails substantially higher computational cost without delivering proportionally greater benefits, which limits its practicality for community use; hence, all main experiments are conducted with TIPO-200M.
+C.5 Impact of Model Size on Inference Speed
 We conducted comprehensive speed tests of our 100M, 200M and 500M parameter models using the inference pipeline described in Section 5.2.
@@
 Our primary metric is average tokens generated per second, which reflects real-world task performance rather than theoretical maximum throughput.
-The evaluation was performed using llama.cpp [23], an efficient C++ implementation that provides optimized support for various hardware accelerators including CUDA, HIP, and Apple Metal.
+The evaluation was performed using llama.cpp (Gerganov, 2023), an efficient C++ implementation that provides optimized support for various hardware accelerators, including CUDA, HIP, and Apple Metal.
 TIPO-100M TIPO-200M TIPO-500M
@@
 Tokens per second (tok/sec) represents the average generation speed, while generation time (gen time) shows the average time in seconds required for a complete two-step prompt optimization process.
-Appendix C Evaluation Statistics
+Appendix D Evaluation Statistics
 In this appendix, we provide more statistics for the result obtained in Section 5.
-C.1 In-domain test regarding scenery tag
+D.1 In-domain test regarding scenery tag
 (a) The box plot for the Aesthetic Score result of scenery tag test.
@@
 The analysis shows that TIPO significantly outperforms all other methods, demonstrating a considerable margin of improvement.
-C.2 In-domain prompt generation test
+D.2 In-domain prompt generation test
 (a) The box plot for the Aesthetic Score result of short prompt input.
@@
 Figure 9: The distribution of aesthetic and AI corrupt score for truncated long prompt input.
-Figures 11 and 12 display the box plots and KDE plots of aesthetic scores and AI corruption scores obtained from the In-domain prompt generation test detailed in Section E.2.
+Figures 8 and 9 display the box plots and KDE plots of aesthetic scores and AI corruption scores obtained from the In-domain prompt generation test detailed in Section F.2.
 While the box plots reveal subtle differences in performance between various methods, the AI corruption scores provide valuable insights.
 Specifically, these scores indicate that implementations supported by TIPO produce more stable output images than other methods.
-C.3 Out-of-domain evaluation
+D.3 Out-of-domain evaluation
 (a) The box plot for the Aesthetic Score result of out-of-focus test.
@@
 Figure 12: The similarity matrix between 100 images of worst aesthetic generated results of SD3.5-Large experiments.
-Figures 14 and 15 present similarity matrices for different prompt generation methods and their corresponding aesthetic outputs on SD3.5-Large  [21].
+Figures 11 and 12 present similarity matrices for different prompt generation methods and their corresponding aesthetic outputs on SD3.5-Large  (Esser et al., 2024b).
 A matrix with predominantly lower similarity values (brighter appearance) indicates high diversity among generated images, while higher values (darker appearance) suggest consistent but less diverse outputs.
 Please refer to Table 2 in Section 5.
-C.4 Ablation Test
+D.4 Ablation Test
 (a) The box plot for the Aesthetic Score result of NL ablation test.
@@
 Both of them are better than the original tag, which indicates that the tag TIPO aspect helps control corruption and promotes the aesthetic score.
-Appendix D TIPO example
+Appendix E TIPO example
 In this section, we provide some text example of TIPO’s input and output.
-TIPO Format Template User Input: 1girl, ciloranko, maccha (mochancc), ningen mame, ask (askzy), solo, masterpiece, absurdres, newest, safe A girl sits in a cozy cafe, cradling a cup of coffee in her hand Formatted TIPO Input for Expand Tags: meta: absurdres rating: safe artist: ciloranko, maccha (mochancc), ningen mame, ask (askzy) quality: masterpiece, newest aspect_ratio: 1.0 target: <|short|> <|short_to_tag|> short: A girl sits in a cozy cafe, cradling a cup of coffee in her hand tag: 1girl, solo Formatted TIPO Output after Expand Tags and Expand Natural Prompt: meta: absurdres rating: safe artist: ciloranko, maccha (mochancc), ningen mame, ask (askzy) quality: masterpiece, newest aspect_ratio: 1.0 target: <|short|> <|tag_to_long|> tag: 1girl, solo, sitting, closed mouth, jewelry, long hair, looking at viewer, crossed legs, plant, table, couch, bracelet, cup, smile, teacup, indoors, blue eyes, blonde hair, holding long: A girl sits in a cozy cafe, cradling a cup of coffee in her hand.
+TIPO Format Template User Input: 1girl, ciloranko, maccha (mochancc), ningen mame, ask (askzy), solo, masterpiece, absurdres, newest, safe A girl sits in a cozy cafe, cradling a cup of coffee in her hand Formatted TIPO Input for Expand Tags: meta: absurdres rating: safe style: anime-style illustration, digital art, character design, fantasy concept art quality: masterpiece, newest aspect_ratio: 1.0 target: <|short|> <|short_to_tag|> short: A girl sits in a cozy cafe, cradling a cup of coffee in her hand tag: 1girl, solo Formatted TIPO Output after Expand Tags and Expand Natural Prompt: meta: absurdres rating: safe style: anime-style illustration, digital art, character design, fantasy concept art quality: masterpiece, newest aspect_ratio: 1.0 target: <|short|> <|tag_to_long|> tag: 1girl, solo, sitting, closed mouth, jewelry, long hair, looking at viewer, crossed legs, plant, table, couch, bracelet, cup, smile, teacup, indoors, blue eyes, blonde hair, holding long: A girl sits in a cozy cafe, cradling a cup of coffee in her hand.
 The cafe has large windows with green plants on the walls and a wooden table in front of her.
@@
 Figure 16: An example formatted content we used for training and inference in TIPO.
-Appendix E Image Examples
+Appendix F Image Examples
 In this section, we present sample images from the experiments described in Section 5 to visually demonstrate the improvements achieved by TIPO.
-E.1 In-domain test regard to scenery tag
+F.1 In-domain test regard to scenery tag
 Figure 17: Comparison of generated images using simple input (left) vs. TIPO-enhanced input (right) for the scenery tag
 Figure 17 demonstrates the difference in output diversity between simple input and TIPO-enhanced input for the scenery tag.
-As observed, TIPO significantly expands the range of generated sceneries, better reflecting the variety present in the Danbooru2023 dataset [69].
+As observed, TIPO significantly expands the range of generated sceneries, better reflecting the variety present in the Danbooru2023 dataset (Yeh, 2024a).
 The left column shows results from simple input (scenery tag only), while the right column illustrates the enhanced diversity achieved with TIPO-enhanced input.
-E.2 In-domain prompt generation test
+F.2 In-domain prompt generation test
 (a) Short Caption
@@
 However, the use of TIPO to refine or extend these prompts still yields noticeable improvements in aesthetics and overall quality.
-Appendix F Human Preference
+Appendix G Human Preference
 (a) ELO rating on Illustrious
@@
 Original vs. PromptExtend 45:66 0.405 0.0572 Marginally
-Original vs. Tipo 28:80 0.259 $<0.0001$ Yes***
+Original vs. Tipo 28:80 0.259 <0.0001<0.0001 Yes***
 Original vs. Promptist 35:64 0.354 0.0046 Yes**
 PromptExtend vs. Promptist 41:80 0.339 0.0005 Yes***
-Promptist vs. Tipo 30:101 0.229 $<0.0001$ Yes***
-PromptExtend vs. Tipo 30:90 0.250 $<0.0001$ Yes***
+Promptist vs. Tipo 30:101 0.229 <0.0001<0.0001 Yes***
+PromptExtend vs. Tipo 30:90 0.250 <0.0001<0.0001 Yes***
 MagicPrompt vs. Promptist 14:34 0.292 0.0055 Yes**
-MagicPrompt vs. Tipo 11:48 0.186 $<0.0001$ Yes***
+MagicPrompt vs. Tipo 11:48 0.186 <0.0001<0.0001 Yes***
 MagicPrompt vs. Original 15:28 0.349 0.0660 No
@@
 Each evaluation asked participants to compare two generated images—labeled A and B—and select which they preferred (or a tie) according to specific criteria (e.g., prompt adherence, image quality, or aesthetic appeal).
-F.1 User Interface for Human Preference Evaluation
+G.1 User Interface for Human Preference Evaluation
 We developed a specialized survey interface to facilitate efficient and unbiased human evaluation of generated images.
@@
 After submitting their evaluations, the interface reveals the transformed prompts used to generate each image, providing transparency about how the original prompt was modified by each method.
-F.2 Extended Human Evaluation.
+G.2 Extended Human Evaluation.
 Participants assessed each image’s performance on prompt adherence, image quality, and aesthetic appeal, with visually shown unmodified and image pairs.
@@
 Notably, it attained a 64.4% peak win rate (against MagicPrompt) under the Full scenario and 57.5% (also against MagicPrompt) under SD35-medium, emphasizing TIPO’s proficiency in generating images that closely follow prompt specifications while maintaining visual coherence.
-F.3 ELO Ratings.
+G.3 ELO Ratings.
 We computed theoretical ELO ratings from the aggregated pairwise comparisons to quantify overall performance differences among the five methods.
@@
 The result is depicted in Figure 19, TIPO has secured the highest ELO rating over other models.
-F.4 Human Preference ELO Method
+G.4 Human Preference ELO Method
 We computed theoretical ELO ratings from human-judged pairwise preference data to quantitatively evaluate the relative performance of each prompting method.
@@
 Human evaluators assessed comparisons between methods, resulting in one of three outcomes:
-  • Method $i$ wins: assigned a score of $1$ for method $i$, and 00 for method $j$.
-  • Method $j$ wins: assigned score $1$ for method $j$, and 00 for method $i$.
+  • Method $i$ wins: assigned a score of $1$ for method $i$, and 0 for method $j$.
+  • Method $j$ wins: assigned score $1$ for method $j$, and 0 for method $i$.
   • Tie: assigned score $0.5$ to both methods.
@@
 Win and tie rates were converted to ELO rating differences using:
-Adjusted Win Rate=Win Rate+Tie $\text{Adjusted Win Rate}=\text{Win Rate}+\frac{\text{Tie Rate}}{2}$
-ELO $\text{ELO Difference}=400\times\log_{10}\left(\frac{\text{Adjusted Win Rate}}{1-\text{Adjusted Win Rate}}\right)$
+$\text{Adjusted Win Rate}=\text{Win Rate}+\frac{\text{Tie Rate}}{2}$
+$\text{ELO Difference}=400\times\log_{10}\left(\frac{\text{Adjusted Win Rate}}{1-\text{Adjusted Win Rate}}\right)$
 To ensure numerical stability, extreme adjusted win rates were constrained as follows:
-ELO $\text{ELO Difference}=\begin{cases}-800,&\text{Adjusted Win Rate}\leq 0.001\\ +800,&\text{Adjusted Win Rate}\geq 0.999\\ \end{cases}$
+[…]
 Calculating Method ELO Ratings.
 Final ELO ratings were determined by averaging each method’s pairwise ELO differences and centering these averages around a baseline rating (e.g., $1000$):
-$\text{ELO}_{\text{method}_{i}}=\text{Base Rating}+\left(\text{Average ELO Difference for method }i-\text{Overall Mean ELO Difference}\right)$
+[…]
 Interpretation of ELO Scores.
 Methods with higher ELO scores consistently outperform lower-scored methods. A rating difference of 400 points corresponds to a 90% expected win probability for the superior method.
-F.5 Statistical Significance
+G.5 Statistical Significance
 As summarized in Table 11, we conducted two-sided binomial and McNemar’s tests (p <0.05) to assess the statistical significance of observed differences.
@@
 Collectively, these findings illustrate TIPO’s robust, model-agnostic effectiveness and underscore the model-sensitivity of alternative methods, particularly Promptext and MagicPrompt.
-F.6 Survey Response Examples
+G.6 Survey Response Examples
 In this section we provided some responses of our human preference survey as reference.
@@
 Figure 25: Some survey responses on SD3.5-medium generated image with different prompt optimization method
-F.7 Conclusion
+G.7 Conclusion
 The extended evaluations presented here reinforce TIPO’s standing as a reliable and effective prompt-optimization strategy.
 Its consistent performance gains under diverse model conditions highlight its potential for broad applicability, with its strong alignment with user-specified prompts, high image quality, and favorable aesthetic outcomes.
-Appendix G Ablation Study on TIPO
+Appendix H Ablation Study on TIPO
 In this section, we investigate the effect of incorporating TIPO (Tags + Inferred Prompt Objects) across various generation settings.
 Our primary goal is to validate whether additional structured information (e.g., core tags and minimal spatial/contextual cues) can improve image quality, reduce artifacts.
-G.1 Experimental Setup
-Prompt Variants.
+H.1 Experimental Setup
+Prompt Variants
 To systematically analyze TIPO’s contribution, we consider four types of input prompts improvement task:
@@
 In each case, we compare the baseline prompts (without TIPO cues) against prompts incorporating TIPO’s structured, tag-based critical information and minimal spatial hints.
-Data Preparation.
+Data Preparation
 We start by randomly sampling core words from a word table to represent a diverse range of topics (e.g., objects, environments, descriptors).
@@
 Overall, the six prompt variants are tested on 4,000 images, ensuring a balanced comparison.
-Inference Procedure.
-Prompts are fed into our image-generation pipeline under identical model settings (classifier free guidance, sampler, steps, etc.), using the v-parameterized variant of Illustrious v3.5[49].
+Inference Procedure
+Prompts are fed into our image-generation pipeline under identical model settings (classifier free guidance, sampler, steps, etc.), using the v-parameterized variant of Illustrious v3.5(Park et al., 2024).
 We focus on how TIPO modifications alter the generation outcomes and whether they introduce additional computational overhead.
-G.2 Evaluation Metrics
+H.2 Evaluation Metrics
 (a) Corrupt Score Distributions
@@
 Figure 26: Side-by-side comparison of Corrupt (left) and Aesthetic (right) score distributions across prompt types.
-Aesthetic Score.
+Aesthetic Score
 We employ an off-the-shelf aesthetic predictor to estimate image quality. In the following paragraph, we discuss the model’s bias.
-AI Corruption Score.
+AI Corruption Score
 Using an automated ’ AI corruption ’ detection model, we measure generation artifacts, such as distorted objects and unnatural shapes. Higher scores imply cleaner, more coherent outputs.
-G.3 Results & Discussion
-Impact on Aesthetics.
+H.3 Results & Discussion
+Impact on Aesthetics
 Figure 26(b) shows that TIPO-enhanced prompts generally achieve higher aesthetic scores than their non-TIPO counterparts, albeit with some variance.
 Notably, we observe a correlation between wider color ranges and higher aesthetic scores discussed in Figure 27, suggesting a bias toward more colorful or varied compositions.
-Improvements in Corruption Score.
+Improvements in Corruption Score
 As shown in Figure 26(a), TIPO-based prompts yield significantly lower corruption scores, indicating fewer artifacts.
@@
 However, saturation shows no notable correlation with corrupt score (Pearson $r=0.0125$).
-G.4 Speed Test and Overhead Analysis
+H.4 Speed Test and Overhead Analysis
+Inference Speed
 A key concern for production pipelines is whether TIPO generation imposes a substantial time overhead.
@@
 MAGICPROMPT GPT2-125M 1000 1.1398 0.4043
-Memory Footprint.
-We also confirm that TIPO’s overhead in terms of VRAM usage is minimal. The practical adoption of TIPO in pipelines has shown no critical memory concerns, which aligns with our measurements.
-G.5 Conclusion of Ablation
+Memory Footprint
+We also confirm that TIPO’s overhead in terms of VRAM usage is minimal (e.g., << 0.5 GB for TIPO-200M and << 1.5 GB for TIPO-500M) with the quantization supported by llama.cpp.
+The practical adoption of TIPO in pipelines has shown no critical memory concerns, which aligns with our measurements.
+Training Costs
+[…]
+Although TIPO requires more total GPU-hours than PAE, it is trained on over 30 million prompts—two orders of magnitude larger than both Promptist and PAE.
+After normalization, TIPO achieves the lowest cost per 1k prompts, demonstrating strong scalability.
+It is also worth noting that Promptist and PAE rely on reinforcement learning with external T2I rollouts.
+Even for SD1.5, each rollout takes roughly five seconds, and the cost increases dramatically for larger models such as SDXL, SD3, or Flux.
+By contrast, TIPO requires no rollouts, so its training cost scales linearly with corpus size and remains independent of the target T2I model.
+Table 13: Training cost comparison between TIPO and RL-based prompt optimization methods. GPU-hours per 1k prompts are normalized for fairness.
+Method #Params #Prompts #GPUs GPU-h GPU-h /1k
+[…]
+[…]
+[…]
+H.5 Conclusion of Ablation
 Our experiments suggest that TIPO (1) consistently lowers AI corruption artifacts, (2) can boost aesthetic scores, and (3) remains computationally inexpensive.
@@
 In short, a concise natural language prompt with core tag-based critical information appears to be an effective, suggested form for most use cases.
-Appendix H Topic Distribution Visualization
-Latent Dirichlet Allocation (LDA) [9] is a generative probabilistic model for topic modeling [32], which assumes that each document is a mixture of topics, with each topic represented by a distribution over words.
-LDA uncovers hidden thematic structures by analyzing word co-occurrence patterns, while methods like TF-IDF and TextRank [43] enhance its ability to extract meaningful insights from large textual datasets.
+Appendix I Topic Distribution Visualization
+Latent Dirichlet Allocation (LDA) (Blei et al., 2003) is a generative probabilistic model for topic modeling (Jelodar et al., 2018), which assumes that each document is a mixture of topics, with each topic represented by a distribution over words.
+LDA uncovers hidden thematic structures by analyzing word co-occurrence patterns, while methods like TF-IDF and TextRank (Mihalcea and Tarau, 2004) enhance its ability to extract meaningful insights from large textual datasets.
 We implemented a multi-stage topic modeling and clustering methodology using LDA to extract varying numbers of topics (20, 30, 50, and 100) from the corpus.
@@
 We empirically assessed whether the resulting topics were sufficiently large and diverse by employing multi-level topic analysis.
-This iterative process mitigates potential challenges such as substantial topic overlap, which can diminish distinctiveness when extracting a large number of topics [62].
-To address the potential overlap and further assess the diversity and meaningfulness of the topics, we performed a secondary clustering [70].
+This iterative process mitigates potential challenges such as substantial topic overlap, which can diminish distinctiveness when extracting a large number of topics (Stevens et al., 2012).
+To address the potential overlap and further assess the diversity and meaningfulness of the topics, we performed a secondary clustering (Zhao and Karypis, 2002).
 We grouped the initially extracted topics into five major clusters using k-means clustering.
-We evaluated the clustering performance by calculating the inertia (Sum of Squared Distances) [27], shown in Table 10, 11, and 12.
+We evaluated the clustering performance by calculating the inertia (Sum of Squared Distances) (Hartigan et al., 1979), shown in Table 14, 15, and 16.
 Since the topics have already been filtered for meaningful content, a higher inertia value indicates greater diversity among the clusters, reflecting a broader range of valid and meaningful topics across the dataset.
@@
 100 1656.74 1245.60 1987.63 1628.32 877.36 657.14 1622.79 1777.61
+Z
 Table 14: Inertia for COYO-Dataset inference, higher is better
@@
 The analysis, conducted using Gephi, focuses on nodes with a degree greater than 600 to highlight critical components.
-Nodes are color-coded by modularity class by Fast Unfolding Algorithm [10], revealing clusters of closely associated tags.
-Node size reflects Eigenvector Centrality[11], emphasizing highly connected and influential tags within their network.
-Appendix I Discussion and Future Work
-Despite the promising performance demonstrated by TIPO, several areas merit further investigation:
-Personalization and User Adaptation.
-TIPO currently does not directly incorporate personalization methods (e.g.
-LoRA [30]).
-Incorporating such modules could further align generated outputs to individual user preferences, enabling more tailored creative assistance.
-Investigating joint training strategies that integrate personalized fine-tuning represents a valuable future direction.
-Generalization to Uncommon Compositions.
-While TIPO performs effectively on typical and broadly represented concepts, it may still struggle with highly unconventional combinations (e.g., “avocado chair”).
-Future research should systematically evaluate TIPO’s performance on rare or unusual prompts (but considered creative prompts), potentially exploring auxiliary training or adaptation methods to ensure broader compositional robustness.
-Capacity and Backbone Scalability.
-Given TIPO model’s architecture [66, 65], its capacity for handling highly complex or lengthy prompts could be inherently limited.
-Future work could investigate more extensive or specialized architectures enhance compositional and semantic capacity, ideally without compromising real-time inference efficiency, are crucial for practical deployment.
+Nodes are color-coded by modularity class by Fast Unfolding Algorithm (Blondel et al., 2008), revealing clusters of closely associated tags.
+Node size reflects Eigenvector Centrality(Bonacich, 1972), emphasizing highly connected and influential tags within their network.
+Appendix J Discussion and Future Work
+Despite the promising performance demonstrated by TIPO, several limitations and future directions remain open for further study.
+Distribution Dependence and OOD Generalization.
+The optimization behavior of TIPO inherently depends on the distributional bias of the open text–image corpus used for training.
+As a result, it may exhibit unstable behavior on extremely rare or stylistically unconventional prompts.
+Besides, when applied to out-of-distribution (OOD) T2I models whose training data or prompting conventions deviate significantly from LAION-style distributions, aesthetic performance may degrade slightly.
+In addition, our current OOD evaluation uses GPT-4o-mini–generated prompts, which are high-quality but do not fully represent real-world user queries, thus limiting external validity.
+Addressing the generalization to unseen models and long-tail prompts remains an important direction for future work, which can be pursued from two complementary perspectives: improving the model’s generalization capacity via stronger backbones and domain-diverse fine-tuning, and enhancing evaluation through large-scale collection of authentic user prompts and cross-model benchmarking.
+Stronger Backbone Initialization.
+Our current implementation trains a mid-sized LLaMA variant from scratch.
+Future work could instead initialize from stronger open-source LLM backbones and fine-tune them on T2I corpora, potentially improving robustness on long-tail and domain-specific distributions.
+This could also mitigate failures on small or highly biased datasets by leveraging more general linguistic priors.
+Model-Specific Adaptation and Style Variance.
+For models that require structured or JSON-formatted prompts, TIPO can be combined with a lightweight adapter or fine-tuned on a small set of model-specific data.
+Extending TIPO with such adaptation modules could better accommodate systems whose prompt syntax diverges from mainstream diffusion models.
+In the longer term, TIPO can also serve as a backbone integrated with RL-based refinement for model-specific alignment.
+Personalization and Style Preservation.
+The current TIPO is a general-purpose optimizer and does not incorporate user-level or stylistic conditioning.
+Building on prior personalization techniques such as LoRA (Hu et al., 2022), future work could explore lightweight adapters or online learning mechanisms that track user preferences and maintain project-level stylistic consistency, enabling personalized and context-aware prompt optimization.
+Image-Aware and Feedback-Driven Refinement.
+TIPO currently operates purely in the text domain without utilizing generated images or user feedback.
+A promising extension is to incorporate vision–language models like Qwen3-VL (Bai et al., 2023) for image-aware refinement, allowing iterative refinement with optimized prompts, visual outcomes, and user instructions.
+However, such integration requires non-trivial data curation and training pipelines, which we leave for future work.
+Scaling Behavior and Model Capacity.
+From 200M to 500M parameters, TIPO continues to yield improvements in FDD and Aesthetic scores.
+Due to limited compute resources, we could not explore larger configurations to observe scaling saturation.
+A systematic study of TIPO’s scaling behavior and architectural variants would be a valuable direction for future research.
+Such analysis could reveal scaling laws unique to prompt optimizers and guide practical model sizing for future deployments.
+Appendix K Disclosure of LLM Usage
+We used GPT-5 only to polish writing by improving the readability and grammar correctness.
+No LLMs were used in the main contributions of this work, such as ideation, experiment design, or result analysis.